A Complete Bibliography of Publications in Isis, 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.06

Title word cross-reference

+2000 [570]. −2500 [570].

1 [722, 880, 711]. 1165 [1340]. 1305 [702]. 1519 [368]. 1540 [138]. 1548 [1087]. 1564 [753]. 1618 [292]. 1627 [336]. 1630 [729]. 1641 [542]. 1675 [161]. 1687 [162, 481]. 1692 [372]. 1694 [513]. 16th [763]. 16th-Century [763]. 1711 [499]. 1711-ig [499]. 1720 [755]. 1757 [409]. 1763 [334]. 1764 [929]. 1766 [763]. 1767 [160]. 1772 [1290, 411]. 1790 [439]. 1794 [854]. 1795 [140]. 18.Jahrhundert [32]. 1800 [652, 1311, 953]. 1803 [216]. 1805 [629]. 1809 [682]. 1820 [158, 505]. 1828 [1346]. 1830 [1353]. 1838 [1355]. 1840 [959, 274]. 1848 [1088]. 1850 [1149]. 1851 [932]. 1854 [764]. 1855 [717]. 1858 [687]. 1859 [51]. 1860 [1266, 221, 610, 1312]. 1865 [376]. 1866 [736]. 1870 [75]. 1873 [74]. 1880 [202, 760]. 1885 [848]. 1891 [1193]. 1894 [1315]. 1896 [653]. 1900 [776, 995, 971, 476, 1140, 95]. 1902 [1181]. 1904 [1326]. 1907 [485]. 1912 [885]. 1913 [1313]. 1915 [552]. 1918 [886]. 1919 [276]. 1920 [1018]. 1925 [103]. 1926 [1021]. 1927 [1161]. 1930 [844]. 1932 [340, 1246, 685]. 1935 [1056]. 1937 [305]. 1940 [721]. 1944-ig [719]. 1945 [689]. 1950 [1360, 1317, 1247]. 1951 [1041]. 1951/1952 [1041]. 1952 [1041]. 1957 [516, 268, 286]. 1958/1962 [707]. 1959 [35, 191]. 1960 [142, 78, 796, 346, 442]. 1961 [166]. 1963 [801, 1022]. 1965 [1266, 1075]. 1966 [1265, 1209, 1142]. 1969 [1325]. 19th [373, 1214].

20th [373, 1214]. 250th [483].

467 [1109].

59 [165].

7 [676].

83 [730].

A-E. [1143]. A. [946]. A.D [1123, 476, 510, 479, 1085, 1069]. A.D. [702, 544, 565]. aan [1092]. Aaron [761, 904]. Abbe [925]. ’Abdallah [26]. Abdul [478]. Abdul-Kader [478]. Abich [1040]. Abraham [1021, 337]. Abteilung [369]. ’Abu [26, 967, 25, 918]. Abulcasis [908]. Academiae [1013, 1014]. Academician [8]. Academie [732]. Academies [652]. Academy [1056, 1246, 1178]. Accelerators [1057]. Acceptance [559]. according [603]. Account [932, 588]. Achievement [466]. Achievements [319, 890]. Acid [743]. Acidum [991]. Ackerknecht [1088]. Ackerman [472]. Act [841]. Action [885, 251, 350]. activité [453]. Activities [1032]. Ad [966, 26]. Ad-Dinawari [26]. Adam [219, 548]. Additional [801]. Address [146]. Adhan [908]. Administration [784, 1022]. Administrative [233, 167, 348, 79, 1022]. Admiralty [1356]. Adnan [285]. Adolf [641]. Adriaan [829]. Adulteration [1005]. Advancement [1202]. Adversus [335]. Aerial [525]. Aerodynamics [1009]. Aeronautics [784, 1022]. Aesthetics [59]. aevi [1085]. AFCRL [945]. Affair [1203]. affaire [1203]. Africa [1046]. against [1030]. Agassi [536]. Agassiz [628, 514, 343]. Age [579, 52, 418, 850, 617, 1154, 712, 1127, 1300, 613, 1084]. Ages [597, 965, 1064, 545, 68]. Agricola [70]. Agricultural [1152]. Agriculture [1284]. Aharon [549]. Ahmad [918, 390]. Air [1290]. Aitchison [278]. Aiton [463]. Akademie [160, 334]. Akademii [216]. al [1261, 406, 908, 918, 651, 388, 698, 789, 354, 390, 603, 865, 918, 135, 285, 478, 1261, 918]. al-Biruni [603, 285]. al-buldan [135]. al-Din [406]. al-Fusul [918]. al-Hasan [918]. al-Haythams [651]. al-Hindi [918]. al-Jwarizmi [390]. al-Kashi [406, 478]. al-Khwarizmi [390, 1261]. Al-Kindi [698, 789]. al-Mutanna’ [390]. al-Muthanna [1261, 390]. al-Shatir [354, 865]. al-Uqlidisi [918]. al-Zahrawi [908]. Alan [158, 855]. Alani [44]. Alaska [971]. Albert [1306, 515, 1059, 417, 1341]. Albertus [1273]. Alchemy [1112, 1034, 1123]. alchimistes [99]. Aldrovandi [621, 621]. Alex [1080]. Alexander [1269, 610, 105, 1274, 924, 1341, 57, 640, 1060, 1133, 758, 148, 1244, 56, 484, 55]. Alexandre [1235, 99, 363, 861]. Alexandrien [1339]. Alfred [467, 780, 813, 455]. Algebra [828, 967]. Algebre [589]. Algorismus [745]. Alhazen [1028]. Alien [1297]. Aligned [664]. alio [663]. aliquo [148]. Alistair [672]. aljabr [967]. alla [847]. Allard [121]. Alle [481]. allemand [99]. Allen [1128]. Allesandro [1252]. Alphonse [61]. Alt [1115]. Alt-Augsburger [1115]. alten [122]. altera [810]. Altholz [995]. Alvan [1277]. Alvar [52]. Alvin [577]. Amacher [964]. Ambroise [847]. Ambrose [371, 881]. Ameisenowa [225]. America [1266, 815, 1004, 1082, 221, 165, 1049, 735, 344, 634, 1202, 754, 402, 1276, 734, 688, 274]. American [647, 1216, 1266, 776, 628, 418, 300, 809, 299, 1073, 577, 970, 1076, 505, 889, 1147, 623, 976, 254, 241, 84, 695, 1321, 149, 1138, 201, 1337, 1127, 590, 1183]. Amos [71]. Ampère [1229]. Amphorismes [770]. Analecta [731]. analitico [1186]. Analogies [1170]. Analysis [832, 386, 252, 593]. Analytic [1021]. Analytical [1334]. Ananias [1102]. Anatomica [1340]. Anatomical [327]. Anatomie [364, 511]. Anatomy [930, 1064, 137, 138, 928, 799, 73]. Anaxagoras [1109, 806]. Anaximander [544, 507, 328]. Ancient [1177, 616, 1135, 676, 1083, 129, 726, 63, 42, 440, 668, 725, 194, 1271]. Anderson [139]. Andersson [653]. Andrade [564]. Andre [105]. Andreas [1055, 511, 847, 753, 138]. Andreas-Salome [1055]. Andrew [324, 1351]. Andrews [631]. Andrzej [225]. Angelo [1186]. Anglo [1262, 678]. Anglo-Saxon [1262, 678]. Angular [827]. Angus [643]. Animal [1135, 798, 1198, 377, 911, 574]. Animalium [336, 1068]. Animaux [331, 1191, 551]. Anne [571]. Anniversary [483]. Annotated [1149, 812, 133]. Anonymi [697]. ans [1081]. Anstalten [276]. Antarctic [274]. Anthemius [217]. Anthony [1310]. Anthropology [938, 191, 645, 556, 845]. Anticipation [1232]. antiguedad [445]. Antike [1342, 364]. Antiken [195]. antiquite [447]. Antiquity [60, 672, 440, 619, 581]. Antoine [854]. Antologia [1357]. Antoni [481, 164]. Antonio [72]. Antony [715]. ’anwa’ [26]. any [148, 663]. Apeiron [507]. Apes [391]. Aphrodisias [148]. Apostolic [742]. Apothecaries [1268]. Apotheker [76]. Apparatus [601, 1311]. apparentium [1264]. apparitione [1264]. ’Appennino [878]. Application [911]. Applied [426]. Appraisal [1288]. Approaches [1299]. April [142]. Aquinas [808, 10]. ar-Razi [25]. Arabic [388, 918]. Arabische [1342]. Arabo [965]. Arabo-Latin [965]. Arabs [26]. Aram [197]. Archaeology [580, 725]. Archeological [1120]. archéologique [724]. Archimedes [207, 965]. Architect [742]. Architecture [539, 300, 1083]. Archive [1292, 532]. Archives [257, 669, 945, 461]. Archivi [1013, 1014]. Archivist [244]. Arctic [516]. arcu [1264]. Arcueil [1354]. ardente [786]. Argentina [531, 531]. Argles [932]. Arguments [1030]. Aristarchus [185]. Aristoteleos [335]. Aristoteles [450]. Aristoteliciens [335]. Aristotelis [1011]. Aristotle [615, 808, 1238, 1344, 618, 906, 282, 405]. Arithmetic [918]. Arithmetics [86]. Armenia [1102]. Armin [1090]. Armitage [643]. Armstrong [355, 516]. Armytage [952]. Arnald [941]. Arsbok [1142]. Art [1145, 201, 28, 224, 1332, 795, 471]. Artes [1067, 195]. Arthur [625, 754, 841, 1194, 270]. Articles [440]. Artist [1086]. Artists [1277]. Arts [1047, 472]. Arzt [281]. Arzte [1046]. Asa [686]. Ascanio [792]. Ascribed [506]. Ash [438]. Ashmole [1124]. Asimov [718, 756, 756]. Aspects [1211, 1114, 516, 199, 170]. Association [720]. Astrofizik [774]. Astrologia [1173]. Astrological [462]. Astrology [1066, 916, 603, 427]. Astronautics [784]. Astronom [774]. Astronomical [367, 1261, 588, 509, 407, 849, 462, 390]. Astronomicas [390, 583]. astronomie [760]. Astronomique [363]. astronomische [1115]. Astronomischen [443]. Astronomy [996, 40, 613, 1169, 1042, 603, 427, 346, 416, 503, 762, 1072]. Astrophysics [1110, 942]. Asylums [1275]. Atlantic [104]. Atlanticus [743]. Atmosphere [862]. Atmospheric [1070, 943]. Atom [1063, 960]. Atomic [689, 662, 1218, 1134, 1089, 1229, 657, 707, 826, 980]. Atomism [1026, 559, 492, 442, 693, 1349]. Atoms [1012, 1077]. Attempt [262, 351]. Attempts [741]. Attitudes [165]. Attributed [1291, 528]. Audubon [410]. Aufklaerung [296]. Auger [453]. Augmented [1012]. Augsburger [1115]. August [142, 374, 1161, 1174]. Augustis [1158]. Aurel [1133]. Aurele [1189]. Aurora [1065]. ausgewahlt [369]. Ausgewählte [70]. ausgewahlten [814]. Auslegung [369]. äusserlichen [116]. Austin [505]. Austrialia [1017]. Auswahl [369]. Autenticidad [390]. Authenticity [390]. Authorities [564, 564]. Authority [681]. Authorship [1070, 635, 528, 943]. Autobiographical [955, 1124, 1219]. Autobiography [421, 307, 1160, 1091]. Automata [1322]. Automation [617]. Automotive [301]. Autos [301]. aux [1191]. avant [1016]. Avenue [1313]. Aviation [1164]. Avogadro [1229]. Award [521]. Axel [1003, 474]. Axiomatization [834]. Aydin [273].

B [778, 1347, 654, 337, 568, 430, 303, 1245, 971, 594, 194, 121, 509, 1333, 1040, 802, 487, 1344, 1180, 505, 26, 515, 327, 847, 307, 933, 510, 889, 1147, 1162, 812, 963, 1326, 927, 856]. B.C [544]. B.C. [1109, 131]. Babbitt [957]. Babini [58, 58]. Baboons [391]. Babylon [613, 668, 320]. Babylonian [1035]. Back [36, 80, 107, 144, 168, 179, 205, 234, 238, 308, 347, 380, 422, 456, 489, 519, 553, 595, 632, 660, 691, 738, 785, 822, 859, 891, 912, 936, 974, 1024, 1061, 1094, 1097, 1130, 1166, 1195, 1221, 1225, 1248, 1280, 1318, 1361, 1365]. Background [909, 411, 678, 1114]. Bacon [586, 290, 867]. Badawy [1083]. Bagrow [714]. Bahnbrecher [685]. Baird [628]. Bakelants [511]. Bakr [25]. Balbin [531]. Baldasar [138]. Baldner [561]. Baliani [1242]. Ballenegger [719]. Baltzer [552]. Band [587, 810, 369, 1156, 1240, 484]. Bank [1279]. Banks [526]. Banse [55]. Baptiste [1242]. Barbara [1019, 1311, 1040]. Barlow [1312]. Barnabas [1187]. Barnett [585]. Barometer [384, 642]. Baron [1185]. Barrett [886]. Bartholomew [548]. Barton [1255]. Based [910]. Basel [832]. Basic [424]. Basis [3]. Basle [125]. Bates [605]. Bath [22]. Bats [3]. Battuta [510, 479]. Bau [1148].

Bauer [121]. Bauleute [366]. be [148, 663]. Beagle} [263]. Beardsley [776]. Beasts [1179]. Beaujouan [1084, 728]. Beck [56, 484]. Beckinsale [715]. Becquerel [872]. Bedeutung [450]. Bedini [809]. Beer [639, 103]. Beeson [1219]. Before [41, 33, 158, 953, 715, 542, 928, 524, 565]. Begegnung [32]. Beginning [175, 60]. Beginnings [648, 311]. Begründung [1348]. Behaviorists [1198]. behind [352]. beiden [450]. Being [548]. Beitrag [1348, 1359, 614]. Beitrage [397, 366, 32]. Beitragen [1040]. Beldomandis [1253]. Belgie [374]. Bell [72, 199]. Beloved [305]. Bends [636]. Benedetti [1343, 430]. Benfey [1077, 759]. Benison [1215]. Benjamin [648, 72, 800, 1309, 923, 1255]. Benoeming [374]. Bentley [51]. Berengario [137]. Berg [777]. Bergbau [366]. Bergbaus [366]. Bericht [847]. Beringer [548]. Berkeley [623, 669, 623, 978, 291]. Berkner [712]. Berlandier [1353]. Berliner [160, 334]. Berman [1080]. Bern [139, 104, 322]. Bernard [1279, 408, 123, 1021, 1192, 335, 1162, 344, 191, 304, 1191, 562, 1278, 1314, 159]. Bernhard [1148]. Bernoulli [198]. Berossos [1110, 1329, 942]. Berry [288]. Bert [54]. Berthelot [781]. Berzelius [1019]. bespozvonotschnie [373]. Better [683]. Betty [1355]. between [628, 426, 603]. Bewegungslehre [450]. Beyond [1164, 1228, 558]. Beziehungen [1359]. Beziehungenzwischen [32]. Bibby [141]. Bibliographia [1345, 796]. Bibliographical [812]. Bibliographie [1156, 1240, 1240, 760]. Bibliography [996, 1149, 1258, 1219, 338, 339, 658, 820, 934, 1096, 1224, 1364, 854, 106, 204, 378, 517, 1180, 799, 783, 1118, 1209, 1338]. bibliothèques [1154]. Biennial [191]. Bigness [434]. Bildung [1148, 195]. Billings [884, 884]. Bio [1156, 1240]. Bio-Bibliographie [1156, 1240]. Biographical [414, 533, 992, 903, 1255, 1178, 756, 342]. Biographien [716]. Biography [1217, 962, 816, 887]. Biologen [687, 552]. Biología [58]. Biological [1114, 199]. Biologie.Bd [1041]. Biologist [231, 813]. Biology [178, 564, 93, 475, 189, 718, 1299, 1117, 675]. Bird [43]. Birth [850, 123, 972, 838, 806, 483]. Biruni [603, 285]. bis [281, 369, 614, 1326]. Blackwell [808]. Blackwood [1313]. Blake [1180]. Blame [353]. Blaser [369]. Blending [460]. Bliakher [373]. Blood [615]. Board [1227]. Boas [840, 708, 412, 729, 29]. Boatfield [72]. Bobinger [1115]. Bochner [958]. Bodenheimer [231, 189]. Boder [856]. Bodies [8, 265, 749, 1199, 295]. Body [910]. Boerhaaviana [731]. Boerhaavii [731]. Bohemian [537]. Böhm [1339]. Bohr [1040, 707, 1129, 707]. Bohrs [1040]. Bologna [138, 1340, 461, 1066]. Bolzano [777]. Bomb [1218]. Bombelli [461, 742]. Bond [1000]. Bondaroy [370]. Bonkis [496]. Bonser [678]. Book [1360, 469, 640, 419, 647, 1056, 709, 467, 956, 336, 1092, 579, 1015, 1159, 1211, 547, 59, 331, 414, 472, 1220, 368, 485, 774, 1051, 1053, 34, 689, 780, 1216, 1317, 1077, 996, 15, 122, 340, 448, 688, 1266, 127, 804, 1074, 139, 583, 724, 959, 1007, 1115, 1145, 776, 957, 1093, 1239, 372, 722, 1279, 573, 628, 1088, 932, 230, 815, 767, 890, 1004, 964, 120, 329, 94, 157, 373, 1346, 631, 1308, 324, 92, 134, 297, 398, 445, 500, 589, 713, 615, 538, 1237, 20, 408, 714, 421, 447, 721, 1246, 367, 1177, 818, 995, 620]. Book [903, 1157, 1268, 1261, 571, 22, 483, 1079, 1315, 1349, 1356, 320, 644, 1148, 882, 1006, 687, 733, 52, 639, 737, 778, 958, 1008, 1355, 301, 418, 1054, 285, 323, 878, 1149, 1357, 1347, 648, 1263, 284, 679, 880, 229, 817, 930, 1114, 783, 801, 580, 72, 119, 155, 161, 359, 371, 729, 926, 415, 446, 514, 654, 1003, 1234, 1358, 300, 546, 1270, 337, 549, 585, 1164, 911, 543, 846, 420, 1019, 850, 1206, 30, 611, 809, 1188, 1267, 1348, 225, 406, 685, 717, 287, 622, 840, 907, 1353, 283, 568, 102, 616, 617, 482]. Book [910, 14, 997, 101, 413, 619, 123, 8, 196, 303, 332, 466, 1212, 708, 13, 885, 219, 1218, 1017, 574, 54, 141, 784, 972, 1009, 650, 711, 851, 881, 1154, 1306, 201, 281, 886, 1176, 1244, 293, 772, 1262, 296, 725, 626, 761, 781, 888, 1247, 712, 299, 773, 1082, 1337, 498, 674, 816, 922, 221, 757, 777, 64, 70, 1120, 1126, 1163, 1359, 1245, 764, 971, 677, 1269, 1052, 19, 1304, 906, 1086, 844, 53, 302, 536, 811, 1345, 564, 1142, 1190, 1073, 65, 798, 1278, 1089, 27, 474, 591, 594, 610, 132, 544, 680, 728]. Book [808, 1185, 218, 904, 33, 403, 652, 194, 1184, 1219, 503, 1021, 1055, 345, 769, 58, 97, 105, 140, 370, 682, 91, 100, 121, 162, 269, 272, 338, 339, 399, 504, 586, 800, 1235, 203, 1158, 1342, 763, 730, 1354, 473, 577, 923, 676, 1083, 642, 50, 192, 1210, 775, 875, 1023, 1165, 1186, 1311, 98, 292, 587, 588, 613, 810, 1072, 1264, 341, 970, 77, 333, 755, 509, 441, 961, 690, 643, 74, 274, 362, 374, 400, 516, 1060, 951, 289, 342, 410, 1128, 625, 1005, 1080, 1333, 552, 575, 718, 1191, 1192, 1078, 1012, 268, 135]. Book [165, 136, 646, 1046, 1123, 1182, 223, 1331, 476, 1140, 537, 407, 76, 731, 629, 683, 1151, 1187, 1310, 60, 1076, 1274, 93, 216, 471, 195, 924, 1161, 1238, 1058, 909, 960, 1214, 852, 158, 273, 570, 719, 1178, 325, 126, 328, 17, 452, 671, 1314, 202, 326, 397, 1146, 720, 1213, 1343, 1217, 1271, 758, 1049, 291, 802, 723, 753, 29, 103, 129, 298, 411, 487, 501, 541, 627, 759, 1116, 130, 1241, 797, 375, 813, 1312, 1010, 231, 481, 1307, 159, 1344, 1091, 1042, 1298, 857, 858, 766, 197, 294, 409, 681, 279, 805, 511, 1236]. Book [756, 199, 276, 765, 1045, 768, 884, 1041, 592, 1153, 1305, 278, 125, 1180, 475, 657, 450, 1265, 220, 21, 1119, 1242, 69, 200, 376, 486, 505, 762, 28, 1309, 99, 1047, 214, 732, 760, 848, 953, 1273, 224, 26, 908, 477, 1044, 1205, 1335, 515, 707, 1059, 1129, 966, 1020, 137, 138, 327, 847, 1340, 71, 104, 307, 551, 929, 417, 1189, 715, 369, 508, 542, 1143, 726, 933, 502, 540, 506, 754, 282, 782, 582, 510, 849, 1299, 96, 335, 401, 402, 641, 277, 438, 1118, 855, 1022, 624, 198, 584, 1350, 295, 673, 998]. Book [1316, 67, 439, 609, 876, 889, 952, 955, 999, 1050, 1127, 1147, 215, 361, 455, 507, 1276, 655, 360, 1016, 1209, 286, 363, 443, 1156, 1240, 1336, 735, 967, 969, 1075, 1048, 1144, 1297, 845, 548, 1272, 271, 566, 131, 545, 24, 222, 678, 1208, 1117, 1162, 1352, 470, 306, 653, 968, 31, 319, 512, 590, 1125, 156, 807, 217, 796, 1000, 973, 567, 710, 56, 734, 770, 877, 736, 228, 344, 854, 841, 931, 1155, 716, 1183, 1160, 95, 1243, 1300, 1334, 1338, 1090, 449, 618, 62, 124, 779, 191, 330, 75, 656]. Book [581, 1152, 227, 404, 623, 771, 925, 928, 1002, 1087, 1121, 1179, 343, 684, 812, 614, 963, 128, 853, 645, 806, 799, 366, 480, 883, 346, 1194, 1207, 1313, 1084, 16, 73, 565, 675, 1303, 1332, 814, 133, 280, 66, 651, 270, 442, 1043, 32, 160, 226, 334, 444, 1013, 1014, 1339, 1277, 23, 164, 193, 364, 1018, 1081, 1275, 1301, 499, 405, 843, 954, 479, 686, 1150, 1181, 1341, 1085, 593, 962, 887, 927, 57, 484, 612, 569, 578, 842, 1011, 803, 1122, 412, 513, 1124, 1351, 1302, 290, 63, 576, 795]. Book [994, 453, 468, 1113, 550, 288, 305, 322, 365, 454, 539, 672, 1001, 1141, 68, 189, 304, 377, 572, 1057, 1215, 488, 1193, 416, 440, 25, 478, 965, 856, 18, 61, 190, 275, 321, 621, 630, 879, 905, 451, 51, 55, 649, 727, 1054, 1273]. Books [1157, 953, 542, 1143]. Boole [589]. Boorse [960]. Borealis [1065]. Borelli [363]. Boring [421]. Born [1163]. Boscovich [483, 669, 1231, 987]. Bose [871]. Botanical [1176, 606, 905]. Botanist [1307]. Botanists [824]. Botany [886, 788, 19, 623]. Boucher [773]. Bouguer [89, 39]. Boveri [552]. Boyer [568, 1333]. Boyle [840, 338, 29, 228, 559, 990, 732]. Bradley [413]. Bradwardine [941]. Bradway [884]. Brahe [979, 1173, 939]. Brain [20, 1168, 930, 592, 16]. Braintree [325]. Brand [885, 847]. Brandl [1046]. Braunschweig [1292]. Breves [137]. Brian [620]. Bridge [855]. Bridgman [303]. Brief [708]. Briefe [1108, 1326, 1161, 671, 1090]. Briefwechsel [74, 929, 160, 334, 1055]. Brieven [1092, 481]. Brighter [34]. Brigitte [1359]. Brinitzer [289]. Britain [717, 318, 260, 543, 1005, 689]. British [721, 1356, 52, 580, 885, 590, 697, 799]. Brock [1089, 401]. Bronowski [269]. Brooke [1082, 883]. Brooklyn [855]. Brooks [94]. Brophy [466]. Brown [1239, 1309, 512, 1002, 569]. Browne [414, 868]. Bruce [269]. Bruins [880]. Bruno [622]. Brussels [753]. Buch [96]. Buffon [624, 207, 45, 1016]. Builders [126]. Building [201, 1337, 1183]. buldan [135]. Bulletin [1056, 1181]. Bullock [23]. Bullough [1266]. Bunge [13]. Bureau [1073, 341]. Buridan [1068]. Burke [1078]. Burkert [649]. Burmeister [1156, 1240]. Burning [207]. Burning-Mirrors [207]. Burroughs [764]. Bush [472, 1002]. Bush-Brown [1002]. Butlerov [487]. Bykov [298, 487]. Bylica [225]. Byrne [1272]. Byron [763, 763]. byzantinisches [772]. Byzantium [440].

C [804, 52, 801, 72, 846, 1353, 219, 725, 677, 564, 132, 680, 194, 1184, 1021, 345, 292, 1060, 1271, 1042, 766, 1153, 762, 1309, 760, 327, 782, 510, 624, 295, 889, 1208, 779, 1332, 1150, 569, 1011, 1124, 672, 95]. Cable [104]. Cadet [527]. cadrans [1074]. caeteris [369]. Cagnard [210, 431]. Cahn [733]. Cajori [983]. Calcareous [420]. Caldwell [342]. Calendar [637]. Calhoun [299]. California [426]. Calvert [1049]. Cambridge [549, 1075, 688, 30, 412]. Camera [742]. Camp [725, 418, 616, 725]. Campanella [287]. Campbell [19, 557, 1170]. Canada [957, 783, 1093]. Canadian [1239]. Canal [126]. Candole [61]. Canguilhem [121]. Cannon [560]. Cantimpre [464, 436]. capacity [430]. Cape [1355]. Capillaries [1104]. Captain [139]. Cardanus [1105]. Cardini [727]. Care [769, 683]. Carl [568, 201, 1337, 289, 1333, 1310, 1213, 1031]. Carletti [763]. Carlo [1015]. Carlos [1243]. Carlson [1003]. Carlyle [420]. Carmody [509]. Carnot [82, 210, 431]. Carol [1190]. Carozzi [654, 417]. Carpi [137]. Carpine [1265]. Carpine-Lancre [1265]. Carroll [1140, 889, 1141]. Carson [1311]. Cartesiana [796]. Cartographic [17]. cartographica [1085]. Cartography [717, 714]. Case [636, 1200, 209, 696]. Caspar [1345]. Catalogue [1157, 679, 376, 1314, 542, 1143, 565]. Catalogus [133]. Catastrophist [7]. Catesby [410]. Cathedral [530]. Cathedrals [546]. Catherine [725]. Causal [13]. Causality. [13]. causis [1264]. Cautionary [987]. Cavendish [288, 1283]. Cecil [214]. Cecily [973]. Celestial [225]. Cell [1044, 963]. Celsus [1017]. Centaur. [192]. Centenaire [1203]. Centenary [1203, 487]. Centre [1211]. Centrifugal [115]. Centuries [767, 373, 131, 645]. Century [472, 815, 1157, 1079, 30, 910, 763, 1005, 1214, 1343, 29, 1307, 740, 735, 1048, 590, 95, 656, 227, 1070, 4, 600, 1065, 697, 1036, 1292, 1228, 741, 182, 1102, 241, 1283, 426, 938, 147, 943, 1030, 492, 606, 703, 696, 824, 42, 350, 1177, 1356, 201, 1262, 1023, 537, 60, 681, 31, 1125, 931, 1141]. Ceremonioso [583]. Ces [365]. Ch [121, 334, 1013]. Chabrier [100]. Chain [1257]. Chain-Clock [1257]. Chaldean [1110, 1329, 942]. Chambers [33]. Chandelier [1327]. Chandler [94]. Change [672, 477]. Changed [1040]. Chaos [449]. Chappe [1052]. Chapters [135, 17, 1302]. Character [529]. Characteristics [916]. Characters [417]. Chardin [203]. Charles [628, 648, 54, 906, 91, 1342, 50, 1060, 342, 328, 1344, 1180, 475, 24, 736, 95, 684, 812, 66, 1301, 54, 108, 460, 639, 778, 1217]. Charleton [1012, 559, 1012]. Charlton [75]. Charltoniana [1012]. Chart [1356, 717]. Chasis [890]. Chasles [1203]. Checkland [848]. Chemical [129, 298, 487, 759, 592, 125, 1000, 428, 826, 662, 953]. Chemie [1348]. Chemiker [76]. Chemist [571, 682, 399]. Chemistry [1051, 776, 911, 491, 1247, 1120, 523, 173, 504, 386, 541, 696, 761, 218, 904, 29, 501, 1116, 515, 228, 1334, 129]. Chemists [1292]. Chemurgic [1284]. Chesneaux [121]. Chevreul [515]. chez [297]. Chickens [621]. Chief [1159]. Childs [1216]. China [136, 1123, 700, 560, 385, 332, 27, 407, 994]. Chinese [931, 637, 1064, 916]. Ch’ing [700, 1155, 385]. Chinoise [223]. Chirurgie [1267, 1081]. choisies [1212]. Cholera [736]. Chorley [715]. Chr [1339]. Chretienne [547]. Christian [685, 929, 321]. Christians [403]. Christine [846]. Christoph [289]. Christopher [688, 853]. Chronicles [763]. Chronology [784]. Chrysostomos [281]. chuan [931]. ciel [1341]. Ciencia [398]. Ciencias [58]. Cimento [732]. Cinquecento [1186]. circa [1134]. Circulation [615, 862]. Circumnavigation [763]. Citation [818]. Citizen [846]. Civil [299]. Civilisation [136, 1140, 994, 576]. Civilization [1210, 1141]. Civilizations [1122]. Clagett [268, 1341, 68, 965]. Clair [723]. Clairaut [527]. Clap [149]. Clarence [1177]. Clark [764, 1277, 817, 711, 402]. Clarke [209]. Classic [1155, 275]. Classical [293, 298, 444, 293]. Classics [564, 438, 564, 759]. Classification [1324, 514]. Claude [472, 1052, 1278, 1191, 1314, 1192, 1162, 304]. Claus [1041]. Clayton [623]. Clerk [425]. Cleugh [34]. Clifford [371]. Climatological [832]. Cline [1040]. Clock [1327, 1257]. Clocks [407, 849]. Clockwork [407]. Cnidian [194]. Co [1147]. Co-operation [1147]. Cochrane [652]. Codex [880, 743]. coelesti [1264]. Cohen [121, 123]. Cohesive [956]. Colbert [1335]. Colcestrensis [875]. colegios [728]. Coleman [572]. Colin [968]. Collander [886]. Colleagues [1129]. Collected [1009, 481, 1309, 1009, 888].

Collecting [253]. Collection [697, 250, 894, 487, 955, 412]. Collections [252, 482]. College [711, 376, 277, 1313, 149]. Colloque [286]. Colodny [566]. Colonial [1337, 410]. Color [747, 313]. coloribus [1241]. Colors [387]. colorum [1264]. Colours [732]. Combination [759]. Combining [10]. Combustion [411]. Comenius [71]. Comentario [390]. Comes [1276]. cometis [810]. Comets [292]. Coming [579]. Commemoration [861]. Commensurator [1291]. Comment [986]. Commentaires [770]. Commentaries [133]. commentariorum [133]. Commentarius [1011]. Commentary [1261, 390, 967, 1160, 808]. commenté [69]. Comments [671, 836, 751]. Communist [1246]. communs [1191]. Company [325]. Completas [1243]. Completes [408]. Complexionibus [435]. Compton [1194]. Comptons [1313]. Computer [16]. Concede [1037]. Concentric [698]. Concept [899, 844, 786, 1171, 1144, 593, 1001]. Conception [655]. Conceptions [424, 1228]. Concepts [303, 330, 42, 444]. Conceptual [1040]. concerning [1159, 683, 183]. Concise [293, 361]. Condit [201, 1337]. Conditions [672]. Conference [259, 169, 1254]. Conflict [375, 426, 299, 1049]. Congrès [1265]. Congress [1254, 793]. Conjecture [747]. Conjectures [674]. Connection [1106]. Connely [300]. Conolly [683, 1275]. Conon [1035]. Conquest [848]. Conscience [375]. Considerations [732]. Consistent [1295]. Constantinopolitanus [880]. constituendo [1264]. Constitution [592]. Construction [557, 644, 546, 1275]. Contacts [870]. Containing [438]. contemporaine [454]. Contemporary [394, 242, 1124]. Continued [831]. Contra [599]. Contra-Equivalence [599]. Contraction [864, 984]. contre [335]. Contribution [1266, 150, 84]. Contributions [301, 1172, 1056]. Control [1005]. Controversial [183]. Controversy [523, 741, 292, 1132, 377]. Conversation [852]. Conway [93, 675]. Coomaraswamy [1258]. Cooperative [954]. Cope [277]. Copeman [766]. Copenhagen [979]. Copernic [363, 284]. Copernican [709]. Copernicus [185, 599]. Copy [116]. coquillages [482]. Corn [630]. Cornelius [413, 923]. Corner [767]. Cornered [968]. Corpuscular [990]. Correction [309, 318, 358, 464, 608, 920, 1070]. Corrections [1112]. Correspondance [1161, 452]. Correspondence [117, 605, 356, 1355, 161, 48, 1292, 212, 266, 162, 187, 945, 11, 88, 163, 1160, 1326, 513, 893, 1124, 316, 153, 628]. Cosmas [547]. Cosmogony [351]. cosmographicum [810]. Cosmologiques [284]. Cosmology [1173, 328]. Cosmos [1194, 1228]. Costa [515]. Costabel [121, 372]. Costello [30]. Cotton [697]. Coulston [91]. Council [1093, 1284]. Count [1309, 569]. Couper [487]. courants [500]. Cours [223]. Course [564]. Cowen [768]. Cowley [337]. Cox [688]. Cradle [22]. Crafts [1122]. Cragg [681]. Cranefield [94]. Crawford [495]. Creation [1144, 841]. Creeping [1259]. crepusculis [1070, 943]. Crescini [1186]. Critical [414, 1003, 65, 797, 658, 820, 934, 1096, 1224, 1364, 796, 106, 204, 378, 517, 268]. Criticism [834, 978]. Critique [1344, 1242]. Crombie [132, 672]. Crone [416]. Crookes [863]. Crosland [1354, 501]. Crowther [1316, 1050, 290].

Crucial [1250, 411]. Crystals [1078]. Cultivated [61]. Cultural [712, 658, 820, 934, 1096, 1224, 1364, 106, 204, 378, 517]. Culture [956, 1177, 1206, 610, 131, 1303, 611, 1076]. Cultures [1049]. cum [810]. cura [807]. Curran [995]. Curriculum [30]. Curt [1054]. Customs [894]. Cutting [127]. Cuvier [572]. Cycle [748, 835]. Cyclotron [1216]. Cyril [141, 773, 279]. Cytochrome [1044]. Czerwinski [709].

D [1360, 1220, 957, 882, 778, 287, 102, 345, 292, 342, 158, 570, 1314, 28, 999, 734, 779, 75, 1152, 814, 280, 1043, 1081, 1341, 1001, 406]. Dagnelie [1045]. Dailey [995]. Daily [831]. Dakhel [478]. Dale [401]. d’Alembert [730]. Dales [1011]. Dalton [826]. Daltonian [208, 494, 826]. d’Andre [511]. Daniel [299, 1147, 548, 734, 806, 526, 1031, 1310, 486, 198]. Daniels [1127, 1276]. Danmark [474]. Dann [76]. Dannenfeldt [1307]. dans [331, 1154, 551, 453]. Danto [270]. d’Apotekers [1158]. d’après [589, 100]. Darlington [345]. Darnton [1234]. Darwin [52, 778, 54, 764, 33, 50, 812, 51, 605, 263, 529, 740, 684, 460, 694, 1106, 639, 1350, 54, 53, 345, 1312, 199, 321]. Darwinian [53]. Darwiniana [686]. Darwinism [686, 1276]. Darwinismus [190]. Data [818]. Date [838]. Daumas [121, 155, 576]. d’Avicenne [69]. David [340, 1355, 1218, 886, 1311, 768, 1044, 439, 999, 271, 927, 305, 856, 567, 846, 883]. Davids [369]. Davis [715]. Davy [121, 895, 987, 571]. Dean [1079]. Death [219, 1069]. Debate [7, 1282]. Debates [662, 1089]. Deborah [1277]. débuts [1267]. Decans [367]. Decay [224]. December [35, 78, 1289]. Dechend [627]. Deductive [939]. Deeming [1355]. Defender [1066]. Defends [893]. Defense [599]. Definitive [1254]. Defremery [510]. del [583, 445, 1017, 1186, 390, 732]. Delaunay [121]. D’Elia [332]. dell [878]. della [1344, 502]. dell’Ottocento [1357]. Delorme [409]. Delta [899]. Dember [720]. Democritus [442]. Demography [1135]. Demonic [287]. Demonstratio [430]. Demonstration [430]. Denker [77]. Denmark [1317]. Dennis [854]. Denomy [1341]. Department [1284]. deque [1264]. Derek [320, 407]. Derivation [1100]. Derry [476]. Descartes [551, 796, 866, 1322, 1257, 297, 101, 360, 969]. Described [637]. Describing [526]. Descriptio [1013]. Description [252, 683]. Desde [445]. Desiderio [445, 58]. Desmond [1350]. Destombes [1085]. Detection [1005]. Determination [631]. d’Etienne [733]. Deutsch [626, 32]. Deutsch-Russische [32]. Deutsche [1359]. deutschen [32]. deutscher [76]. Development [59, 1317, 1266, 94, 301, 761, 798, 27, 263, 1040, 1042, 214, 715, 439, 1147, 545, 963, 578, 1122, 313, 1057, 230]. Developments [1138]. d’Histoire [1265, 624, 1181, 1235]. Dialexis [1264]. Dialogue [873, 1021, 224, 1159]. Dialogues [1212]. Diamond [895]. Diaries [1355]. Dibner [104, 322]. Dicks [280]. Dictionary [1262, 1047]. Dictionnaire [724, 66]. Did [1109, 748, 1035, 1037]. Didactica [71]. Diesel [301]. Difform [1253]. difformis [1199]. Diffusion [448]. Digestion [868]. Dijksterhuis [982, 538, 834]. Dilemma [1103, 586]. Dillon [887]. Dimension [600]. Din [406]. Dinawari [26]. Dindimi [1185]. Dinosaurs [1335]. Dioscorides [130]. Direct [914]. direction [121]. Dirigibles [1009]. Discarded [680]. disciple [1242]. Discourse [224]. Discoverer [105]. Discoverers [615]. Discoveries [588, 927, 1051]. Discovery [780, 324, 1200, 473, 115, 428, 939, 842, 322, 672, 15, 17, 758, 1237]. Discussion [243, 258, 255, 249]. Disease [1274, 200]. Diseases [766]. d’Isis [383]. Disjunctive [956]. Diskussion [1306]. Dispersion [690]. Dissertations [335]. Distance [350]. Distillation [601]. Distinction [603]. Diverse [196]. Division [710]. Dlicht [1158]. Dmitrii [485]. Doby [615]. Doctor [1188, 291]. Doctors [885, 723]. Doctrines [284]. Documentary [735, 1301]. Documents [233, 1017, 764, 167, 348, 79, 461, 742]. Dokumenten [1148]. domaines [453]. Domingo [1199]. Don [583]. Donald [688, 801, 764, 341, 889, 655, 856]. Dora [1245]. Dorothea [64]. Dorothy [291, 1273, 623, 440]. Double [1220]. Doublet [1134]. Doubts [208, 494, 826]. Douglas [472, 1174]. Down [1079]. Dr [463, 339, 629, 548, 73]. Drabkin [926, 790, 1343]. Drachmann [581]. Drake [292, 1343, 486, 734, 1159, 926]. Drazen [1267]. Dream [1072, 901, 811]. Drebbel [413]. Drieux [1084]. Driven [1257]. Drug [1005]. Drugs [62]. Duality [1295]. Duane [102]. Duckworth [327]. Dugas [121]. Duhem [284, 1170]. Dujarric [100]. Dulieu [121]. DuMond [631]. d’une [511]. Dunn [715]. Dupree [686]. Dureau [1084]. Dureau-Lapeysonnie [1084]. during [263, 700, 75, 385]. Durling [1157]. Durme [452]. Dutch [481, 86]. Duveen [854]. Dyck [226]. Dye [103]. Dyeing [4]. dynamique [372].

E-tu [931]. E. [1143]. Earlier [677]. Earliest [476, 1140, 5, 1289, 214, 918]. Early [235, 1093, 367, 862, 648, 30, 1067, 1082, 560, 132, 1198, 1184, 182, 186, 202, 84, 28, 508, 845, 700, 1002, 939, 193, 385, 809, 505, 645]. Earth [433, 262, 463, 862, 598, 1101, 751, 1347, 136]. East [225, 440, 1083]. Eastern [1122]. Ebenezer [1321, 775]. Ebers [183]. ecentricorum [1169]. Echinus [699]. Ecology [694]. Economic [815, 1147, 803]. Economics [176]. écrits [100]. ed [480]. Edelstein [1271, 898]. Edge [91]. Edinburgh [495]. Edited [1160, 671]. Editing [240]. Editio [810]. Editions [1067, 1105]. Editorial [555]. Edmund [420, 808, 76, 1272, 623, 963]. Edson [1164]. Eduard [1051, 1214, 1193, 982]. Education [932, 964, 113, 612]. Educator [141]. Edward [709, 776, 514, 1206, 1262, 1304, 852, 966, 726, 215, 271, 1179, 343, 344]. Edwards [988]. Edwin [421, 1042, 1335]. Eecke [47]. eerste [374]. eeuwse [329]. Egypt [1083, 881]. Egyptian [194, 367, 849]. égyptienne [128]. Ehrenfried [293]. Eichholz [882]. eigenen [627]. Eighteenth [1177, 910, 681, 1048, 31, 1125, 227, 4, 1065, 741, 1283, 147]. Eighteenth-Century [910, 1048, 227, 4, 1065, 741, 1283, 147]. Eighth [517]. Eighty [658, 106, 204, 378, 517]. Eighty-Eighth [517]. Eighty-Fifth [106]. Eighty-Ninth [658]. Eighty-Seventh [378]. Eighty-Sixth [204]. Eileen [995]. Einführung [1269]. Einshtein [594]. Einstein [594, 671, 1250].

Einzeldarstellungen. [614]. Eis [1305]. Eiseley [586]. Eisenstein [924]. eiusdem [1264]. Ekirch [754]. Elastic [295]. Eleatic [582]. Electrical [617, 1288]. Electrodynamics [350, 1172]. Electroluminescence [1200]. Electromagnetic [914, 524, 746]. Electromagnetic-Wave [524]. Electromagnetism [322]. Electron [631]. Elektrizitat [1193]. Elektrizitatslehre [488]. Elementary [386]. elementis [697]. Elements [428, 1237, 10]. Eleventh [1070, 943]. Eleventh-Century [1070, 943]. Elfriede [925]. Elias [1124]. Elihu [305]. Elizabeth [1346]. Elizabethan [28]. Ella [1358]. Ellegard [52]. Ellery [1202, 962]. Elliptical [1100]. Elmer [628]. Elof [1003]. Eloge [152, 429, 830, 355, 982, 389, 109, 1174, 47, 790, 110, 898, 829, 1325]. Eloges [869, 46, 108]. Elskamp [1074]. Elting [1006]. Elvira [1237]. Embriologii [373]. Embryology [373, 1117, 157, 800, 1205]. Emergence [1140, 103, 451]. Emergent [976]. Emil [122, 782]. Emmet [342, 486]. Emmons [1321]. Emperor [1155]. Empire [970, 75]. Empiricism [1283]. emptrismo [480]. Encyclopedia [388, 756]. End [1177, 1234, 537]. Endeavour [1310, 1031]. Endocrinologist [1259]. Ends [254]. Energy [689, 657, 1232, 1057, 1001]. Engelbreth [1317]. Engelbreth-Holm [1317]. Engine [301, 82, 210, 431]. Engineer [299, 742, 1049, 773]. Engineer-Architect [742]. Engineering [1009, 994, 326, 1075]. Engineers [82, 616]. Engines [301]. England [959, 903, 1349, 28, 848, 277, 678, 559, 1227, 325, 1153]. English [932, 1262, 1047, 1048, 853, 927, 86, 315, 917, 824, 1008]. Enigma [1199]. Enlightenment [1234, 652, 1065, 159]. Enquiry [815]. entomologiques [482]. Entwichlung [488]. Entwicklung [122]. Environment [1270]. epiciclorum [1169]. Epicurean [559, 693]. Epicuro [1012]. Epicuro-Gassendo-Charltoniana [1012]. Epicurus [1012]. Epicyclical [40]. Episode [740]. Epistle [698, 789]. Epitome [1185]. époque [128]. Epoques [624]. Equatorium [406]. Equivalence [599, 1232]. Era [293]. Erasmus [1350, 360]. Erde [369]. Erection [546]. Erforscher [1193]. Ergebnisse [614]. Eric [229, 652]. Erich [156]. Erik [1019]. Eriksson [138]. Erna [857, 858]. Ernan [1211]. Ernest [467, 1216, 972, 1154, 498, 1151, 306, 803]. Ernst [574, 1021, 1055, 1005, 1326, 416, 1108, 1326, 814]. Erratum [89]. Errors [650]. Erschliesser [55]. Erste [369]. Erwin [1088, 657, 795, 973, 1165]. Eskew [22]. Esmond [647]. Espanola [398]. Especially [1292]. Espiritu [1017]. Esploratori [1357]. Essai [128]. Essay [420, 840, 91, 87, 656, 514, 442]. Essayes [732]. Essays [997, 1235, 192, 1303, 686, 707, 1004, 421, 200, 955, 455, 1117, 856]. Essential [1350]. est [148]. Esther [995]. Estudios [398]. Eternal [62]. Ethiopie [1182]. Ethnology [1073, 894]. Ethology [1198]. Etienne [1278]. Etiology [653]. Études [1235]. Euclidean [352]. Eugen [301, 1009]. Eugene [818, 515, 1147, 1338]. Eugeniusz [709]. Euler [929, 32, 1013, 1014]. Euleriana [1013, 1014]. Eulers [160, 334]. Eupalinos [1107]. Eupalinus [664]. Europas [765]. European [150, 656]. Europeene [1045]. Eva [1054, 355]. Evans [1355, 119]. Evenings [662]. Everett [448, 798]. Evidence [1184, 1030]. Evolution [52, 219, 54, 646, 5, 1289, 1205, 675, 572, 639, 574, 403, 93]. Evolutionism [740]. Ewald [55]. Ewan [573, 573]. Ewers [1353]. Exact [537]. Exactas [58]. Excitement [955]. Exercitationes [335]. Exile [1239]. Expedition [764, 764, 415, 764]. Experiment [636, 1250]. Experimental [800, 1162, 1030, 914]. Experimente [1348]. Experiments [3, 362, 411, 749, 732, 732]. Explanation [640]. Explanations [1065, 862]. Exploitation [516]. Exploration [516, 400, 971, 970, 762]. Explorer [970, 17, 962]. Exponent [2]. Expressed [954]. Extant [918]. External [417]. Extraction [601, 478]. Extraordinary [569]. Eye [263].

F [1015, 1266, 94, 500, 20, 447, 580, 1358, 287, 1009, 888, 904, 338, 410, 60, 758, 231, 847, 361, 1272, 1125, 567, 799, 160, 189]. F.R.S [483]. Fabrica [847]. Fabrick [1012]. Fabry [847]. Fachliteratur [1305]. Fact [855]. Fahrenheit [704, 667]. fait [365]. Fall [873]. Falling [749, 1199]. Falsely [1291]. Far [516]. Faraday [816, 151, 947, 113, 779]. Farber [1051, 1214]. Farm [1284]. Farquhar [109]. Fascination [955]. Fate [18]. Father [1216, 228, 149]. Fayet [140]. February [166, 790, 1352]. Fehl [1206]. Feigl [997]. Felix [192, 222]. Fellow [338]. Ferenc [1334]. Ferguson [1338]. Fernandez [1017, 97]. Ferrier [1168]. Festschrift [1305]. Festschriften [440]. Feuer [757]. Feverabend [997]. fi [967, 918]. Ficino [287]. Field [961, 342, 486, 1335]. Fielding [1056]. Fifteenth [606]. Fifteenth-Century [606]. Fifth [106]. Fifty [19]. Figure [1347]. Figures [980]. Figurovskii [485]. Filliozat [121]. fin [1084]. Finaly [719]. Finch [326]. Fineman [72]. fingo [315]. Finitude [789]. Finland [886]. Finlay [1243]. Finzi [967]. Fire [272, 1290]. First [992, 86, 315, 585, 1337, 1214, 411, 934, 138, 870, 352, 893, 290, 531, 1254, 1202]. Fisherman [561]. fisica [445, 502]. Fisk [291]. Fi’t [25]. Fi’t-Tibb [25]. Fitch [1183]. Fitchen [546]. FitzGerald [864, 984]. Five [301]. fizika [499]. Fleming [105, 655]. Flemming [1317]. Flesh [1076]. Flexible [295]. Flight [1009, 577]. Floating [8]. Flora [625]. Florence [530]. Florentine [763]. Florkin [419]. Flow [636]. Flowers [625]. Flugel [801]. Fluid [636, 371]. fluxu [917]. Flying [1009]. Folie [1336]. Folklore [678]. Folta [537]. Fontenelle [159, 409]. Footed [1179]. Forbes [538, 676, 202, 1010, 63]. Force [1077, 983, 115, 712]. Forces [1183, 956]. Foreign [1147]. Forel [1161]. Forerunners [51]. Foresight [469]. Forgotten [744]. Forman [1128]. Formation [1266]. forme [335]. Forms [1332]. Formulas [1077]. Forschers [156]. Fortleben [1342]. Fossilien [116]. fossilium [70]. Foster [1043]. Fothergill [403]. Fougeroux [370]. Found [1126]. Foundations [498, 951, 1208, 1321, 800]. Founded [1012]. Founders [358, 333]. Founding [149]. Four [747, 1179, 709]. Four-Color [747]. Four-Footed [1179]. Fourcroy [682]. Fourth [600, 1364]. Fragments [280, 226]. Frammenti [727]. Francaise [331, 1181, 370, 551, 140]. France [75, 1234, 1154, 1300].

Frances [622]. Francesco [878, 1357, 763, 1241]. Francis [1145, 52, 509, 629, 877, 152, 1200, 586, 290]. Franciscan [1017]. Francisco [1162]. Franck [99]. Francois [850]. Frank [414, 884, 550]. Franklin [648, 72, 775, 402, 1288, 211]. Franz [1342, 924, 449]. Fraser [954]. Fray [1017]. Fred [737]. Frederic [589]. Frederick [906, 376]. Free [873, 294]. Freeman [812]. Frege [1020]. Freiberg [134]. Freiesleben [685]. French [1354, 159, 877, 869, 1203, 562, 556, 294]. Freud [1021, 1055, 306, 653, 1238, 306, 964]. Frick [410]. Friedr. [1292]. Friedrich [301, 64, 74, 198, 1090, 226, 405]. Friend [355, 775]. Friends [1129, 683]. Frings [281]. Fritz [120, 296, 552, 795]. Front [1, 37, 81, 111, 145, 180, 181, 206, 309, 236, 261, 310, 349, 381, 423, 457, 490, 520, 554, 596, 633, 661, 692, 739, 819, 823, 860, 892, 913, 937, 975, 1025, 1062, 1098, 1099, 1131, 1167, 1196, 1223, 1226, 1249, 1281, 1319, 1363]. Frontiers [566]. Frueh [811]. Fullerton [628]. Fulton [338, 109]. Function [172, 507]. Functions [868, 1168, 20]. Furon [121]. Further [986, 494]. Fusul [918].

G [774, 959, 589, 421, 1268, 1315, 1356, 446, 543, 1176, 773, 844, 65, 594, 403, 121, 896, 985, 1078, 731, 1274, 1038, 802, 298, 487, 376, 848, 137, 847, 417, 946, 540, 1316, 952, 999, 1050, 566, 470, 217, 124, 75, 581, 1303, 160, 1013, 1014, 23, 290, 377]. G. [430, 871]. Gabir [96]. Gachot [1052]. Gage [229]. Gaius [1069]. Galapagos [1293]. Galaxies [737]. Galba [697]. Galen [650, 805, 327, 148]. Galilee [948, 1242, 1212]. Galilee. [1211]. Galilei [1159, 926, 292]. Galileo [1159, 926, 292, 755, 852, 433, 665, 1015, 1211, 1201, 262, 463, 459, 83, 8, 264, 873, 749, 870, 351, 792, 466, 1343, 1015, 332]. Gall [1168]. Gallagher [763]. Gambling [567]. Games [567]. Gamow [1040]. Garboe [474]. Garden [1121]. Garfield [818]. Garnett [22]. Garrett [18]. Garrison [1056]. Garzoni [1066]. Gases [425]. Gasking [1346]. Gassendi [294, 335, 459, 634, 825, 480]. Gassendo [1012]. Gaub [910]. Gaudin [1229]. Gavin [639]. Gear [127]. Gear-Cutting [127]. Gebhardt [765]. Gedanken [1348]. Gedenkschrift [57]. Gegenstand [1306]. Geistes [369]. Geistesgeschichte [1348, 1040]. Geisteswelt [1342]. Gene [1003]. General [52, 1267, 273, 475, 271]. generale [576, 454, 760, 121]. Generalization [6, 827]. Generalizations [938]. Generation [1346, 551]. Genesis [695]. Genetics [1054, 275, 879]. Genetik [614]. Genevieve [783, 1219]. Genius [1216]. Gent [374]. Geoffrey [339, 471]. Geoffroy [733]. Geographer [1133]. Geographical [899, 17, 280]. Geography [609, 392]. Geologic [1032]. Geological [1315, 182, 426]. Geologie [323, 366]. Geologiens [474]. Geology [1321]. geometriae [1185]. Geometric [354, 1042]. géométrique [66]. Geometry [352, 1341, 217]. Geomorphology [715]. Geophysical [248]. Georg [218, 76, 276, 1156, 1240]. George [139, 767, 1149, 711, 289, 410, 629, 720, 1040, 375, 21, 1127, 1276, 131, 962, 1302, 521, 1202, 813, 110, 1092]. Georges [572]. Georgius [70]. Gerald [956, 14, 681, 1018]. Gerard [105]. Gerate [1115]. Gerhard [302, 397, 1305, 443]. Germaine [770]. German [585, 1218, 103, 973, 1292, 1326]. Germany [1218]. Gershenson [806].

Gertrude [53]. Gesammelte [98, 587, 810]. Gesamtwerk [369]. Geschichte [397, 1041, 449, 614, 366, 323, 1269, 276]. Geschiedenis [374]. Gesprache [57]. Gessendus [1012]. Gestalt [720]. Ghiyath [406]. Gianfrancesco [1344]. Gibb [510, 479]. Gibbs [1125]. Gibridami [1053]. Gift [114]. Gifte [96]. Gilbert [330, 102]. Gilberti [875]. Gille [770]. Gilles [453]. Gillis [374]. Gillispie [91]. Gillot [589]. Gilman [876]. Gilpin [1300]. Gingerich [570]. Giordano [622]. Giorgio [544]. Giovanni [1066]. Glacken [1177]. Glass [51]. Gleanings [8, 264]. Glenn [1005, 908]. Gliozzi [502]. Globe [225]. Gloom [59]. Glory [59]. Glossary [606]. Gnomon [530]. Gnosis [584]. Gods [567]. Goetzmann [970]. Goichon [69]. Gold [743]. Goldammer [369]. Goldbach [929]. Goldbeck [301]. Goldfarb [1355]. Goldring [890]. Goldschmidt [307, 307]. Golino [1015]. Gonzalo [97]. Goodfield [842, 539, 377]. Gorgias [582]. Goshen [1301]. Gothic [546]. Gottfried [687]. Gotthilf [1090]. Gouhier [101]. Gout [766]. Government [885, 182, 1227, 657, 1275, 1315]. Gowing [689]. Graduate [1033]. Graham [1246]. Grammar [905]. grammatica [1185]. Grammatikos [1339]. Grande [1162]. grands [500]. Granit [1217]. Grant [966]. Graphic [311]. Grassi [292]. Gray [686]. Great [773, 407, 1238, 1002, 362]. Greco [881]. Greco-Roman [881]. Grecs [66]. Greece [440, 726]. Greek [881, 677, 272, 328, 130, 508, 217, 581, 899, 392, 1304]. Green [1002]. Greenberg [1147, 806]. Greene [219]. Greenwood [1304, 271]. Gregor [1053, 796, 614]. Gregory [911, 480]. Grene [1299, 618]. Grew [2, 928]. Griechischen [281]. Grimaldi [1241]. Grimsley [730]. Grimwood [1060]. Grinding [127]. Grmek [1267, 1314, 1081]. Grob [1018]. Groote [1336]. grossen [552]. Grosser [473]. Grosseteste [1011, 184, 528, 917, 602]. Grotjahn [924]. Ground [1067]. Groups [710, 657]. Grover [997]. Growth [175, 674, 604, 575, 1312, 1197, 377]. Grubbe [782]. Gruber [375]. Grunbaum [641]. Grundlagen [122, 614, 364]. Grundlicher [847]. Grundzüge [1269]. Gudehus [1359]. Guerlac [411]. Guidance [648]. Guide [1033, 342, 505, 796]. Guides [133]. Guido [1343]. Guiducci [292]. Guilielmi [875]. Guiyath [406]. Gundolf [1305]. Gunpowder [272]. Gunther [130]. Gustav [301]. Gustave [128]. Gustavus [1323]. Guterman [1358]. Guthrie [677]. Guy [728, 1084]. Guyénot [121]. Gweneth [336]. Gyula [1334].

H [1056, 776, 1088, 92, 818, 778, 72, 161, 717, 102, 482, 141, 498, 906, 1089, 162, 800, 970, 362, 1161, 1058, 328, 1307, 1153, 1119, 744, 762, 1335, 327, 782, 582, 952, 1127, 215, 455, 360, 545, 590, 1108, 513, 1124, 1193, 879]. H. [896, 985, 1038]. Haak [585]. Haase [1213]. Habakkuk [590]. Haber [52]. Hagecius [1264]. Hahn [1091]. Hakluyt [927]. Hale [1202, 962]. Hali [412]. Halieutica [506]. Hall [1355, 840, 708, 95, 708, 755, 999, 412]. Hamarneh [908]. Hamby [847]. Hamilton [510, 479]. Handbuch [296]. Handlist [812]. Hanes [856]. hange [584]. Hanks [1016]. Hanno [56, 484]. Hanoverian [959]. Hans [1348, 685, 77, 1305, 716, 614, 1122]. Hanson [593]. Hansruedi [1188]. Happy [391]. Hardaway [1267]. Hardenberg [226, 1090]. Hardie [968]. Hardin [18]. Harig [397, 443]. Hariot [1349]. Harkins [650]. Harless [1360]. Harlow [922, 346]. Harms [1151, 716]. Harold [995, 617, 906, 333, 1000]. Harre [271]. Harris [1325, 413, 1047]. Harrison [1266, 1004, 221, 1126]. Harry [712, 227]. Hart [1000]. Hartley [333, 325]. Hartmann [77]. Harvard [1311]. Harvey [336, 1114, 1104, 615, 1113]. Hasan [918]. hasta [445]. Hastings [973]. Hauksbee [1200]. haut [1154]. Hawi [25]. Hayck [1264]. Hayes [630]. Haythams [651]. Hayy [69]. Hayyan [96, 96]. Health [885, 768, 21]. Hearnshaw [721]. Heat [798, 1042, 1309, 377, 798]. Heavenly [407]. Heavens [136, 644]. Heckscher [73]. Hedgehog [699]. Hedges [803]. Hedley [472]. Hegel [269]. Heilbron [1128]. Heinrich [76, 361, 1090, 488]. Heinz [369, 1156, 1240]. Heissen [847]. Hele [1350]. Helen [72, 922, 962]. Helene [331, 72]. Helix [1220]. Hellenistic [131]. Heller [814]. Helmholtz [1172]. Helmont [787]. Helmut [323]. Hennemann [302]. Henri [1074, 1145, 101, 1212, 1052, 940]. Henrietta [713]. Henry [1237, 466, 1247, 1219, 960, 508, 222, 710, 395, 1169, 1283, 549, 288]. Henslow [1312]. Heraclitus [283]. Herbal [130]. Herbaria [1087]. Herber [628]. Herbert [1216, 890, 119, 997, 772, 763, 630]. Herbes [1087]. Heritage [577, 709]. Herivel [909]. Herman [1009, 906, 282, 451]. Hermann [281, 1090, 1172]. Hermetic [622]. Heroic [418]. Herrick [532]. Herschel [644, 1355, 1103, 643, 1355]. Hertha [627]. Hertz [524, 746]. Heseler [138]. Hesse [1169]. Hideomi [578]. Hiebert [657]. Higgins [399, 399]. High [1057]. High-Energy [1057]. Hijab [478]. Hilaire [733]. Hilda [1021]. Hilden [847]. Hilken [1075]. Hill [688, 853]. Himmel [369]. Himmelfarb [53]. Himmelsphotographie [685]. Hindi [918]. Hindle [1082, 883]. Hindu [508, 843]. Hinman [337]. Hinrichs [1323]. Hipparchus [280]. Hippocrates [807]. hippocratique [404, 24]. Hisab [918]. Histoire [284, 1016, 447, 155, 446, 121, 576, 454]. Historia [398, 296, 445]. Historian [355, 245, 239, 247, 550]. Historians [355, 869]. Historical [1004, 94, 1114, 784, 797, 542, 1143, 1183, 565, 1320, 1303, 1301, 1106, 1124, 1302, 672, 501, 578]. Historie [474]. Historiography [536, 1138, 1080]. historique [66]. Historischen [364]. History [647, 1056, 395, 636, 804, 1093, 1239, 253, 425, 767, 157, 373, 538, 20, 721, 818, 984, 620, 22, 921, 648, 783, 119, 1270, 611, 1267, 1033, 717, 708, 1218, 711, 886, 293, 491, 64, 764, 677, 844, 65, 610, 904, 503, 769, 97, 91, 272, 504, 577, 642, 192, 613, 341, 242, 400, 1060, 793, 1128, 1333, 718, 268, 476, 60, 273, 719, 202, 1146, 758, 177, 298, 541, 1298, 766, 279, 278, 475, 657, 21, 1119, 384, 658, 820, 934, 1096, 1044, 1224, 1364, 715, 508, 1287, 1118, 1022, 439, 361, 360, 1209, 131]. History [222, 678, 1117, 174, 106, 95, 1338, 771, 928, 1043, 1018, 1301, 696, 795, 189, 304, 572, 1215, 169, 204, 378, 237, 517, 856, 879, 905, 127, 1145, 1003, 1073, 345, 768, 277, 952, 735, 656, 803, 1317, 127, 230, 92, 714, 543, 1023, 537, 1334, 1179]. Hodgen [645]. Hodges [782]. Hoeppli [193]. Hoffmann [334, 160].

Hofmann [293]. Hohenheim [369]. Holder [323]. Holland [771]. Holly [1194]. Holm [1317]. Holmstedt [802]. Holmyard [95]. Holton [956, 14]. Holy [1137]. Homage [1015]. Homer [890, 890]. Homme [197]. Homocentric [1169]. Honor [997, 1303]. Honourable [338]. Hooke [1037, 828, 915, 387, 339]. Hope [59]. Hopkins [105]. Horatio [292]. Horine [342, 486]. Horizontal [89, 39]. Horned [989]. Horsky [537, 1264]. Hortulus [1121]. Hosack [846]. Hoskin [644]. Hospital [1088, 1018]. Hospitals [1275]. Hossein [1210]. Host [792]. Houghton [995]. House [1313]. Houzeau [760]. Hovey [299]. Howard [344]. Howell [472]. Hoyle [737]. hsing [931]. Huang [1155]. Huard [1267, 223, 1081]. Hub [688]. Huber [198]. Hugh [1349]. Hughes [1187]. Hugonis [1185]. Hull [92]. Humaine [1084]. Human [923, 707, 722]. Humanist [141, 1156]. Humanities [954]. Humboldt [57, 484, 55, 56]. Humboldts [57]. Humphrey [413, 1126]. Humphry [571, 895]. Hundred [709, 801]. Hung [1123, 1294]. Hungarian [719]. Hunger [772]. Hunter [573, 629, 686]. Huntley [414]. Hutten [498]. Hutton [1101]. Huxley [778, 141, 217, 612]. Hybrid [630]. Hydrogéologie [598]. Hydrogeology [654]. Hydrography [1356]. Hyperaspistes [810]. Hypnotism [342]. Hypotheses [315]. Hypothesis [1012, 1229].

Ia [373]. Ibanez [222, 192]. Ibn [1327, 26, 510, 96, 479, 25, 527, 69, 390, 918, 6, 827, 843, 354, 865, 1261, 651]. Ibrahim [918]. Iconographia [731]. Iconological [1263, 73]. Idea [1312, 159, 220, 928]. Ideas [1114, 214, 1231, 1048, 91, 192, 507, 304, 1239]. Idee [449]. Identification [150]. Identity [792]. Iena [276]. Ier [1181]. Ifram [285]. ig [719, 499]. Ignaz [858]. Ihde [904, 761]. ihre [1269]. ihrer [323]. II [1009, 370, 121, 162, 369, 508, 1143, 1240, 1014, 484, 994, 1141, 621, 832, 1136, 504, 479]. III [1009, 369, 454, 416, 541, 1243]. Ill [1018]. Illinois [1149]. Illness [1302]. Illumination [1227]. Illustrated [1270, 377]. Illustrating [1112, 1034]. Illustrations [235, 851]. Ilza [1155]. im [1306, 626, 302, 1342, 857, 443, 32, 160, 334]. Image [680]. Imagery [1048]. Images [1086]. Immunization [1119]. Impact [219, 712, 657]. Implements [1152]. imprimeur [511]. Imprints [505]. Incipits [679]. incognitae [1264]. inconnus [365]. Index [995, 440, 1041]. India [427]. Indian [1112, 1034]. Indians [1353]. Indications [683]. Indicopleustes [547]. Indies [97]. Indiscerpible} [988]. Individualitat [1040]. Industrial [1140, 848]. Industry [1360, 103, 125, 301]. Inertia [597, 825]. Infections [193]. Inference [1208]. Infinite [1228, 59]. Influence [964, 545, 344, 694, 1122]. Influences [658, 820, 934, 1096, 1224, 1364, 106, 204, 378, 517]. Informal [613]. Information [143, 232, 379, 518, 659, 821, 935, 1095, 1222, 1362]. Inheritance [460, 963]. Inner [1093]. Innovations [448]. Inquiry [498, 683]. Insane [683, 683, 1275]. Insanity [683]. Insects [1179]. Instance [939]. Institute [268, 1176]. Institutio [805]. Institution [683, 376]. Institutions [1266]. Instruments [1145, 1008, 809]. Intellectual [757, 269, 656, 696, 672, 853]. intellectuelle [453]. Interesting [662]. Interior [734]. Internal [1155]. International [1211, 1254, 793, 1265, 286]. internationale [1359]. Interpretation [235, 1184, 186, 1103, 455, 183]. Introduction [620, 1267, 680, 1165, 475, 137, 215, 271, 92, 189]. Introductory [135]. inusitatae [1264]. Invention [648, 871, 672, 1239, 418]. Inventioni [196]. Inventions [590]. Inventor [23]. Inventory [1128]. Invertebrate [373]. Investigations [1346]. Invisible [164]. Iohannes [702]. Ippocrate [807]. Ira [512]. Iran [427]. Irena [1269]. iride [1241]. Iron [22, 27, 815]. Ironworks [325, 325]. Irregular [980]. Irrweg [190]. Irving [818, 1059, 1218]. Isaac [161, 162, 718, 756, 412, 513, 1351, 550]. Isaag [513]. Isagogae [137]. ISIS [382]. Islam [1342, 273, 1210]. Islamic [637]. Islands [324]. Isler [1188]. Isles [580]. Isochrony [1201, 915]. Isolation [631]. Israel [231, 790]. Istoriia [373, 591, 216]. Istoriya [298]. Istvan [719]. Italian [493]. Italiano [1357]. Italy [72, 1343]. Itard [121]. IV [369, 697]. Ivanovich [485, 352]. Ivory [307].

J [1077, 122, 957, 1239, 329, 1308, 538, 1177, 818, 1157, 1079, 320, 1355, 1263, 229, 801, 72, 446, 654, 1358, 420, 1019, 910, 303, 761, 906, 65, 808, 904, 194, 1184, 121, 269, 272, 399, 504, 676, 775, 896, 1023, 985, 509, 374, 646, 537, 407, 60, 202, 1038, 541, 294, 592, 1305, 1119, 200, 214, 760, 224, 847, 715, 506, 582, 1316, 955, 1050, 1075, 548, 470, 590, 1000, 95, 1243, 191, 771, 565, 675, 1013, 954, 1341, 290, 63, 288, 377, 321]. J. [425]. Jac. [1019]. Jackson [375, 764, 1127]. Jacob [375]. Jacopo [137]. Jacqueline [1265]. Jacques [551, 624, 655]. Jaffe [344]. Jahn [548]. Jahren [74, 1326]. Jahrhundert [1306, 626, 302, 857, 443]. Jahrhunderts [1348, 772]. Jahrtausenden [1148]. James [472, 1220, 34, 466, 54, 1060, 1123, 326, 889, 845, 306, 1183, 1313, 275, 1101]. Jammer [1040, 444]. Jamshid [406]. Jan [777, 982, 1352]. Jane [800, 1117]. January [658, 820, 934, 1096, 1224, 1364, 106, 204, 378, 517]. Japan [578]. Jar [871]. Jean [1353, 720, 511, 1045, 1242, 1277, 1068, 730]. Jean-Baptiste [1242]. Jeanne [885, 1084]. Jeanne-Marie [1084]. Jenner [291]. Jerome [910]. Jersey [768]. Jesse [631]. Jessie [415]. Jet [1009]. Jevons [589]. Jh [1339]. Jirí [1264]. Joachim [1156, 1240, 23, 1348]. Joannes [71]. Jodrell [1279]. Johan [829]. Johann [292, 1187, 548, 1090]. Johannes [687, 98, 587, 810, 1264, 1041, 968, 1339]. John [573, 995, 1355, 546, 1353, 219, 906, 811, 338, 763, 1128, 1078, 629, 683, 909, 103, 805, 884, 1180, 1265, 1047, 848, 953, 282, 877, 16, 1326, 1275, 451, 109, 84, 493, 827, 1326, 773, 1126, 884, 623]. Johnson [152, 455]. Johnston [1194]. Jolan [499]. Joly [404]. Jonathan [988]. Jones [1347, 1048]. Joravsky [340]. Jorpes [1019]. Jose [583, 398, 58, 58]. Josef [995, 281, 195]. Joseph [1266, 573, 157, 483, 293, 536, 27, 140, 108, 1192, 136, 407, 103, 1160, 994, 512, 1125]. Josephy [577]. Josten [1124]. Journal [573, 1017, 763, 531, 370, 9]. Journals [415, 629]. jours [447]. Jozef [709]. Ju [334]. Judd [1073]. Juillet [286]. Jules [297]. Julian [778]. Julio [429].

July [1174, 1290]. June [842, 539, 831]. Jungiu [1348]. Jungk [34]. Juskevic [929, 160, 334, 1013, 1014]. jusqu’en [760]. Just [33]. Justus [911, 74, 627]. Jwaideh [135]. Jwarizmi [390].

Kac [1236]. Kader [478]. Kahn [906, 328]. K’ai [931]. K’ai-Wu [931]. kalten [847]. Kamil [967]. Kangro [1348]. Kanner [769]. Kantor [646]. Kapiteln [814]. Kaplan [998]. Kaplon [1015]. Kargon [1349]. Karl [15, 688, 674, 1021, 671, 1307, 369, 1156, 1240, 1090, 488, 1326]. Karl-Heinz [369]. Karman [1164]. Kartenbildes [716]. Kashi [406, 478]. Kazan [696]. Kazansky [487]. Keil [1305]. Keilin [1044]. Kekule [374, 487, 374]. Kelly [1017]. Kelvin [779]. Kemble [1042]. Kennedy [406, 478]. Kennissen [1092]. Kennzeichen [116]. Kensington [932]. Kepler [292, 363, 968, 1230, 98, 587, 810, 699, 870, 901, 114, 1100, 588, 811, 1072, 852]. Kepleriana [1345]. Kett [1266]. Kettlewell [778]. Keynes [339]. Keys [388]. Khachatur [856]. Khalaf [908]. Khimicheskogo [298]. Khwarizmi [388, 390, 1261]. Kibre [679]. Kiefer [805]. Kieffer [904]. Kilgour [376]. Kindi [698, 789]. Kinetic [425]. King [625, 575, 762, 1350, 31]. King-Hele [1350]. Kinnersley [775, 211]. Kip [326]. Kirchenvatern [281]. Kircher [207]. Kirk [65]. Kirkwood [811]. Kirwan [400]. Kitab [967, 918]. Kitabu’l [26, 25]. Kitabu’l-’anwa’ [26]. Klaproth [76]. Klassicheskoi [298]. Klassifikation [946]. Klaus [1040]. Klee [972]. Klemm [64, 1090]. Klibansky [795]. Kn’asev [1013]. Knjazev [1014]. Know [748]. Knowledge [20, 707, 674, 707]. Knowles [642, 1146]. Ko [1123, 1294]. Koch [195]. Koestler [841]. Koller [687]. Komarov [1176]. Kommentar [369]. Kompassmacher [1115]. Kompfner [648]. Komplementaritat [1040]. Kopelevic [1013, 334]. Kopernikus [443]. Korrespondenz [1040]. Koshtoiants [856]. Kosmologische [1148]. Kosmos [449, 484]. Kossel [1063]. Koyré [1235, 861, 99, 363, 121, 1235]. Kranzberg [1140, 1141]. Krisenzeiten [1359]. kristallografii [591]. Kristan [856]. Kristeller [133]. Kronick [439]. Kronig [888]. Krutikova [1013]. Kuhn [1128, 441]. Kulikovskii [774]. Kultur [32]. kultureller [1359]. Kung [931]. Kunstler [716]. Kureze [946]. Kurt [772, 369, 361]. Kurze [614]. Kushyar [843]. kutalas [719]. Kuznetsov [594].

L [1015, 1266, 804, 373, 92, 20, 447, 721, 995, 418, 616, 910, 413, 885, 725, 712, 816, 1021, 121, 961, 400, 1128, 537, 1058, 1040, 768, 592, 137, 327, 1340, 1022, 889, 306, 217, 160, 334, 621, 51, 967]. l-mugabala [967]. Labaree [72]. Labban [843]. laboratorium [374]. Laboratory [1335]. Labour [590]. Labour-Saving [590]. Laer [710]. Lafeuille [770]. Lagowski [1077, 1000]. Laidlaw [22]. Lamarck [654, 598, 751]. Lamarckism [695]. Lambert [89, 39]. Lammli [449]. Lancaster [760]. Lancelot [483, 442]. Lancre [1265]. Land [1008]. Landforms [715]. Lands [537]. Langmuir [1059]. Language [501, 226]. Lapage [471]. Lapeysonnie [1084]. lapidibus [882]. Laplace [522]. Lardomshistoriska [1142]. Large [421]. LaRocque [1189].

Larrabee [906]. Lasserre [850]. Last [131]. Late [619, 897, 700, 95, 1199, 385]. Later [545, 217, 1064]. Latin [881, 133, 965, 905, 430, 1169, 699, 148, 663, 679]. latins [1154]. Latitudes [865]. Laurent [854, 1324]. Lavoisier [100, 854, 990, 1290, 1032, 370, 411]. Lavoisiereana [9]. Law [1230, 2, 483, 1285, 115, 442, 440, 941]. Lawn [620]. Lawrence [1331]. Laws [866]. Lear [811]. Leben [419, 687, 552, 1213]. Leber [364]. Leberforschung [364]. Leclercq [1045, 1353]. Lecons [1191]. Lecture [930]. Lectures [928, 120]. Lee [1164, 438]. Leeuwenhoek [481, 164]. Lefebvre [128]. Leibniz [209, 372]. Leibniz. [1213]. Leicester [1237, 1247]. Leidecker [361]. Leighly [1265]. Leistung [76]. Leithauser [23]. Lelland [200]. Lemay [775]. l’Encyclopedie [551]. Lenoble [121, 46]. Leo [714, 769, 775]. Leon [453, 781]. Leonard [72, 159, 673]. Leonardo [1086, 269, 743, 530, 368]. Leonhard [32, 160, 334, 1013, 1014, 561, 1307, 929]. Leopard [165]. Leopold [732]. Lesky [857, 858]. Lesley [1016]. Leslie [1268, 543, 278]. L’esplorazione [878]. Lester [575, 31, 995]. Letter [264, 526]. lettera [847]. letteraria [1357]. Letters [949, 1190, 1021, 1312, 481, 493, 653, 764, 306]. lettre [511]. Lettres [419, 1212]. Leverrier [314]. Levey [843, 1120, 129, 967]. Levine [734]. Lévy [121]. Lewis [757, 764, 887, 680]. Lexicographic [701]. Lexicon [1047]. L’expérience [1242]. Ley [613, 1091]. Leyden [871]. Libellus [1087]. Liber [1264, 369, 1070, 943]. Liberales [195]. Librarian [239]. Library [1157, 495, 376, 542, 1143, 1118, 1209, 565, 412]. libri [70]. libros [635, 1011]. Licensing [1266]. Licht [369]. Lichtenberg [289]. Lichtenstein [549]. Lichtenthaeler [24]. Liebig [911, 627, 977]. Liebigs [74]. Life [1216, 139, 662, 571, 483, 1019, 773, 1163, 1282, 798, 105, 399, 1274, 375, 813, 200, 1129, 782, 526, 1352, 1294, 343, 164, 1124, 1113, 288, 1215, 655]. Light [112, 344]. Lighthouse [1227]. Lighthouses [158]. Lightning [947]. Lilian [1271]. Liljestrand [802]. Lind [137, 1340, 621]. Lindborg [969]. Lindeboom [731]. Lindsley [856]. Lineage [696]. Ling [407, 994, 136]. Lini [1128]. Lining [84]. Link [1064]. Linnaeus [1027, 1171]. Linne [446]. Lionnais [500]. Lippman [726]. Lists [133]. Literary [890, 881]. Literature [899, 1076, 150, 1197, 796, 620, 680, 226, 1297]. Lithographiae [548]. Little [1121]. Lived [300]. Lives [1002]. Living [475]. Livingston [1057]. Livingstone [414]. Livre [1341]. livres [511]. Lloyd [960, 124, 1303]. Lobachevskii [352]. local [430]. Locali [336]. Localium [430]. l’Océanographie. [1265]. Lockemann [218]. loco [1264]. Loeb [655]. Loenen [582]. Loewenberg [54]. Logic [15, 1236, 271, 777, 361, 467]. logica [805]. Logical [1020]. logique [589, 589]. Lomonosov [774, 990, 626]. London [711]. Long [1274, 647]. Longitude [1126]. Longitudes [570]. Loren [1246, 851, 586]. Lorentz [984, 671]. Lost [87, 585]. Losungen [450]. Lou [1055]. Louis [628, 514, 1353, 511, 47, 300, 343]. Louise [933, 1002]. Lovell [1279]. Lovett [1040]. Lowenberg [54]. Lowies [1108]. Loyd [1060]. Lucas [893]. Lucien [370]. Luck [955]. Ludwig [1046, 1271, 765, 898]. lumine [1241]. luminis [1264]. Luna [702]. Lunar [668, 1072, 939]. Lunatic [1275]. Lurie [514, 343]. Lychnos. [1142]. Lyell [312]. Lying [548]. Lynn [679, 907, 325, 477, 830]. Lyon [573]. Lyonet [482]. Lyons [327].

M [448, 583, 329, 631, 398, 1237, 1149, 880, 119, 332, 1009, 1247, 1073, 121, 800, 577, 871, 1060, 342, 537, 452, 1314, 487, 501, 159, 69, 327, 582, 1117, 656, 675, 1013, 1014, 1081, 499, 1057]. M. [774]. M.-A [121]. MacDonald [779]. Mach [814, 1326, 1108, 1326]. machen [369]. Machine [1360, 804, 230, 1076, 589, 127, 127, 197]. Machines [1006, 1009]. MacKinney [851]. Madden [215]. Maddison [1145, 1145]. Made [732]. Madeleine [100]. Madness [1302]. Magazine [533]. Magic [287, 226]. Magie [1182]. Magnete [102]. magnitudinis [1264]. Magnus [1273]. magyar [719]. magyarorszagi [499]. Maille [744]. Major [214, 319]. Makers [688, 809]. Making [617]. Malaria [734]. Malgaigne [847]. Malpighi [1205]. Malthusian [1036]. Man [1211, 1270, 1126, 1040, 200, 1259, 62, 954, 352, 18, 203, 754]. Management [688]. Mandelbaum [797]. Mandler [720]. Mani [364]. Manning [1315]. Manuel [550]. Manures [420]. Manuscript [235, 253, 832, 186, 252, 239, 462]. Manuscripta [1013, 1014]. Manuscripts [318, 259, 464, 996, 245, 256, 250, 260, 565, 436, 237, 851]. manuscrits [1154, 1314, 728]. Manzoni [703]. Mappemondes [1085]. Maps [225, 17]. Marcel [419, 1085]. Marcello [1205]. March [78, 790]. Marci [1264]. Marcus [1264]. Marek [1264]. Margaret [689, 722, 645]. Margery [1144]. Marginalia [163]. Maria [1241, 727]. Marie [729, 840, 708, 29, 1084, 412]. Marine [717]. Mario [13, 292, 502, 807]. maris [917]. Marjorie [59, 1274, 1299, 618, 1194]. Mark [1236, 410]. Markovnikov [487]. Marsak [159]. Marsh [1021]. Marshall [268, 1341, 68, 965]. Marston [1183]. Marti [222, 192]. Marti-Ibanez [222, 192]. Martin [225, 1017, 1120, 924, 129, 967, 770, 226, 843, 76]. Marvin [843]. Marxian [93]. Marxism [340]. Marxist [523]. Mary [1237]. Masao [611]. Mass [1232, 444]. Massachusetts [1018]. Matematica [531]. Material [533, 992, 848]. Materials [1337]. Mathematical [959, 424, 1304, 40, 1197, 147, 1286, 1351]. Mathematics [996, 958, 850, 293, 136, 60, 226, 720, 713, 568, 293, 1333, 531, 1236]. mathématique [453, 500]. Mathematiques [297]. Mathematische [587, 1115]. mathesis [1241]. Mathison [62]. Matrix [1270]. Matter [36, 80, 107, 144, 1, 37, 81, 111, 168, 179, 205, 234, 145, 180, 181, 206, 238, 308, 347, 380, 309, 236, 261, 310, 349, 422, 456, 489, 519, 381, 423, 457, 490, 553, 595, 632, 660, 520, 554, 596, 633, 691, 738, 785, 822, 661, 692, 739, 819, 859, 891, 912, 936, 823, 860, 892, 913, 974, 1024, 1061, 1094, 1097, 937, 975, 1025, 1062, 1098, 1130, 1166, 1195, 1221, 1225, 1099, 1131, 1167, 1196, 1223, 1248, 1280, 1318, 1361, 1365, 1226, 1249, 1281, 1319, 1363, 997]. Matter [1101, 38, 720, 990, 933, 539]. Matthew [332]. Matthews [1268, 543]. Matthias [450, 651]. Maurice [155, 1354, 797, 1162, 576]. Maurois [105]. Maurolyco [87]. Max [1074, 1163, 1345, 1040, 444, 685, 77]. Maxime [1182]. Maximilian [1115]. Maxwell [779, 425, 981, 997, 914]. Maya [748]. mayores [728]. Mayr [574]. Maziarz [1304]. Mazlish [269]. McC [94]. McDermott [629, 877]. McKenzie [319]. McKie [1174]. McLanathan [1086]. McMullin [1211]. Meade [1267]. Meaning [124]. Measurement [175, 631, 1235, 172]. Measures [1262]. Measuring [1008, 164]. mécanique [453]. Meccaniche [926]. mechanica [1014]. Mechanical [666, 1049, 1042, 581, 994]. Mechanics [371, 834, 314, 295, 68, 926, 1040, 1343]. Mechanist [1282, 377]. Mechanistic [655]. méconnu [453]. Medal [521]. Medecine [223, 1181, 1154, 1182, 24]. Mediaeval [679, 133]. Medical [1266, 903, 1263, 885, 192, 575, 723, 376, 505, 953, 542, 1143, 438, 678, 31, 174, 565, 1266, 851, 1180]. medici [875]. Medicina [58]. Medicine [1056, 1266, 767, 783, 1033, 910, 1123, 1066, 508, 1209, 565, 193, 1215, 128, 1118, 1209, 222, 1266, 1157, 194, 1271, 508, 1118, 222, 1162, 1155, 1088, 1004, 221, 768, 1153, 1303]. Medicine. [1084]. Medieval [1266, 620, 903, 546, 851, 680, 407, 897, 1153, 1272, 1341, 171, 1228, 1199, 663, 132, 477]. médiévaux [728]. Medizin [1269, 1046, 1041, 281]. Medizinische [369, 857, 858, 584]. Meine [973]. Mel [1093]. Melancholy [795]. Melissus [582]. Mel’kumov [1009]. Melvin [1140, 548, 1141]. Memoir [1290, 1215]. Memoires [624]. Memoirs [1178]. Memorable [662]. Men [737, 1006, 119, 1335, 1059, 1002]. Mendel [1054, 1053, 1106]. Mendeleev [485, 1232, 1029, 1323]. Mendelism [1150]. Mendels [614]. Mendelsohn [798]. Mens [1092]. Mensch [77]. Menschen [369]. Mental [1302]. Mentally [769, 1018]. mentis [910]. Menut [1341]. Mercatorienne [452]. Merchant [763]. Mercurial [571]. Mercury [1060]. Meriwether [764, 887]. Merk [972]. Merton [1253]. Mesmerism [1234]. Mesopotamia [129]. Messenger [852]. Metallography [279]. Metallurgy [580, 580]. Metals [278]. Metaphysical [507]. Metaphysics [1235]. Metaphysik [1306]. Metaphysique [297]. Meteorology [744, 26, 1146, 609]. Method [997, 394, 1027, 224, 377, 330, 540]. méthodologie [1278]. Methodology [467, 134]. metodo [1186]. Metonic [748, 835]. Metric [986, 1254, 836, 393, 837]. Metrie [197]. Mette [1269]. Meyer [1040]. Meyer-Abich [1040]. Michael [932, 644, 1040, 999, 23, 113, 907, 816]. Michel [1145, 1212, 1074, 1212, 515]. Michèle [1336]. Michelson [344, 1250, 945, 944]. Microbiologists [402]. Microbiology [401]. Middle [597, 1064, 545, 68, 965]. Middleton [642, 1146]. Midonick [713]. Midwest [486]. Midwife [1010]. Miessner [648]. Mikhailov [1013, 1014]. Milestones [1051, 401]. Millas [583, 398]. Millas-Vallicrosa [583, 398]. Mille [1081]. Miller [783, 1219, 720]. Millhauser [33]. Millikan [631]. Milling [230]. Milton [33, 222]. Mind [14, 910, 997]. Mineralien [70]. Minerals [1273, 417]. Ming [223, 700, 385]. Minimal [693]. Mirandola [1344]. Mirko [1267]. Mirrors [207]. Missile [972]. Missing [209]. Missionary [1017]. Mississippi [877]. Mistranslation [432]. Mittel [765]. Mittel-Europas [765]. Mittelalters [1305, 195]. Mitterling [274]. Mivart [375]. Models [354, 1170]. Modern [1051, 1266, 1006, 611, 14, 303, 13, 761, 757, 132, 586, 1140, 1151, 172, 1316, 444, 451]. moderne [1267, 121]. modernen [1156, 814]. Mohammad [285]. Moholy [1270]. Moholy-Nagy [1270]. Molecular [1229]. Møller [1317]. Moments [119]. Monaco [1265]. Mond [1115]. Mond- [1115]. Monde [284, 1341]. Monkeys [391]. Monro [723]. Monte [1049]. Monumenta [1085]. Moon [391, 433, 463]. Moon-Happy [391]. Moorat [565]. Moore [1040]. Moorish [908]. Morazé [121]. Mordecai [967]. Morgan [1358, 971, 889, 493]. Morgenbesser [1207, 270]. Morison [1006]. Morley [344]. Morris [362]. Morse [50]. Mortem [1132]. Morton [1015, 473]. Moscovici [1242]. Moseley [1256, 896, 985, 1038]. Moses [52]. Most [732]. motibus [1068]. Motion [1230, 866, 926, 597, 148, 978, 1206, 430]. Motions [1341]. Motte [458, 1351]. Motu [336, 83]. Motuum [430]. Motz [960]. Mountain [59, 59]. mouvement [1242]. moved [148, 663]. Mover [871]. moveri [148]. Moves [262]. movetur [148, 663, 663]. Moyen [1154, 1084]. MS [697, 1340]. Mss [565]. Mueller [447]. Muffet [1179]. mugabala [967]. Mugler [66]. Muhammad [26, 25]. Mu‘jam [135]. Muller [687, 1187, 160]. Multhauf [1116, 1303]. Mundi [296]. mundo [875]. Munilla [1017]. Murray [955]. Museum [697, 624]. Music [226]. Musical [726]. Muslim [26]. Mutanna’ [390]. Muthanna [1261, 390]. My [1163, 1274, 973, 1163, 763]. Mysterium [810]. Mystery [933]. Mystiques [99]. Myth [568].

N [485, 996, 20, 564, 194, 690, 325, 657, 1075, 567, 1013, 1014, 1018, 1339]. Nachtwey [190]. nad [1053]. Nadir [314]. Nagel [467]. Nagy [1270]. Nais [331]. Name [1058]. Names [1087]. Napoleon [1354]. Nartov [334]. Nash [673]. Nasr [1210]. Natalie [856]. Nathan [735]. Nation [927]. National [1093, 1157, 784, 1118, 1022, 1147, 1209, 624, 1178]. Natur [369]. natura [70, 1264]. Natural [2, 340, 639, 840, 54, 764, 1012, 1027, 1283, 732, 673, 470, 771, 1351, 795, 97]. Naturalist [1190, 1310, 1031]. Naturalisti [1357]. Naturalisti-Esploratori [1357]. naturalistica [878]. Naturalists [628]. Nature [722, 1177, 1309, 754, 673, 271, 710, 895, 750, 271, 906, 624, 18]. naturelle [624, 1016]. Naturis [464, 436]. Naturphilosophie [302]. Naturwissenschaft [397]. Naturwissenschaften [122, 1156]. naturwissenschaftliche [369]. Nauk [216]. Naval [1356]. Navigation [28, 416]. Navigations [927]. Nazi [1218]. Neal [330]. Near [1083, 440, 683]. necessary [148, 663]. necesse [148]. Nedelandse [329]. Needham [121, 136, 407, 994, 157, 27]. Needs [1082]. Nef [848]. Neglected [1184]. Nehemiah [928, 2]. Nei [1123, 1155, 1294]. Neil [1073]. Neo [695]. Neo-Lamarckism [695]. Neptune [558, 473]. Nesta [573]. Netherlanders [413]. Neu [953]. Neue [366]. neuen [55]. Neugebauer [367, 849]. Neumann [16]. Neuplatonismus [584]. Neurological [964]. Neurophysiologist [1217]. News [794, 12, 49, 90, 118, 154, 188, 213, 267, 317, 357, 396, 437, 465, 497, 608, 535, 563, 607, 638, 670, 706, 920, 839, 874, 902, 919, 950, 993, 1039, 1071, 1139, 1175, 1204, 1260, 1296, 1330, 1111, 1233, 752]. Newsome [771]. Newton [161, 162, 755, 909, 412, 513, 1351, 550, 235, 112, 315, 458, 983, 834, 38, 1037, 115, 186, 265, 209, 163, 434, 1295, 978, 313, 387, 893, 833, 1349]. Newtonian [314]. Niccolo [196]. Nicolaas [73]. Nicolaus [930]. Nicole [966, 1341, 1253, 85, 745, 1341]. Nicolson [1274, 59]. Nieburg [1058]. Niels [1040, 707, 1040, 1129]. Nikolai [352]. Nikolaus [364]. Nineteenth [815, 1356, 201, 1262, 1005, 537, 60, 735, 590, 95, 656, 600, 1292, 182, 241, 426, 938, 492, 696, 350]. Nineteenth-Century [815, 1005, 735, 600, 1292, 182, 241, 426, 938, 492]. Ninetieth [820]. Ninety [934, 1096, 1224, 1364]. Ninety-First [934]. Ninety-Fourth [1364]. Ninety-Second [1096]. Ninety-Third [1224]. Ninth [658]. Nissen [1041]. Niter [525]. Nitric [743]. No [880, 534, 1209, 705]. Noah [1206]. Non [352, 315]. Non-Euclidean [352]. Nora [1219, 1312]. Norbert [1358]. Norman [998, 734, 557]. North [516, 455, 734]. Norwood [593]. nos [447]. nostro [875]. Note [984, 1291, 983, 834, 83, 9, 835, 701, 1255, 792]. Notes [117, 356, 86, 458, 48, 212, 266, 791, 187, 11, 88, 177, 1124, 316, 153]. notis [810]. nova [875]. novae [1264]. Novalis [226]. Novy [537]. Nuclear [1218, 690]. Nuee [786]. Nuevos [398]. Numerical [354, 394]. Nurseryman [573]. Nymphis [369].

O [1077, 713, 367, 194, 759, 849, 305]. O.F.M [1017]. Oberth [1009]. Object [3]. Objectivity [91]. Objects [1112, 1034]. Obras [1243]. O’Brien [856]. Observation [758]. Observations [1149, 1035, 241, 1289, 1295, 524]. Observatory [495, 273, 273]. Observe [1109]. Observer [602]. Occasional [707]. occident [1267, 1081]. Ocean [862, 1060]. Oceanography [1079]. Ockham [635]. O’D [758]. Odington [1286]. Oersted [322]. oeuvre [733, 100, 409, 1211]. Oeuvres [408]. Oil [1137]. Okapi [392]. Ola [653]. Olby [1150]. Old [868, 440]. Oldenburg [353]. Oldest [150]. Olkusz [225]. O’Malley [292]. Omne [148, 663]. Omnia [71]. omnimodis [1187]. Onderrichts [374]. Onderrichts-laboratorium [374]. Ong [224, 321]. ont [365]. Ontologie [1306]. Opera [1185, 1014, 71]. operation [1147]. Opere [807]. Opinion [1272]. Oporino [847]. Oporinus [511]. Oppenheimer [800, 1117, 1331]. Opportunities [1082]. Opposite [1251]. Opposition [700, 385]. Oprichting [374]. Optics [833, 1277]. Optique [1052]. Opty [1053]. Oral [253, 1215]. Order [214, 337]. Oresme [966, 1341, 1253, 85, 745]. Organic [911, 1324, 515, 1332]. Organization [1147]. Origin [8, 50, 786, 411, 747, 1189, 507, 667, 684, 826, 61]. Original [1051]. Originality [599]. originaux [589]. origines [446, 576]. origini [1186]. Origins [358, 1054, 299, 498, 1288, 757, 544, 333, 961, 386, 1324, 328, 1116, 439, 545, 853, 1078, 1151, 1150]. Orlando [1216]. Ornithologen [765]. Ornithology [621]. ortu [1264]. Oskar [133]. Ostrovitianov [216]. Other [1017, 868, 928, 435, 1124]. Otto [972, 1091]. Ovid [506, 506]. Oviedo [97]. Owen [570, 740]. Owsei [1271, 1303, 51]. Oxidation [1214]. Oxygen [1214].

Oye [1092].

P [368, 774, 521, 1317, 1149, 359, 303, 121, 1311, 400, 410, 501, 1116, 199, 929, 715, 710, 1303, 160, 334, 1013, 1014]. P. [744]. Pace [72]. Pacific [324]. Pack [881]. Padua [451]. Page [743]. Pagel [1114, 584]. Palaontologie [323]. Palatii [880]. Palisca [472]. Palissy [408]. Pallas [626, 1190]. Palmer [844]. Pamela [585]. Pannekoek [416, 503]. Panofsky [795]. Pao [1123, 1294]. Pao-p’u [1123]. Paolucci [466]. Paper [151]. Papers [240, 722, 949, 72, 254, 487, 884, 412, 275, 532, 888]. Papp [58, 445]. Papyrus [183]. Paracelse [99]. Paracelsian [525]. Paracelsus [369, 584, 369]. Paradise [585, 43]. paradoxes [335]. Paradoxicae [335]. paradoxie [450]. Parasites [193]. Parasitic [193]. Parasitology [1043]. Pare [847]. Parents [1328]. Paris [1088, 9]. Parish [1119]. Park [722]. Parker [367, 849]. Parliament [986, 836, 393, 837]. Parmenides [582]. Part [996, 801, 1214, 710, 994, 745]. Partial [1073]. Participants [259]. Particle [1295]. Parting [743]. Partington [399, 272, 504, 541]. Parts [693]. Party [1246]. Pasquale [332]. Passages [435]. Passions [650]. Pastor [429]. Path [1100]. Pathfinder [1164]. Pathologie [364]. Pathology [647, 911]. Patricia [1353, 811]. Patrick [467]. Pattern [1322]. pauca [966]. Paul [1092, 94, 997, 1212, 650, 1128, 137, 782, 402, 507, 133, 47]. Paul-Henri [1212]. Paul-Louis [47]. Pauli [888]. Pavon [625]. Payne [126, 1121]. Peale [415]. Pearce [816, 961]. Pearl [679]. Pearson [1326, 1326]. Peckham [50]. Pedro [583, 1017]. Peiresc [264]. Pellat [1342]. Pemberton [458]. Pendulum [1201]. Penicillin [105]. Penick [889]. Pennant [1190]. Pennsylvania [767]. Pens [1002]. pensée [500, 1235, 551]. pensées [101]. Pentagram [1263]. Perception [3, 797, 720]. Perelmuter [1014]. Period [976]. Periodical [52]. Periodicals [995, 439]. Perrault [1189]. Persia [1122]. Persian [508]. Personal [116]. Personne [453]. Persons [683]. Perspectives [321, 1045]. Pertaining [686]. Peru [625]. Peter [815, 964, 1190, 220]. Peters [275]. Petersburger [160, 334]. Petit [446]. Petroleum [202]. Petruck [843]. Petrus [1012, 1308]. Petty [666]. Pfeiffer [1055]. pharaonique [128]. Pharmaceutical [953, 908, 1080]. Pharmacology [802]. Pharmacy [543]. Phenomenes [1191]. Phenomenon [203]. Philemon [771]. Philip [283, 403, 274, 1332]. Philipp [858]. Philolaos [649]. Philoponos [1339]. Philosopher [1184, 813]. Philosophers [65]. philosophia [875]. philosophiam [1185]. Philosophical [533, 828, 1299, 955, 928, 593, 951, 641]. Philosophie [1278]. philosophique [453]. philosophische [369]. Philosophy [467, 92, 20, 997, 677, 666, 1283, 215, 470, 710, 226, 795, 840, 1304, 220, 582, 566, 1351, 797, 1207, 270]. Phoenix [43]. Photographic [872]. Photographs [1256]. Phrenology [342]. Physica [528]. Physical [992, 619, 1292, 1328, 172, 214, 938, 540, 147, 609, 405, 994, 468, 1042]. Physician [1307, 486]. Physicians [711]. Physicist [569]. Physicists [1040]. Physico [1241]. Physico-mathesis [1241]. physicorum [635, 1011].

Physics [424, 600, 459, 171, 1134, 1128, 1040, 707, 1059, 147, 1199, 663, 994, 123, 303, 808, 362, 806, 444, 951, 67]. Physik [814, 651]. Physikers [1090]. Physiologe [198]. Physiologia. [1012]. Physiological [94, 386, 977, 183, 377]. Physiologie [364]. Physiology [911, 604, 868, 377, 856]. physique [453]. picatrix [791]. Piccolomini [792, 1252]. Pico [1344]. P’ien [1123, 1294]. Pierre [372, 284, 1267, 203, 223, 1045, 1189, 335, 1081, 365, 459, 1200, 482]. Pilkington [228]. Pilot [242]. Pioneer [628, 388, 1164, 1328, 744, 486, 515, 623, 402]. Pioneers [301, 1009, 924, 1007]. Pitagorici. [727]. Place [13, 345, 273, 384, 946, 304]. Plan [1017]. Planck [77, 671]. Planet [758, 558]. Planetary [354, 1230, 406, 1009, 570, 897, 865]. Planets [865]. plans [589]. Plant [2, 573, 824]. Plants [1293, 925, 928, 1002, 61]. Plate [1244]. Plates [872]. Plato [980, 850]. Platon [284, 649]. Platonist [549]. Plinius [1069]. Pliny [788, 771]. Pneumatic [1287]. Poesch [415]. Poesie [331]. Poetry [1048]. Poggendorff [900]. Poincaré [940]. Poison [124]. Polar [400]. Policy [688, 426, 1147, 784, 1147]. Polinière [1200]. Politics [889, 1147, 690]. politischen [1359]. Polydore [894]. Pomerans [1023]. Pope [1274]. Popkin [360]. Popov [648]. Popper [15, 674]. por [390]. Portrait [618]. Portrats [716]. Portsmouth [412]. Positivism [656]. Positivist [75]. Positron [593]. possession [1182]. Post [1132, 1272]. Post-Medieval [1272]. Post-Mortem [1132]. Postcript [704]. Posthumous [1072]. Potassium [428]. Potential [981]. Potocki [225]. Poulle [1084]. Poulle-Drieux [1084]. Pouring [1137]. Powell [1048]. Power [40, 688]. Poynter [20]. Practica [1185]. Practice [1112, 1034]. Practices [1008]. Practitioners [959, 903]. Precision [1007]. Precursor [2, 1323]. Preface [511]. Prefazione [847]. Preformation [491]. Prehistory [580, 653]. Premier [1265]. Premières [101]. Present [833, 1337, 214, 889, 672]. Preservation [254, 256, 244]. President [146, 149]. Presocratic [65]. Presocratics [677]. Press [52, 439]. Presuppositions [1272]. Price [320, 407]. Priestley [512, 1125, 1160]. Primary [1286]. Prime [871]. primera [531]. Primitive [601]. Prince [732]. Principal [416]. Principall [927]. Principia [458, 432, 909, 983]. Principle [13, 597, 825, 663]. Principles [1351, 1301, 843]. Printed [86, 1157, 505, 953, 542, 1143]. prius [1264]. Private [254]. Prize [395]. Probability [1285, 1272, 1272]. Problem [318, 620, 1104, 260, 428]. Problemata [1340]. Probleme [614]. problemes [1192]. Problems [268, 177, 641, 183, 377, 1009]. Procedures [327]. Proceedings [793, 268]. Process [4, 976]. Proclus [544]. Profession [12, 49, 90, 118, 154, 188, 213, 267, 317, 357, 396, 437, 465, 497, 608, 535, 1266, 776, 885]. Professional [1049]. Professionalization [976]. Professor [1066, 630]. Progress [867, 176, 23]. Project [1060, 45]. Prolegomenon [1198]. Proof [1253]. Propaedeutica [1185]. Properties [631]. Proportionibus [85, 966]. Proportionum [430, 745, 966, 85]. Proposals [251]. Propulsion [1009]. Prosdocimus [1253]. Prospect [1236].

Protection [683, 732]. protestantisme [948]. Protoplasmic [1282]. Prove [262]. Proxy [1328]. Przibram [671]. Psalmen [369]. Psalters [369]. pseudo [404]. pseudo-hippocratique [404]. Psychiatry [1301, 1151]. Psycho [1021]. Psycho-Analytic [1021]. Psychoanalysis [653]. Psychoanalytic [964, 924]. Psychological [757]. psychologie [447]. Psychologist [949, 421]. Psychologists [1238]. Psychology [175, 721, 801, 646, 678, 1138, 720]. Psychoneuroses [653]. p’u [1123, 1294]. Public [885, 21, 426, 138]. Publication [893]. Publications [142, 78, 35, 166, 438]. publiée [121]. Pure [426, 1147]. Pursell [889, 1140, 1141]. Purver [1144]. Pygmaeis [369]. Pythagoras [649, 185, 1184, 1184]. Pythagorean [6, 827]. Pythagoreanism [1184]. Pythagoreans [677].

Qattara [601]. Qualities [1286, 1341]. Quality [1197]. Quantentheorie [1040]. Quantification [171, 178, 173, 177, 174, 176, 170, 169]. Quantities [147]. Quantum [1128, 671, 1040, 940]. Queer [1251]. Quellen [1240]. Query [605, 1258, 701, 534, 1259, 705]. Quesiti [196]. Questio [917]. Question [8]. Questioners [1040]. Questiones [44]. Questions [620]. Quill [1126]. Quinn [927]. Quintessence [1059]. Quiricus [1158]. Quiros [1017]. quod [148, 663]. Qurra [527, 6, 827, 509]. Qutayaba [26].

R [647, 15, 959, 1145, 957, 538, 714, 367, 1079, 1054, 580, 585, 888, 674, 100, 121, 272, 399, 504, 676, 341, 690, 646, 1123, 17, 202, 541, 1010, 681, 199, 137, 1340, 510, 849, 271, 344, 95, 812, 1313, 280, 193, 1150, 63, 621]. Race [165, 938]. Raderuhren [1115]. Radiation [1103, 524]. Radio [648]. Radioactivity [780]. Radiology [1317]. Radiometer [863]. Radiotelegraphy [648]. Raef [1121]. Rafael [461, 742]. Ragnar [1217]. Rainbow [568, 867]. Rainer [1305]. Raleigh [927]. Ramesside [849]. Ramsay [415]. Ramus [1308, 224]. Rand [688]. Randall [906, 282, 451]. Rapport [922]. Rarick [152]. Raskin [1013, 1014]. Raspail [1245]. Rastitel’nymi [1053]. Rather [910, 200]. Rational [371, 549, 1283, 295]. Rauschenberg [1310]. Rauwolf [1307]. Raven [65]. Rayleigh [945, 944]. Raymond [764, 410, 795]. Razi [25]. Razumovskij [334]. Re [1087]. Reactive [1009]. Read [928]. Reader [52]. Reading [1353, 564]. Readings [904, 1298, 802, 845]. Reality [312]. Reappraised [1015]. Reason [263, 224, 681]. Reasonable [289]. Rebel [289]. Received [142, 78, 35, 166]. Reception [52, 1028, 1231]. Rechenbuch [772]. recherches [482, 404]. Récit [69]. Reckoning [843]. Recognition [1322, 824]. Reconsidered [199]. Record [17, 831]. Recorde [838, 1067]. Recording [311]. Records [246, 244]. recours [948]. Redfield [722, 722]. Redivivus [43]. Reduction [354, 1285]. Reductionism [977]. Reed [1076]. Reference [1180, 908]. Reflections [1320, 1215]. refluxu [917]. Reformer [1245]. Refraction [1070, 943]. Refutation [148]. Refutations [674]. regii [875]. Regime [404]. regimine [910]. Regiomontanus [1291, 1187]. Register [903]. Regress [867]. Reingold [735]. Reino [304]. Reinterpretation [582]. Reisewerk [484]. Rekenboeken [329]. Related [764, 653]. Relating [1017, 461, 1124]. Relation [185, 220, 226]. Religion [1123, 226, 795]. religions [369]. religions-philosophische [369]. Religious [1112, 657, 1034]. Remarks [987]. Rembert [419]. Rembrandt [73]. Reminiscences [231]. Renaissance [620, 729, 680, 133, 331, 330]. Rene [1023, 454, 121]. Rennie [773]. Repaired [1012]. Reply [463, 534, 837, 705, 387]. Report [1128]. Reports [1051]. Representation [147]. reprobatione [1169]. rerum [1105, 464, 436]. Research [1093, 1033, 1218, 719, 241, 833, 1176]. Resolve [741]. Resolved [826]. Resonance [523]. Resources [1149]. respicientes [966]. Respiration [1044, 42, 377]. Responses [657]. Responsibility [257]. Retarded [769]. Retiring [146]. Retreat [683]. Retrospect [1236]. Review [1360, 469, 640, 419, 647, 1056, 709, 467, 956, 336, 1092, 579, 1015, 1159, 1211, 547, 59, 331, 414, 472, 1220, 368, 485, 774, 1051, 1053, 34, 689, 780, 1216, 1317, 1077, 996, 15, 122, 340, 448, 688, 1266, 992, 127, 804, 1074, 139, 583, 724, 959, 1007, 1115, 1145, 776, 957, 1093, 1239, 372, 722, 1279, 573, 628, 1088, 932, 230, 815, 767, 890, 1004, 964, 120, 329, 94, 157, 373, 1346, 631, 1308, 324, 92, 134, 297, 398, 445, 500, 589, 713, 615, 538, 1237, 20, 408, 714, 421, 447, 721, 1246, 367, 1177, 818, 995]. Review [620, 903, 1157, 1268, 1261, 571, 22, 483, 1079, 1315, 1349, 1356, 320, 644, 1148, 882, 1006, 687, 733, 52, 639, 737, 778, 958, 1008, 1355, 301, 418, 1054, 285, 323, 878, 1149, 1357, 1347, 648, 1263, 284, 679, 880, 229, 817, 930, 1114, 783, 801, 580, 72, 119, 155, 161, 359, 371, 729, 926, 415, 446, 514, 654, 1003, 1234, 1358, 300, 546, 1270, 337, 549, 585, 1164, 911, 543, 846, 420, 1019, 850, 1206, 30, 611, 809, 1188, 1267, 1348, 225, 406, 685, 717, 287, 622, 840, 907, 1353, 283, 568, 102, 616, 617]. Review [482, 910, 14, 997, 101, 413, 619, 123, 196, 303, 332, 466, 1212, 708, 13, 885, 219, 1218, 1017, 574, 54, 141, 784, 972, 1009, 650, 711, 851, 881, 1154, 1306, 201, 281, 886, 1176, 1244, 293, 772, 1262, 296, 725, 626, 761, 781, 888, 1247, 712, 299, 773, 1082, 1337, 498, 674, 816, 922, 221, 757, 777, 64, 70, 1120, 1126, 1163, 1359, 1245, 764, 971, 677, 1269, 1052, 19, 1304, 906, 1086, 844, 53, 302, 536, 811, 1345, 564, 1142, 1190, 1073, 65, 798, 1278, 1089, 27, 474, 591, 594, 610, 132, 544, 680, 728]. Review [808, 1185, 218, 904, 33, 403, 652, 194, 1184, 1219, 503, 1021, 1055, 345, 769, 58, 97, 105, 140, 370, 682, 91, 100, 121, 162, 269, 272, 338, 339, 399, 504, 586, 800, 1235, 203, 1158, 1342, 763, 730, 1354, 473, 577, 923, 676, 1083, 642, 50, 192, 1210, 775, 875, 1023, 1165, 1186, 1311, 98, 292, 587, 588, 613, 810, 1072, 1264, 341, 970, 77, 333, 755, 509, 441, 961, 690, 643, 74, 274, 362, 374, 400, 516, 1060, 951, 289, 342, 410, 1128, 625, 1005, 1080, 1333, 552, 575, 718, 1191, 1192, 1078, 1012, 268, 135]. Review [165, 136, 646, 1046, 1123, 1182, 223, 1331, 476, 1140, 537, 407, 76, 731, 629, 683, 1151, 1187, 1310, 60, 1076, 1274, 93, 216, 471, 195, 924, 1161, 1238, 1058, 909, 960, 1214, 852, 158, 273, 570, 719, 1178, 325, 126, 328, 17, 452, 671, 1314, 202, 326, 397, 1146, 720, 1213, 1343, 1217, 1271, 1040, 758, 1049, 291, 802, 723, 753, 29, 103, 129, 298, 411, 487, 501, 541, 627, 759, 1116, 130, 1241, 797, 375, 813, 1312, 1010, 231, 481, 1307, 159, 1344, 1091, 1042, 1298, 857, 858, 766, 197, 294, 409, 681, 279, 805, 511]. Review [1236, 756, 199, 276, 765, 1045, 768, 884, 1041, 592, 1153, 1305, 278, 125, 1180, 475, 657, 450, 1265, 220, 21, 1119, 1242, 69, 200, 376, 486, 505, 762, 28, 1309, 99, 1047, 214, 732, 760, 848, 953, 1273, 224, 26, 908, 477, 1044, 1205, 1335, 515, 707, 1059, 1129, 966, 1020, 137, 138, 327, 847, 1340, 71, 104, 307, 551, 929, 417, 1189, 715, 369, 508, 542, 1143, 726, 933, 502, 540, 506, 754, 282, 782, 582, 510, 849, 1299, 96, 335, 401, 402, 641, 277, 438, 1118, 855, 1022, 624, 198, 584, 1350, 295, 673]. Review [998, 1316, 67, 439, 609, 876, 889, 952, 955, 999, 1050, 1127, 1147, 215, 361, 455, 507, 1276, 655, 360, 1016, 1209, 286, 363, 443, 1156, 1240, 1336, 735, 967, 969, 1075, 1048, 1144, 1297, 845, 548, 1272, 271, 566, 131, 545, 24, 222, 678, 1208, 1117, 1162, 1352, 470, 306, 653, 968, 31, 319, 512, 590, 1125, 156, 807, 217, 796, 1000, 973, 567, 710, 56, 734, 770, 877, 736, 228, 344, 854, 841, 931, 1155, 716, 1183, 1160, 95, 1243, 1300, 1334, 1338, 1090, 449, 618, 62, 124, 779, 191, 330, 75]. Review [656, 581, 1152, 227, 404, 623, 771, 925, 928, 1002, 1087, 1121, 1179, 343, 684, 812, 614, 963, 128, 853, 645, 806, 799, 366, 480, 883, 346, 1194, 1207, 1313, 1084, 16, 73, 565, 675, 1303, 1332, 814, 133, 280, 66, 651, 270, 442, 1043, 32, 160, 226, 334, 444, 1013, 1014, 1339, 1277, 23, 164, 193, 364, 1018, 1081, 1275, 1301, 499, 405, 843, 954, 479, 686, 1150, 1181, 1341, 1085, 593, 962, 887, 927, 57, 484, 612, 569, 578, 842, 1011, 803, 1122, 412, 513, 1124, 1351, 1302, 290, 63, 576]. Review [795, 994, 453, 468, 1113, 550, 288, 305, 322, 365, 454, 539, 672, 1001, 1141, 68, 189, 304, 377, 572, 1057, 1215, 488, 1193, 416, 440, 25, 478, 965, 856, 18, 61, 190, 275, 321, 621, 630, 879, 905, 451, 51, 55, 649, 727]. Reviews [686, 1147]. revista [531]. Revolution [296, 370, 363, 140, 53, 1235, 853]. Revolutionary [682]. Revolutions [1125, 441]. Rex [723]. Rey [429, 583]. Rhetikus [1156, 1240]. Rhetoric [1048]. Rheumatic [766]. Rhodian [1177]. Rhys [472]. Richard [1266, 1004, 818, 1157, 1267, 221, 1086, 808, 1298, 732, 307, 715, 360, 62, 887, 927, 1011, 468, 740]. Richeson [1008]. Richmond [506]. Ricker [1041]. Riddle [1134]. Ring [1093, 1063]. Rings [492]. Rise [688, 776, 958, 848, 952]. Ristich [1336]. Ritchie [1356]. Rittenhouse [883]. Ritter [1090]. Ritterbush [1332]. Ritual [1112, 1034]. Rivers [1215]. Rivière [100]. Roads [995]. Robbins [932, 846]. Robert [34, 127, 722, 230, 631, 1349, 1149, 1234, 337, 1244, 33, 194, 338, 339, 763, 588, 74, 625, 1238, 719, 126, 1116, 130, 505, 1020, 715, 369, 1022, 566, 1160, 1300, 404, 1303, 1011, 190, 2, 557, 184, 528, 917, 838, 1067, 559, 529, 46, 602, 1108, 840, 29, 228]. Roberval [453]. Robinson [717]. Rochot [335]. Rocket [1009, 1009]. Rockets [1009]. Rodinson [1182]. Rodolico [878, 1357]. Rods [947]. Roemer [222]. Roger [881, 1185, 867, 228, 551, 624, 483]. Rogers [884, 448]. Role [1266, 862, 958, 244]. Rolf [969]. Roller [102]. Rolt [804]. Roman [881, 581, 1152, 440, 788, 545]. romantischen [1090]. Rome [440]. Romer [780]. Ronald [817, 1262, 730]. Roofe [137]. Roos [1180]. Root [787, 478]. Rosecrance [690]. Rosen [852, 21, 1302]. Rosenberg [736]. Rosenfeld [1059]. Rosenthal [1342]. Rosholt [1022]. Rossii [373, 591]. Rostand [121]. Rotating [265]. Rotation [862]. Rounds [440]. Rousseau [1274, 365]. Roy [1310]. Royal [1268, 711, 338, 333, 495, 277, 1144, 928, 358]. Royaumont [286]. Rozental [1129]. Ruben [138]. Rudolf [648, 1305, 200]. Ruffin [420]. Rufus [332]. Ruins [725]. Ruiz [625]. Rule [1253]. Rumford [1309, 569]. Runar [886]. Rupert [708, 755, 412]. Russel [813]. Russell [593, 799]. Russia [373, 1190, 856]. Russian [1176, 610]. Russians [516]. russiche [626]. Russische [32]. russischen [32]. Ruth [1040]. Rybka [709].

S [472, 996, 688, 127, 230, 589, 721, 1356, 1355, 229, 1358, 406, 619, 757, 564, 65, 680, 399, 441, 1060, 1128, 575, 1274, 487, 231, 1298, 766, 294, 805, 848, 1129, 67, 1147, 31, 1338, 73, 565, 803, 189, 478]. S.J [483]. Sa [482, 409, 1211]. Sadi [82, 210, 431]. Saffouri [285]. Saga [723]. Saint [770, 733]. Saint-Gille [770]. Saint-Hilaire [733]. Salamandris [369]. Salamanque [728]. Salernitan [620]. Salmon [1208]. Salome [1055]. Salomon [958]. Sambursky [619, 67]. Samfundets [1142]. Sami [908]. Samos [1035]. Samtliche [369, 369, 369]. Samuel [922, 683, 924]. Sanborn [1309, 569]. Sancto [1185]. Sanger [1009]. Sanguinetti [510]. Sanskrit [996]. Santillana [544]. Santo [1017]. Sarton [1092, 521, 131]. Saturn [758, 795]. Saugus [325]. Saul [1215]. Saunders [194]. savant [453]. Save [247, 248]. Saved [245]. Saving [590]. Saxl [795, 120]. Saxon [1262, 678]. Sayili [273]. Sc.D [890]. Scale [704, 667]. Scene [93]. Scepticism [360]. Scetticismo [480]. Scheikunde [374]. Scheler [370]. Schelling [1358]. Schierbeek [1352, 164]. Schildberger [301]. Schiller [1192]. Schlegel [468]. Schlesinger [540]. Schlozer [626]. Schlunz [1149]. Schmitt [1305, 1344]. Schneer [214]. Schoeters [972]. Schofield [1160]. Scholarship [112]. Scholastic [30]. Scholz [361]. School [767, 451]. Schouten [1263]. Schramm [450, 651]. Schriften [369, 587]. Schristen [369]. Schröder [1359]. Schrödinger [1165, 973, 671]. Schubert [1090]. Schule [857, 858]. Schumacher [334]. Schuman [395]. Schwann [419, 419]. Schwarz [74]. Schweiz [122]. Schwenckfeld [99]. Schwinge [156]. Sciascia [332]. Science [472, 246, 259, 340, 92, 538, 1315, 921, 976, 1033, 997, 784, 712, 498, 256, 564, 140, 869, 121, 1354, 1023, 1311, 182, 1102, 793, 1078, 1012, 268, 1202, 172, 1040, 257, 797, 241, 159, 1227, 1042, 756, 1170, 84, 220, 658, 820, 934, 1096, 426, 1224, 1364, 889, 955, 1127, 1147, 455, 286, 1048, 271, 566, 131, 545, 710, 174, 344, 106, 700, 1207, 703, 565, 1303, 149, 1029, 578, 170, 113, 68, 1215, 45, 385, 204, 378, 237, 517, 547, 454, 1211, 395, 712, 1147, 271, 467, 818, 958, 119, 359, 708, 13, 922, 757]. Science [536, 132, 1058, 1298, 1050, 215, 319, 227, 343, 270, 193, 290, 451, 155, 956, 957, 320, 1206, 611, 14, 610, 923, 1210, 136, 998, 1316, 876, 999, 1147, 735, 612, 994]. Sciences [388, 1246, 121, 136, 537, 1178, 214, 248, 551, 624, 954, 169, 454, 1274, 1047, 540, 673, 271, 710]. Scientiarum [1013, 1014]. Scientific [640, 318, 15, 750, 932, 890, 134, 679, 729, 1254, 184, 697, 809, 888, 712, 674, 757, 604, 544, 244, 91, 1235, 441, 250, 242, 1251, 165, 646, 1091, 1298, 260, 439, 215, 532, 1048, 1208, 653, 1160, 1300, 1252, 412, 288, 672, 377, 1145, 672]. scientifica [1357, 1013]. scientifique [1235]. scientifiques [1278, 728, 1192].

Scientist [688, 141, 1245, 256, 970, 305, 240, 1320, 1188, 1086, 471]. Scientists [254, 1292, 1328, 657, 688, 481]. Scot [907]. Scotland [1231]. Scott [1217, 1165, 60]. Screen [921]. Sea [139, 1079, 717, 1017, 609]. Search [558, 214, 1335, 590, 62]. Seb [99]. Sebba [796]. Second [1230, 1285, 1096, 75]. Sections [342, 827]. Secundus [1069]. Sedrakovich [856]. Seen [1129]. Seidlerova [537]. Sein [76, 1213]. Seine [1213, 584, 76, 450]. seinem [814]. seiner [627]. seines [57]. seizième [286]. Selected [421, 1114, 564, 884, 1180, 200, 455, 1160, 904]. Selection [65, 771, 412, 54, 639, 54]. Selections [957, 840, 1343, 512]. Seligman [507]. Semantic [603]. Semmelweis [858]. Sen [996]. Sensationalism [640]. Sense [797]. Sensitivity [824]. Sensory [3]. September [268, 35, 166, 831]. Serene [732]. Serge [1242]. Serial [438]. Serie [624]. Series [992]. Serpents [1179]. ses [1267, 482, 728, 100, 511]. Seters [482]. Settlements [1337]. Seventeenth [30, 29, 931, 645, 1036, 1228, 703, 42, 472]. Seventeenth-Century [30, 1036, 1228, 703]. Seventeeth [824]. Seventh [378, 1102]. Seventh-Century [1102]. Several [928]. Seyyed [1210]. Shafranovskii [591]. Shall [247, 248]. Shamos [362]. Shaped [1183]. Shaping [300]. Shapley [922, 346]. Sharlin [617]. Sharp [324]. Sharrock [2]. Shatir [354, 865]. Shaw [884]. Sher [818]. Sherrington [1217]. Sherwood [889, 1054, 971]. Shetler [1176]. Shiou [931]. Shiou-chuan [931]. Ships [22]. Shook [1040]. Shore [1177]. Short [804, 721, 718, 476, 766, 137]. Shrubs [150]. Shryock [1266, 1004, 221]. Sibyl [1270]. Sidereal [852]. Sidney [1207, 270]. siècle [99, 551, 286, 547, 365, 454]. Siècles [223, 1267, 1336, 1081]. Siegel [191]. Siena [792]. Sigerist [1219, 222, 1219, 508]. Siglo [58, 445]. Siglos [58]. Sigmund [1021, 1055, 306, 653]. Significance [424, 746]. Silver [921, 743]. Silvio [809]. Simon [1190, 416, 656]. Simultaneous [1200]. Since [320, 932]. Singer [95, 64, 108, 475]. Sir [571, 105, 1133, 868, 526, 895, 414]. Sirak [1102]. Sirks [675]. Sitterson [420]. Six [968]. Six-Cornered [968]. Sixteenth [1157, 1343, 1307, 645, 1030]. Sixteenth-Century [1343, 1307, 1030]. Sixth [204]. Skelton [714, 17, 927]. Sketch [342]. Skies [613]. Skolnikoff [1147]. Slotkin [845]. Small [588]. Smeaton [370, 682]. Smeur [329]. Smirnov [1013, 1014]. Smith [890, 1255, 271, 623, 890, 279]. Smithsonian [359]. Smoke [492]. Smolka [537]. Smollett [991]. Snowflake [114, 968]. Sobre [398]. Social [93, 220, 477, 952, 803, 672]. Société [1181]. Societies [1251]. Society [722, 221, 1354, 333, 1140, 683, 848, 999, 1144, 1029, 1302, 358, 662, 712, 338, 998, 1316, 928]. Sociological [757]. Sociology [246, 177, 222, 1302]. Sodium [428]. Sohn [1292]. Soil [719]. solaires [1074]. Solander [1310, 526, 1031]. Solar [570]. solchen [627]. Solla [320, 407]. Solmsen [405]. Some [424, 631, 9, 251, 208, 742, 150, 241, 199, 868, 653, 987, 170]. someone [148, 663]. Sommerfeld [1063]. Somnium [1072]. Son [453, 1211, 1192, 409, 511]. Sonnedecker [1005, 908]. Sonnen [1115]. Sonnen- [1115]. Sons [1277]. Soto [1199]. Soul [650]. Sound [522]. Source [1054, 1301, 1247, 346]. Sources [1135, 637, 177, 1124, 1128, 1040]. sous [121]. South [1017, 932]. Soviet [1246, 516, 523, 340]. Sowjetische [397]. Soziale [122]. Space [1206, 784, 1228, 613, 641, 1022, 978, 1009, 1164]. Spain [908]. Spallanzani [3]. Spanish [531, 390]. Spath [808]. Special [908]. Species [312, 574, 1171, 50, 684]. Spectrum [701]. Speculation [1229, 704, 667]. Speeches [957]. Speed [522, 344]. Spencer [628]. Spink [294]. Spiritibus [369]. Spiritual [287]. spirituels [99]. Splendid [529]. splendidissimi [1264]. Spots [165]. Sprague [418, 616, 725]. Springs [915, 1189]. SSSR [216]. St [808, 723, 375]. Staff [784]. Stages [5]. Stahl [1325, 545]. Stahlman [570]. Stallo [303, 1326, 1326]. Standards [1254]. Stanislaw [1236]. Stanley [279, 1057]. Stanton [165]. Stanwyn [1176]. Star [87, 849]. State [885, 182, 426, 1018, 833, 1300]. Statement [825]. States [783, 341, 505, 1284, 1147, 736, 999]. Statesman [290]. Statesmen [1050]. Statistiker [198]. Steacie [957]. Steam [82]. Stearn [905]. Stearns [410]. Steel [815, 27, 1076, 579]. Steele [1019, 625]. Steere [19]. Stefansson [389]. Stein [1133]. stellae [1264]. Stellung [488]. Steno [930]. Stephen [469, 472, 1355, 842, 539]. Steps [1298]. Sterling [97]. Stern [1054, 306, 306]. Sternfeld [1020]. Sternuhren [1115]. Steudel [1041]. Steuer [194]. Stevenson [1303, 158, 124]. Stevin [416]. Sticker [1148]. Stieb [1005]. Still [601, 5]. Stillman [1159, 926, 292, 1343]. Stirrups [41]. Stoics [67]. Stones [548]. Storia [502]. Story [1279, 218, 326, 1040, 125, 762, 62, 630]. Stoudemire [97]. Strabo [1121]. Strange [432]. Strasbourg [561]. Strategy [690]. Straus [51]. Stroeniya [298]. Struck [394]. Structural [1077, 523, 1229]. Structure [1077, 441, 698, 298, 487, 1042, 215]. Stuart [28]. Stubbe [614]. Studie [1040]. Studien [649]. Studies [483, 202, 501, 795, 672, 996, 1358, 797, 676, 653, 63]. studio [480]. Studium [1359]. Study [956, 722, 230, 116, 546, 1033, 910, 769, 242, 813, 1283, 475, 715, 507, 1048, 1272, 678, 217, 710, 227, 226, 572, 414, 1263, 1082, 73]. Sturtevant [879]. Subject [252]. sublunari [875]. Sudhoff [369]. Suggestions [683]. suivie [511]. Sullivan [300]. Summa [528]. Summulae [635]. Sun [931, 931]. Sung [931]. Suns [34]. Sunspot [1109]. Suppes [467]. Supplement [854]. Support [182]. Supposed [980]. Surgeons [277]. Surgery [1267, 847]. Survey [1315, 954, 426]. Surveyor [139]. Surveys [182]. Susskind [648]. Suter [332]. Sutton [1149]. Suzanne [409]. Svehla [1334]. Swain [889]. Swammerdam [1352]. Swenson [1060]. Sydney [845]. Sylphis [369]. Symbol [1263, 855]. Symposium [14, 1120]. Syntax [905]. Synthese [1211]. Sypher [1297]. System [433, 463, 986, 836, 393, 837, 405]. Systeme [284]. Systems [1159]. Szabadváry [1334].

T [804, 30, 972, 1009, 773, 399, 1165, 476, 130, 1075, 1179, 645, 954, 905]. T. [141]. Tablas [390, 583]. tableaux [482]. Tables [354, 1261, 390]. talajtani [719]. Talbot [1153]. Tania [306]. Tanquam [209]. Tantric [1112, 1034]. Tarski [467]. Tartaglia [1343, 196]. Tasmania [1251]. Tassel [999]. Tat [443]. Taton [121, 1023, 454]. Taylor [959]. te [374]. Technical [932, 439, 672]. Technicum [1047]. technique [576]. Techniques [1337, 1214, 576, 724]. Technocrats [952]. Technological [230, 439, 746]. Technology [784, 27, 923, 676, 476, 129, 477, 1147, 590, 931, 95, 581, 578, 1122, 63, 994, 1141, 1009, 538, 64, 756, 1297, 1338, 1082, 1140, 1141]. Teich [537]. Teil [369, 160, 334, 364]. Teilhard [203]. Telegraph [617]. Telegraphe [1052]. Temin [815]. Temkin [1271, 51, 1271, 1303]. Temperature [704, 667]. temps [1192]. Tenth [793]. Teorii [298]. Teplov [334]. Terence [1355, 516]. terminologie [66]. Terminology [905]. Terms [603]. terre [624]. Test [914]. Testament [440]. Testimonianze [727]. Testo [847]. Texas [1353]. Text [917]. Texten [323]. textes [372, 589, 69]. Texts [996, 367, 881, 849, 65]. Thâbit [6]. Thabit [509, 527, 827]. Thaddaeus [1264]. Thaumantias [1264]. Their [301, 809, 870, 1122]. Themes [653]. Theobald [1001]. Theodor [419, 1077, 759, 419, 552]. Theodore [579, 1164, 585]. Theodoric [134]. Theodorides [446]. Theologie [547]. Theologische [369]. Theology [549]. Theophile [733]. Theophrast [369]. Theophrastus [882, 369]. Theorem [527, 6, 827]. Theories [4, 303, 1214, 1229, 1272]. Theory [354, 665, 89, 636, 425, 557, 52, 598, 54, 491, 1288, 1282, 1101, 798, 38, 1063, 915, 897, 758, 298, 487, 759, 990, 867, 1028, 1042, 940, 39, 148, 865, 914, 751, 826, 939, 980, 694, 313, 387, 523, 964, 961, 1040, 1020, 572]. Thermal [1103]. Thermodynamics [1026, 1285]. Thermometer [1146]. Theutonicus [702]. thing [148, 663]. Things [1251, 475]. Thinking [720]. Third [1224]. Thirlkel [808]. Thirty [1040]. Thistle [1093]. Thomas [414, 1315, 1149, 1304, 1190, 808, 441, 1128, 1076, 1010, 868, 401, 612, 464, 436, 149, 229, 1188]. Thompson [1309, 229]. Thomson [305, 492]. Thoren [1298]. Thorndike [679, 830, 907]. Thought [1177, 544, 977, 294, 726, 215, 75, 94, 219, 498, 575]. Thousand [34]. Three [869, 131]. Thudichum [592]. Ti [1155]. Tibb [25]. Tibor [615]. Tides [665]. T’ien [931]. T’ien-Kung [931]. Time [1206, 1354, 954, 641, 470, 842, 468]. Timekeeping [1007]. Times [615, 1177, 1262, 476, 1140, 214, 782, 1006, 28, 1002]. Timpanaro [727]. Timpanaro-Cardini [727]. Tin [803]. Titian [415]. Tizard [817]. Today [1207]. Todestages [57]. Tom [1347, 1215]. Tome [624, 1181, 576, 454]. Tomo [1243]. Tomus [1013, 1014]. Tonnelat [121]. Tool [1360]. Tools [804]. Topographie [547]. Topsell [1179]. torlenete [719]. Torpie [818]. Torricelli [384]. tortenete [499]. Totok [1213]. Touching [732]. Toulmin [472, 842, 539, 469]. Tower [307]. Towers [327]. Traces [1177]. Trachtenberg [855]. Tracks [1133]. Trade [1227]. Tradition [269, 1298, 622, 965, 652]. Traditional [1122]. traite [404]. Tralles [217]. Transformation [42]. Transits [227, 285]. Translated [973]. Translation [458, 917, 901, 771]. Translations [996, 983, 133]. translationum [133]. Translator [585]. Transmission [1064]. Transmit [1035]. Transmutation [780]. Traveler [1307]. Traveling [648]. Traveling-Wave [648]. Travels [229, 493, 510, 479]. travers [1336]. Treasury [713, 359, 922]. Treatise [592, 827]. Treatises [183]. Treatment [1286]. Tree [787]. Trees [150]. Trends [177]. Treneer [571]. Trevor [476, 95]. Triangles [1187]. triangulis [1187]. Troubled [1137]. Trough [1287]. True [359]. Truesdell [371, 295]. Tsander [1009]. Tsiolkovsky [1009]. tu [931]. Tube [648]. Tuge [578]. Tuke [683]. Tullio [480]. Tulp [73]. Tunnel [664, 1107]. Turn [1040]. Turnbull [161, 162, 513]. Turner [1087]. turquet [562]. Tuscan [652]. Tuscany [732]. Twelfth [697]. Twelfth-Century [697]. Twentieth [1023, 1141]. Two [1159, 628, 413, 430, 493, 870, 710, 435, 767]. Tycho [979, 1173, 939]. Tylecote [580]. tzu [1123, 1294].

U [688, 848]. U.S [1360, 1315]. U.S.A. [876]. Ubaldo [1343]. übertragen [369]. Uighur [637]. uit [1092]. Ulam [1236]. Ulisse [621]. umbrales [445]. Understanding [469]. Undertakers [325]. Unexposed [872]. Uniformitarian [7]. Uniformiter [1199]. Uniformly [1253]. United [783, 341, 505, 999, 1147, 736, 1284]. Unity [707]. Universal [1047]. Universe [1086, 762, 962, 698, 789]. université [728]. University [1266, 767, 1358, 268, 1340, 1075, 412, 979, 696]. Unknown [8]. Unnoticed [462]. Unpublished [151, 3, 264, 461, 742, 412]. Unusual [532]. upon [1012, 694]. Uppsala [969]. Uqlidisi [918]. Urban [1270]. Urdang [110]. Urness [1190]. URSS [1013, 1014]. Uschmann [276]. Use [818, 983, 601]. Uses [1146].

V [472, 368, 774, 654, 888, 712, 216, 298, 487, 417, 369, 512, 95, 1013, 1014, 373, 591]. Vagum [991]. Valency [844]. Valentin [531]. Valjavec [296]. Valley [734, 877]. Vallicrosa [583, 398]. Value [1136]. Values [923]. Vancouver [139]. Varia [1340]. Variants [491]. varietate [1105]. Various [710]. Vartanian [197]. Vassili [521]. Vault [546]. Ve [1081]. Vector [981]. vegetaux [1191]. Vegetti [807]. Veith [1155]. Velluz [781]. Ventures [325]. Venus [227]. Ver [47]. Verdonk [1308]. Verdun [562]. Vererbungsregeln [614]. Vergil [894]. Vermeer [1305]. Vern [1266]. vero [1264]. Version [315]. versione [847]. Versions [430]. versus [983, 1170]. Vesale [511]. Vesaliu [138]. Vesalius [753, 511, 847]. Vestiges [33]. veterinaire [1084]. Veteris [880]. vetustiorls [1085]. VI [481, 369, 63]. Vibration [915]. Vibrations [434]. Victor [1298]. Victore [1185]. Victorian [995, 1133, 1227]. VIe [547, 733, 100, 1211, 482, 1191, 409, 551, 781]. View [247, 254, 523, 1354, 908, 973]. Vieweg [1292]. Viewpoint [239]. Viewpoints [246]. Views [954, 1029, 1163]. VIII [1011, 8]. Vilhjalmur [389]. Villanova [941]. Vincent [271]. Vinchi [368]. Vinci [615, 1086, 743]. Virchow [200]. Virgil’s [925]. Virtanen [304]. Virtue [1320]. Virville [121]. Vis [741, 1132]. Visibilities [668]. Visibility [89, 39]. Vision [906, 1028, 321, 1297]. Visscher [1162]. Visual [720]. Vital [1077]. Vitalism [977]. Vitalist [1282, 377]. Vitalist-Mechanist [1282]. Vitruvian [1136]. viva [1132, 741]. Vivian [355]. Vocabulary [905]. Vogel [772, 1184]. Voiages [927]. Voices [954]. Void [597, 1228, 1030]. Vol [1009, 370, 121, 676, 646, 1140, 1178, 541, 1042, 1309, 1143, 510, 849, 95, 133, 1085, 513, 994, 1141, 621]. Volcanological [786]. Vollgraff [829]. Voltaire [294]. Volume [143, 232, 379, 518, 659, 821, 935, 1095, 1222, 1362, 722, 367, 72, 161, 1206, 711, 677, 162, 504, 136, 731, 216, 481, 715, 508, 565, 479, 63, 454, 416, 965]. Volumen [1014]. Vom [369, 484, 847, 449]. voor [374]. Vorstellungen [1148, 364]. Voyage [1017, 263, 763, 1031]. Voyages [139]. Vrienden [1092]. Vucinich [610]. Vuillemin [297]. VV [373].

W [688, 1145, 957, 767, 631, 92, 589, 1008, 1149, 817, 72, 161, 1358, 717, 482, 303, 650, 201, 1337, 677, 844, 1089, 808, 652, 370, 162, 642, 1005, 1140, 1146, 375, 766, 592, 28, 327, 889, 952, 1125, 656, 1043, 513, 1001, 1141]. wa [967]. Wadie [135]. Waerden [1040]. Wagoner [1360]. Walahfrid [1121]. Waley [64]. Walker [287]. Wallace [54, 847, 134, 1289, 813]. Waller [732]. Wallis [827]. Walser [1161]. Walter [995, 1114, 1012, 224, 584, 1339, 321, 649, 122, 559, 1286]. Walusinski [121]. Wanda [547]. Wanderer [1244]. Wandlung [443]. Wang [136, 407, 994]. Wants [245]. War [1202]. Ward [330]. Ware [1123]. Warner [1277]. Was [664, 1295]. Wasfi [478]. Watanabe [611]. Watchers [613]. Water [1149]. Watermann [419]. Waters [1137, 28]. Watson [1220, 1238]. Wave [648, 1295, 524]. Wave-Particle [1295]. Waves [746]. Weapons [690]. Weather [832, 341, 831]. Weber [1193]. Webers [488]. Webster [359]. Weeks [1237]. Weg [76, 651]. Wegbereiter [1156, 814]. Weigel [99]. Weight [1257]. Weight-Driven [1257]. Weights [1262]. Weimann [369]. Weiner [1245]. Weinstock [763]. Weisheit [369, 649]. Weisskopf [888]. Wellcome [542, 1143, 565]. Wellenbewegung [1193]. Wellenmechanik [671]. Wellesley [995]. Welt [1213, 55, 156]. Weltalls [1148]. Weltansicht [973]. Weltbild [584]. Weltbildes [443]. Weltgeschichte [296]. Wen [1155]. Werk [1092, 552, 419, 814]. Werke [70, 587, 810, 369, 98]. Werkstatt [156]. Werner [417, 116, 946]. Wertime [579]. Wesley [1208]. West [97, 1028, 801, 1076, 1169, 225, 970]. Western [1177, 219, 498, 64, 269, 1140, 629, 700, 565, 1122, 1141, 385]. Westfall [1298]. Westward [1064]. Wheatland [1311]. Wheeler [399]. Wheelwright [283]. White [1149, 477, 1152]. Whitehead [455]. Whitfield [72]. Whitnah [341]. Whitrow [470]. Whitteridge [336]. Who [1126]. Whyte [483, 442]. Wickersheimer [1154]. Wiederentdeckung [614]. Wiederkehr [57, 488, 1193]. Wiener [857, 858]. Wilderness [1244]. Wilfred [678]. Wilfried [1041]. Wilhelm [1213, 847, 1090, 488, 1193]. Wilkes [415]. Willard [300]. William [336, 134, 911, 30, 102, 19, 904, 682, 399, 1165, 970, 165, 570, 720, 876, 1048, 545, 1087, 73, 1113, 572, 905, 51, 1325, 666, 492, 496, 863, 644, 1114, 643]. Williams [344, 816, 961, 476, 95]. Willier [800]. Willis [1188]. Willy [613, 1091]. Wilma [813]. Wilson [1244, 963]. Winch [220]. Wind [1164]. Winning [970]. Winter [160, 334, 626, 1269, 929, 32]. Wirceburgensis [548]. Wirken [1213]. Wisconsin [268]. Wiskunde [1308]. Wissenschaft [1348, 1359, 195, 32, 649]. Wissenschaften [160, 334].

Wissenschaftsbeziehungen [626]. Witch [1010]. Wöhlers [74]. Wolf [685]. Wolff [1278]. Wolfgang [888]. Wolfram [1305]. Wolska [547]. Woman [989]. Wong [223]. Woodbridge [906]. Woodbury [127, 127, 230, 305]. Woolf [712, 548, 227]. Wooster [1313]. Worcester [1018]. Word [701, 988]. Work [1070, 557, 483, 773, 399, 896, 985, 1072, 199, 1129, 946, 943, 1252, 1019, 1124, 288]. Works [996, 184, 697, 1009, 509, 150, 1180, 505, 1309, 854, 812, 164, 416, 1009, 22, 1352]. World [1159, 2, 688, 52, 337, 619, 763, 960, 158, 1202, 1040, 31, 164, 973, 778, 229, 848, 1152, 405, 468]. Worthies [750]. Wright [723, 922, 962]. Wright-St [723]. Writers [1002]. Writing [818]. Writings [679, 1219, 342, 813, 1350, 653, 840, 1219, 1165, 512]. Written [565]. Wu [931]. Wulff [1122]. Wyckoff [1273]. Wylie [1297].

X [70, 624]. Xerxes [392]. XII [264]. Ximenes [530]. XIV [621]. XIX [373, 58]. XIXe [454]. XVe [1081]. XVIe [1267, 99]. XVII [58]. XVIIe [1267]. XVIII [58]. XVIIIe [1267, 551]. XX [373, 445]. XXXIX [1178].

Yūnis [1327]. Yale [376, 149]. Yaqut [135]. Yaqzan [69]. Yates [622]. Year [411, 475, 114]. Years [709, 801, 1176, 19, 570, 1040, 736, 1326]. Yellow [1155]. Ying [931]. Ying-hsing [931]. York [1056, 683, 846]. Young [151, 933]. Ypocras [770]. Yvonne [1084].

Z [537, 57]. Zachary [277]. Zahrawi [908]. Zakariyya [25]. Zdenek [1264]. Zeitgenossen [627]. Zemplen [499]. Zen [931]. Zenith [314]. zenonischen [450]. Zestiende [329]. Zestiende-eeuwse [329]. Zeugnissen [627]. Zijn [1092]. Zimmermann [1306]. Zirkle [93, 675]. Zofia [225]. Zoologie [276, 1045, 446]. Zoologischen [276]. Zoologist [572]. Zoubov [521]. Zubov [368]. Zuckerman [564]. zum [1359, 449, 484]. Zupko [1262]. Zur [449, 1348, 195, 671, 397, 1040, 1041, 614, 366, 651]. Zusammen [584]. Zusammenarbeit [1359]. zwei [1148]. Zweite [369].

References

Anonymous:1960:FMa

[1] Anonymous. Front matter. Isis, 51(1):1–2, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/227598.

Arber:1960:RSP

[2] Agnes Arber. Robert Sharrock (1630–1684): A precursor of Nehemiah Grew (1641–1712) and an exponent of “natural law” in the plant world. Isis, 51(1):3–8, March 1960. CODEN ISISA4. ISSN 0021-1753 (print),

      1545-6994 (electronic). URL http://www.jstor.org/stable/227599.

Dijkgraaf:1960:SUE

[3] Sven Dijkgraaf. Spallanzani’s unpublished experiments on the sensory basis of object perception in bats. Isis, 51(1):9–20, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227600.

Beer:1960:ECT

[4] John J. Beer. Eighteenth-Century theories on the process of dyeing. Isis, 51(1):21–30, March 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227601.

Levey:1960:ESE

[5] Martin Levey. The earliest stages in the evolution of the still. Isis, 51(1): 31–34, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227602.

Sayili:1960:TIQ

[6] Aydin Sayili. Thâbit ibn Qurra’s generalization of the Pythagorean Theorem. Isis, 51(1):35–37, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227603.

Cannon:1960:UCD

[7] Walter F. Cannon. The uniformitarian–catastrophist debate. Isis, 51(1): 38–55, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227604.

Drake:1960:GGV

[8] Stillman Drake. Galileo gleanings VIII: The origin of Galileo’s book on floating bodies and the question of the unknown academician. Isis, 51(1): 56–63, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227605.

Duveen:1960:NSL

[9] Denis Duveen and Roger Hahn. A note on some Lavoisiereana in the “Journal de Paris”. Isis, 51(1):64–66, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227606.

Larkin:1960:STA

[10] Vincent R. Larkin and Saint Thomas Aquinas. Saint Thomas Aquinas: “On The Combining of the Elements”. Isis, 51(1):67–72, March 1960.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227607.

Koestler:1960:NC

[11] Arthur Koestler, Giorgio de Santillana, Stillman Drake, L. A. Moritz, N. Jasny, Frank M. Albrecht, P. H. Brans, James D. Mack, and Roy G. Neville. Notes & correspondence. Isis, 51(1):73–84, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227608.

Anonymous:1960:NPa

[12] Anonymous. News of the profession. Isis, 51(1):85–87, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227609.

Ducasse:1960:BRB

[13] C. J. Ducasse. Book review: Causality. The Place of the Causal Principle in Modern Science by Mario Bunge. Isis, 51(1):88–90, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227610.

Dingle:1960:BRB

[14] Herbert Dingle. Book review: Science and the Modern Mind: A Symposium by Gerald Holton. Isis, 51(1):90–91, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227611.

Bar-Hillel:1960:BRB

[15] Y. Bar-Hillel and S. Sambursky. Book review: The Logic of Scientific Discovery by Karl R. Popper. Isis, 51(1):91–94, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227612.

Taub:1960:BRB

[16] A. H. Taub. Book review: The Computer and the Brain by John von Neumann. Isis, 51(1):94–96, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227613.

Kish:1960:BRB

[17] George Kish. Book review: Explorers’ Maps. Chapters in the Cartographic Record of Geographical Discovery by R. A. Skelton. Isis, 51(1):

96–97, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227614.

Zirkle:1960:BRBa

[18] Conway Zirkle. Book review: Nature and Man’s Fate by Garrett Hardin. Isis, 51(1):97–98, March 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227615.

Gasking:1960:BRB

[19] Elizabeth B. Gasking. Book review: Fifty Years of Botany by William Campbell Steere. Isis, 51(1):98–99, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227616.

Brookhart:1960:BRB

[20] John M. Brookhart. Book review: The History and Philosophy of Knowledge of the Brain and Its Functions by F. N. L. Poynter. Isis, 51(1):99– 101, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227617.

Meyer:1960:BRB

[21] Karl F. Meyer. Book review: A History of Public Health by George Rosen. Isis, 51(1):101–102, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227618.

Burlingame:1960:BRB

[22] Roger Burlingame. Book review: Cradle of Ships. A History of the Bath Iron Works by Garnett Laidlaw Eskew. Isis, 51(1):102, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227619.

Usher:1960:BRB

[23] Abbott Payson Usher. Book review: Inventors’ Progress by Joachim G. Leithauser; Michael Bullock. Isis, 51(1):103–104, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227620.

Saunders:1960:BRB

[24] J. B. deC. M. Saunders. Book review: La Médecine Hippocratique by Charles Lichtenthaeler. Isis, 51(1):104–105, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227621.

Yahia:1960:BRB

[25] B. Ben Yahia. Book review: Kitabu’l Hawi Fi’t-Tibb by Abu Bakr Muhammad Ibn Zakariyya ar-Razi. Isis, 51(1):105–107, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227622.

Nasr:1960:BRB

[26] Seyyed Hossein Nasr. Book review: Kitabu’l-’anwa’ (On the Meteorology of the Arabs) by ’Abu Muhammad ’Abdallah B. Muslim Ibn Qutayaba Ad-Dinawari. Isis, 51(1):107–108, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227623.

Goodrich:1960:BRB

[27] L. Carrington Goodrich. Book review: The Development of Iron and Steel Technology in China by Joseph Needham. Isis, 51(1):108–109, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227624.

Morison:1960:BRB

[28] S. E. Morison. Book review: The Art of Navigation in England in Elizabethan and Early Stuart Times by D. W. Waters. Isis, 51(1):109–111, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227625.

Leicester:1960:BRBa

[29] Henry M. Leicester. Book review: Robert Boyle and Seventeenth Century Chemistry by Marie Boas. Isis, 51(1):111–112, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227626.

Curtis:1960:BRB

[30] Mark Curtis. Book review: The Scholastic Curriculum at Early Seventeenth-Century Cambridge by William T. Costello. Isis, 51(1):112– 113, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227627.

Schofield:1960:BRB

[31] Robert E. Schofield. Book review: The Medical World of the Eighteenth Century by Lester S. King. Isis, 51(1):113–115, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227628.

Truesdell:1960:BRB

[32] C. Truesdell. Book review: Die Deutsch-Russische Begegnung und Leonhard Euler. Beiträge zu den Beziehungenzwischen der deutschen und der russischen Wissenschaft und Kultur im 18.Jahrhundert by E. Winter. Isis, 51(1):115, March 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227629.

Greene:1960:BRB

[33] John C. Greene. Book review: Just before Darwin; Robert Chambers and Vestiges by Milton Millhauser. Isis, 51(1):116–117, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227630.

Anderson:1960:BRB

[34] Oscar E. Anderson, Jr. Book review: Brighter than a Thousand Suns by Robert Jungk; James Cleugh. Isis, 51(1):117–119, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227631.

Anonymous:1960:PRS

[35] Anonymous. Publications received September–December 1959. Isis, 51 (1):120–123, March 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227632.

Anonymous:1960:BMa

[36] Anonymous. Back matter. Isis, 51(1):124, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227633.

Anonymous:1960:FMb

[37] Anonymous. Front matter. Isis, 51(2):129–130, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226843.

Hall:1960:NTM

[38] A. Rupert Hall and Marie Boas Hall. Newton’s theory of matter. Isis, 51(2):131–144, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226844.

Middleton:1960:BLT

[39] W. E. Knowles Middleton. Bouguer, Lambert, and the theory of horizontal visibility. Isis, 51(2):145–149, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/ stable/226845. See correction [89].

Hanson:1960:MPE

[40] Norwood Russell Hanson. The mathematical power of epicyclical astronomy. Isis, 51(2):150–158, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226846.

deCamp:1960:BS

[41] L. Sprague de Camp. Before stirrups. Isis, 51(2):159–160, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226847.

Wilson:1960:TAC

[42] Leonard G. Wilson. The transformation of ancient concepts of respiration in the Seventeenth Century. Isis, 51(2):161–172, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226848.

Harrison:1960:BPP

[43] Thomas P. Harrison. Bird of paradise: Phoenix redivivus. Isis, 51(2): 173–180, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226849.

Thorndike:1960:QA

[44] Lynn Thorndike. Questiones Alani. Isis, 51(2):181–185, June 1960. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226850.

Wohl:1960:BHP

[45] Robert Wohl. Buffon and his project for a new science. Isis, 51(2):186– 199, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226851.

Popkin:1960:ERL

[46] Richard H. Popkin. Éloges: Robert Lenoble 1900–1959. Isis, 51(2):200– 202, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226852.

Rome:1960:EPL

[47] A. Rome. Eloge: Paul-Louis Ver Eecke. Isis, 51(2):202–203, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226853.

Drake:1960:NC

[48] Stillman Drake, Lynn Thorndike, Rufus Suter, C. Truesdell, and John B. Pierczynski. Notes & correspondence. Isis, 51(2):204–208, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226854.

Anonymous:1960:NPb

[49] Anonymous. News of the profession. Isis, 51(2):208–210, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226855.

Hatch:1960:BRB

[50] Melville H. Hatch. Book review: The Origin of Species by Charles Darwin; Morse Peckham. Isis, 51(2):211–212, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226856.

deTerra:1960:BRBa

[51] Helmut de Terra. Book review: Forerunners of Darwin: 1745–1859 by Bentley Glass; Owsei Temkin; William L. Straus. Isis, 51(2):212– 213, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226857.

Cannon:1960:BRB

[52] Walter F. Cannon. Book review: The Age of the World. Moses to Darwin by Francis C. Haber; Darwin and the General Reader. The Reception of Darwin’s Theory of Evolution in the British Periodical Press, 1859–1872 by Alvar Ellegard. Isis, 51(2):213–215, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226858.

Gillispie:1960:BRB

[53] Charles C. Gillispie. Book review: Darwin and the Darwinian Revolution by Gertrude Himmelfarb. Isis, 51(2):216, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226859.

Dupree:1960:BRBa

[54] A. Hunter Dupree. Book review: Darwin, Wallace, and the Theory of Natural Selection by Bert James Lowenberg; Charles Darwin: Evolution and Natural Selection by Charles Darwin; Bert James Loewenberg. Isis, 51(2):216–217, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226860.

deTerra:1960:BRBb

[55] Helmut de Terra. Book review: Alexander von Humboldt. Erschliesser einer neuen Welt by Ewald Banse. Isis, 51(2):217–218, June 1960. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226861.

Seide:1960:BRB

[56] J. Seide. Book review: Alexander von Humboldt by Hanno Beck. Isis, 51(2):218–219, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226862.

Wassermann:1960:BRB

[57] Felix M. Wassermann. Book review: Gedenkschrift z. 100. Wiederkehr seines Todestages by Alexander von Humboldt; Gesprache by A. von Humboldts. Isis, 51(2):220–221, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226863.

Guerra:1960:BRBa

[58] F. Guerra. Book review: Biología y Medicina en los Siglos XVII y XVIII by Desiderio Papp; José Babini; Las Ciencias Exactas en el Siglo XIX by Desiderio Papp; Jose Babini. Isis, 51(2):221–222, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226864.

Allen:1960:BRB

[59] D. C. Allen. Book review: Mountain Gloom and Mountain Glory: The Development of the Aesthetics of the Infinite by Marjorie Hope Nicolson. Isis, 51(2):222–223, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226865.

Jones:1960:BRB

[60] Phillip S. Jones. Book review: A History of Mathematics. From Antiquity to the Beginning of the Nineteenth Century by J. F. Scott. Isis, 51(2):

224–225, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226866.

Zirkle:1960:BRBb

[61] Conway Zirkle. Book review: Origin of Cultivated Plants by Alphonse de Candole. Isis, 51(2):225–226, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226867.

Sonnedecker:1960:BRBa

[62] Glenn Sonnedecker. Book review: The Eternal Search. The Story of Man and His Drugs by Richard Mathison. Isis, 51(2):226–227, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226868.

White:1960:BRB

[63] Lynn White, Jr. Book review: Studies in Ancient Technology. Volume VI by R. J. Forbes. Isis, 51(2):227–228, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226869.

Forbes:1960:BRBa

[64] R. J. Forbes. Book review: A History of Western Technology by Friedrich Klemm; Dorothea Waley Singer. Isis, 51(2):228–229, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226870.

Goldat:1960:BRB

[65] George D. Goldat. Book review: The Presocratic Philosophers. A Critical History with a Selection of Texts by G. S. Kirk; J. E. Raven. Isis, 51(2): 229–231, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226871.

Toomer:1960:BRB

[66] G. J. Toomer. Book review: Dictionnaire historique de la terminologie géométrique des Grecs by Charles Mugler. Isis, 51(2):231–233, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226872.

Reesor:1960:BRB

[67] Margaret E. Reesor. Book review: Physics of the Stoics by S. Sambursky. Isis, 51(2):233–234, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226873.

Wilson:1960:BRB

[68] Curtis Wilson. Book review: The Science of Mechanics in the Middle Ages by Marshall Clagett. Isis, 51(2):234–237, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226874.

Millas-Vallicrosa:1960:BRB

[69] J. M. Millás-Vallicrosa. Book review: Le Récit de Hayy ibn Yaqz.an commenté par des textes d’Avicenne by A. M. Goichon. Isis, 51(2):237– 238, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226875.

Forbes:1960:BRBb

[70] R. J. Forbes. Book review: Ausgewählte Werke. De natura fossilium libri X, die Mineralien by Georgius Agricola. Isis, 51(2):239, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226876.

Odlozilik:1960:BRB

[71] Otakar Odlozilik. Book review: Opera Didactica Omnia by Joannes Amos Comenius. Isis, 51(2):239–241, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226877.

Cohen:1960:BRBa

[72] I. Bernard Cohen. Book review: The Papers of Benjamin Franklin. Volume I by Benjamin Franklin; Leonard W. Labaree; Whitfield J. Bell; Helen C. Boatfield; Helene H. Fineman; Benjamin Franklin and Italy by Antonio Pace. Isis, 51(2):241–243, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226878.

Temkin:1960:BRB

[73] Owsei Temkin. Book review: Rembrandt’s Anatomy of Dr. Nicolaas Tulp. An Iconological Study by William S. Heckscher. Isis, 51(2):243– 244, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226879.

Hiebert:1960:BRB

[74] Erwin N. Hiebert. Book review: Aus Justus Liebigs und Friedrich Wöhlers Briefwechsel in den Jahren 1829–1873 by Robert Schwarz. Isis, 51(2):244–246, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226880.

Spitzer:1960:BRB

[75] Alan Spitzer. Book review: Positivist Thought in France during the Second Empire, 1852–1870 by D. G. Charlton. Isis, 51(2):246–248, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226881.

Ihde:1960:BRB

[76] Aaron Ihde. Book review: Martin Heinrich Klaproth 1743–1817. Ein deutscher Apotheker und Chemiker, Sein Weg und seine Leistung by Georg Edmund Dann. Isis, 51(2):248, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226882.

Henley:1960:BRB

[77] Ernest M. Henley. Book review: Max Planck als Mensch und Denker by Hans Hartmann. Isis, 51(2):249, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226883.

Anonymous:1960:PRD

[78] Anonymous. Publications received December 1959–March 1960. Isis, 51(2):250–251, June 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226884.

Hindle:1960:AD

[79] Brooke Hindle. Administrative documents. Isis, 51(2):252–254, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226885.

Anonymous:1960:BMb

[80] Anonymous. Back matter. Isis, 51(2):??, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226886.

Anonymous:1960:FMc

[81] Anonymous. Front matter. Isis, 51(3):255–256, September 1960. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226505.

Kerker:1960:SCS

[82] Milton Kerker. Sadi Carnot and the steam engine engineers. Isis, 51 (3):257–270, September 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226506.

Drabkin:1960:NGM

[83] I. E. Drabkin. A note on Galileo’s de motu. Isis, 51(3):271–277, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226507.

Mendelsohn:1960:JLH

[84] Everett Mendelsohn. John Lining and his contribution to early American science. Isis, 51(3):278–292, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226508.

Grant:1960:NOH

[85] Edward Grant. Nicole Oresme and his De Proportionibus Proportionum. Isis, 51(3):293–314, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226509.

Bockstaele:1960:NFA

[86] P. Bockstaele. Notes on the first arithmetics printed in Dutch and English. Isis, 51(3):315–321, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226510.

Hellman:1960:MLE

[87] C. Doris Hellman and Francesco Maurolyco. Maurolyco’s “lost” essay on the new star of 1572. Isis, 51(3):322–336, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226511.

Koyre:1960:NC

[88] Alexandre Koyré, I. Bernard Cohen, Stillman Drake, W. E. K. Middleton, and W. C. Zeek. Notes & correspondence. Isis, 51(3):337–342, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226512.

Anonymous:1960:EBL

[89] Anonymous. Erratum: Bouguer, Lambert, and the theory of horizontal visibility. Isis, 51(3):342, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226513. See [39].

Anonymous:1960:NPc

[90] Anonymous. News of the profession. Isis, 51(3):342–343, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226514.

Hall:1960:BRBa

[91] A. Rupert Hall. Book review: The Edge of Objectivity. An Essay in the History of Scientific Ideas by Charles Coulston Gillispie. Isis, 51 (3):344–347, September 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226515.

Boyer:1960:BRBa

[92] Carl B. Boyer. Book review: History and Philosophy of Science: An Introduction by L. W. H. Hull. Isis, 51(3):347–348, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226516.

Joravsky:1960:BRB

[93] David Joravsky. Book review: Evolution, Marxian Biology and the Social Scene by Conway Zirkle. Isis, 51(3):348–353, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226517.

Bodemer:1960:BRB

[94] Charles Bodemer. Book review: The Historical Development of Physiological Thought by Chandler McC. Brooks; Paul F. Cranefield. Isis, 51 (3):353–354, September 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226518.

Smith:1960:BRB

[95] Cyril Stanley Smith. Book review: A History of Technology. Vol. V. The Late Nineteenth Century, c. 1850–c. 1900 by Charles Singer; E. J. Holmyard; A. R. Hall; Trevor I. Williams. Isis, 51(3):354–356, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226519.

Plessner:1960:BRB

[96] M. Plessner. Book review: Das Buch der Gifte des Gabir Ibn Hayyan by Gabir Ibn Hayyan. Isis, 51(3):356–359, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226520.

Guerra:1960:BRBb

[97] F. Guerra. Book review: Natural History of the West Indies by Gonzalo Fernández de Oviedo; Sterling A. Stoudemire. Isis, 51(3):359, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226521.

Hellman:1960:BRB

[98] C. Doris Hellman. Book review: Gesammelte Werke by Johannes Kepler. Isis, 51(3):360–361, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226522.

Mosse:1960:BRB

[99] George L. Mosse. Book review: Mystiques, spirituels, alchimistes du XVIe siècle allemand. Schwenckfeld, Seb. Franck, Weigel, Paracelse by Alexandre Koyre. Isis, 51(3):361–362, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226523.

Hall:1960:BRBb

[100] Marie Boas Hall. Book review: La Vie et l’oeuvre de Lavoisier d’après ses écrits by R. Dujarric de la Rivière; Madeleine Chabrier. Isis, 51(3):363, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226524.

Doney:1960:BRB

[101] Willis Doney. Book review: Les Premières pensées de Descartes by Henri Gouhier. Isis, 51(3):363–365, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226525.

Dibner:1960:BRB

[102] Bern Dibner. Book review: The De Magnete of William Gilbert by Duane H. D. Roller. Isis, 51(3):365–366, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226526.

Leicester:1960:BRBb

[103] Henry M. Leicester. Book review: The Emergence of the German Dye Industry to 1925 by John Joseph Beer. Isis, 51(3):366–367, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226527.

Oliver:1960:BRB

[104] John W. Oliver. Book review: The Atlantic Cable by Bern Dibner. Isis, 51(3):367–369, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226528.

Guthrie:1960:BRB

[105] Douglas Guthrie. Book review: The Life of Sir Alexander Fleming. Discoverer of Penicillin by Andre Maurois; Gerard Hopkins. Isis, 51 (3):369–370, September 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226529.

Sivin:1960:EFC

[106] Nathan Sivin, Harry Woolf, and Phyllis Brooks Bosson. Eighty-fifth critical bibliography of the history of science and its cultural influences (to 1 January 1960). Isis, 51(3):371–484, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226530.

Anonymous:1960:BMc

[107] Anonymous. Back matter. Isis, 51(3):??, September 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226531.

Hall:1960:ECJ

[108] A. Rupert Hall. Éloges: Charles Joseph Singer, 1876–1960. Isis, 51(4): 486+558–560, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228614.

Leake:1960:EJF

[109] Chauncey D. Leake. Eloge: John Farquhar Fulton, 1899–1960. Isis, 51 (4):486+560–562, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228615.

Sonnedecker:1960:EGU

[110] Glenn Sonnedecker. Eloge: George Urdang, 1882–1960. Isis, 51(4): 486+562–564, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228616.

Anonymous:1960:FMd

[111] Anonymous. Front matter. Isis, 51(4):487–488, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228608.

Cohen:1960:NLR

[112] I. Bernard Cohen. Newton in the light of recent scholarship. Isis, 51 (4):489–514, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228609.

Williams:1960:MFE

[113] L. Pearce Williams. Michael Faraday’s education in science. Isis, 51 (4):515–530, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228610.

Schneer:1960:KNY

[114] Cecil Schneer. Kepler’s new Year’s gift of a snowflake. Isis, 51(4):531– 545, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228611.

Herivel:1960:NDL

[115] J. W. Herivel. Newton’s discovery of the law of centrifugal force. Isis, 51 (4):546–553, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228612.

Carozzi:1960:SWP

[116] Albert V. Carozzi. A study of Werner’s personal copy of Von den äusserlichen Kennzeichen der Fossilien, 1774. Isis, 51(4):554–557, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228613.

Aaboe:1960:NC

[117] Asgér Aaboe, Rufus Suter, Charles Coulston Gillispie, and B. L. van der Waerden. Notes & correspondence. Isis, 51(4):565–568, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228617.

Anonymous:1960:NPd

[118] Anonymous. News of the profession. Isis, 51(4):568–570, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228618.

Cohen:1960:BRBb

[119] I. Bernard Cohen. Book review: Men and Moments in the History of Science by Herbert M. Evans. Isis, 51(4):571–572, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228619.

Bober:1960:BRB

[120] Harry Bober. Book review: Lectures by Fritz Saxl. Isis, 51(4):572–574, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228620.

Hall:1960:BRBc

[121] Marie Boas Hall. Book review: Histoire Générale des Sciences, publiée sous la direction de René Taton. Vol. II, La Science moderne (de 1450 à 1800) by G. Allard; E. Bauer; G. Canguilhem; J. Chesneaux; I. B. Cohen; P. Costabel; M. Daumas; A. Davy de Virville; P. Delaunay; R. Dugas; L. Dulieu; J. Filliozat; R. Furon; É. Guyénot; J. Itard; A. Koyré; R. Lenoble; J. Lévy; Ch. Morazé; J. Needham; J. Rostand; J. Taton; R. Taton; M.-A. Tonnelat; G. Walusinski. Isis, 51(4):574–576, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228621.

Barber:1960:BRB

[122] Elinor G. Barber. Book review: Soziale Grundlagen der Entwicklung der Naturwissenschaften in der alten Schweiz by Emil J. Walter. Isis, 51 (4):576–578, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228622.

Drake:1960:BRB

[123] Stillman Drake. Book review: The Birth of a New Physics by I. Bernard Cohen. Isis, 51(4):578, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228623.

Sonnedecker:1960:BRBb

[124] Glenn Sonnedecker. Book review: The Meaning of Poison by Lloyd G. Stevenson. Isis, 51(4):578–579, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228624.

Meier:1960:BRB

[125] Paul Meier. Book review: The Story of Chemical Industry in Basle. Isis, 51(4):579–580, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228625.

Kirby:1960:BRB

[126] Richard S. Kirby. Book review: The Canal Builders by Robert Payne. Isis, 51(4):580–582, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228626.

Battison:1960:BRB

[127] Edwin A. Battison. Book review: History of the Gear-Cutting Machine by Robert S. Woodbury; History of the Grinding Machine by Robert S. Woodbury. Isis, 51(4):582–583, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228627.

Steuer:1960:BRB

[128] Robert O. Steuer. Book review: Essai sur la médicine égyptienne de l’époque pharaonique by Gustave Lefebvre. Isis, 51(4):583–587, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228628.

Leicester:1960:BRBc

[129] Henry M. Leicester. Book review: Chemistry and Chemical Technology in Ancient Mesopotamia by Martin Levey. Isis, 51(4):587–588, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228629.

Levy:1960:BRB

[130] Jaroslav Levy. Book review: The Greek Herbal of Dioscorides by Robert T. Gunther. Isis, 51(4):588–590, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228630.

Sambursky:1960:BRB

[131] S. Sambursky. Book review: A History of Science; Hellenistic Science and Culture in the Last Three Centuries B.C. by George Sarton. Isis, 51 (4):590–591, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228631.

Grant:1960:BRB

[132] Edward Grant. Book review: Medieval and Early Modern Science by A. C. Crombie. Isis, 51(4):591–593, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228632.

Thorndike:1960:BRB

[133] Lynn Thorndike. Book review: Catalogus translationum et commentariorum: Mediaeval and Renaissance Latin Translations and Commentaries. Annotated Lists and Guides. Vol. I by Paul Oskar Kristeller. Isis, 51 (4):593–594, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228633.

Boyer:1960:BRBb

[134] Carl B. Boyer. Book review: The Scientific Methodology of Theodoric of Freiberg by William A. Wallace. Isis, 51(4):594–596, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228634.

Hourani:1960:BRB

[135] George F. Hourani. Book review: The Introductory Chapters of Yaqut’s Mu‘jam al-buldan by Wadie Jwaideh. Isis, 51(4):596–598, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228635.

Huang:1960:BRB

[136] Su-Shu Huang. Book review: Science and Civilisation in China. Volume 3. Mathematics and the Sciences of the Heavens and the Earth by Joseph Needham; Wang Ling. Isis, 51(4):598–600, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228636.

OMalley:1960:BRBa

[137] C. D. O’Malley. Book review: A Short Introduction to Anatomy (Isagogae Breves) by Jacopo Berengario da Carpi; L. R. Lind; Paul G. Roofe. Isis, 51(4):600–602, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228637.

OMalley:1960:BRBb

[138] Charles D. O’Malley. Book review: Andreas Vesalius’ First Public Anatomy at Bologna. 1540 by Baldasar Heseler; Ruben Eriksson. Isis, 51 (4):602–603, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228638.

Beaglehole:1960:BRB

[139] J. C. Beaglehole. Book review: Surveyor of the Sea; The Life and Voyages of Captain George Vancouver by Bern Anderson. Isis, 51(4):603– 605, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228639.

Hahn:1960:BRB

[140] Roger Hahn. Book review: La Révolution Française et la Science 1789– 1795 by Joseph Fayet. Isis, 51(4):605–607, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228640.

Dupree:1960:BRBb

[141] A. Hunter Dupree. Book review: T. H. Huxley–Scientist, Humanist and Educator by Cyril Bibby. Isis, 51(4):607–608, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228641.

Anonymous:1960:PRA

[142] Anonymous. Publications received April 1960–August 1960. Isis, 51 (4):609–611, December 1960. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228642.

Anonymous:1960:VI

[143] Anonymous. Volume information. Isis, 51(4):612–616, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228607.

Anonymous:1960:BMd

[144] Anonymous. Back matter. Isis, 51(4):617–618, December 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228643.

Anonymous:1961:FMa

[145] Anonymous. Front matter. Isis, 52(1):1–2, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228337.

Guerlac:1961:ARP

[146] Henry Guerlac. Address of retiring President. Isis, 52(1):3–6, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228338.

Ravetz:1961:RPQ

[147] J. Ravetz. The representation of physical quantities in Eighteenth-Century mathematical physics. Isis, 52(1):7–20, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228339.

Pines:1961:OQM

[148] S. Pines. Omne quod movetur necesse est ab aliquo moveri: A refutation of Galen by Alexander of Aphrodisias and the theory of motion. (Latin) [It is necessary for any thing that is to be moved to be moved by someone]. Isis, 52(1):21–54, March 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228340.

Tucker:1961:PTC

[149] Leonard Tucker. President Thomas clap of Yale college: Another “founding father” of American science. Isis, 52(1):55–77, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228341.

Levy:1961:CIS

[150] Jaroslav Levy. Contribution to the identification of some trees and shrubs in the oldest works of European literature. Isis, 52(1):78–86, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228342.

Agassi:1961:UPY

[151] Joseph Agassi and Young Faraday. An unpublished paper of the Young Faraday. Isis, 52(1):87–90, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228343.

Allen:1961:EFR

[152] D. C. Allen. Eloge: Francis Rarick Johnson, 1901–1960. Isis, 52(1):91– 92, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228344.

Woodruff:1961:NC

[153] A. E. Woodruff, Martin Levey, Stillman Drake, O. Neugebauer, L. Sprague de Camp, Norwood Russell Hanson, and Johs. Lohne. Notes & correspondence. Isis, 52(1):93–100, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228345.

Anonymous:1961:NPa

[154] Anonymous. News of the profession. Isis, 52(1):100–105, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228346.

Cohen:1961:BRBa

[155] I. Bernard Cohen. Book review: Histoire de la science by Maurice Daumas. Isis, 52(1):106–107, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228347.

Schramm:1961:BRB

[156] Matthias Schramm. Book review: Welt und Werkstatt des Forschers by Erich Schwinge. Isis, 52(1):107–108, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228348.

Bodemer:1961:BRB

[157] Charles W. Bodemer. Book review: A History of Embryology by Joseph Needham. Isis, 52(1):109–110, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228349.

Kemble:1961:BRB

[158] John Haskell Kemble. Book review: The World’s Lighthouses before 1820 by D. Alan Stevenson. Isis, 52(1):110–111, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228350.

Lytle:1961:BRB

[159] Scott Lytle. Book review: Bernard de Fontenelle: The Idea of Science in the French Enlightenment by Leonard M. Marsak. Isis, 52(1):111– 112, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228351.

Truesdell:1961:BRBa

[160] C. Truesdell. Book review: Die Berliner und die Petersburger Akademie der Wissenschaften im Briefwechsel Leonhard Eulers. Teil I. Der Briefwechsel L. Eulers mit G. F. Muller, 1735–1767 by A. P. Juskevic; E. Winter; P. Hoffmann. Isis, 52(1):113–114, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228352.

Cohen:1961:BRBb

[161] I. Bernard Cohen. Book review: The Correspondence of Isaac Newton. Volume I. 1661–1675 by H. W. Turnbull. Isis, 52(1):114–115, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228353.

Hall:1961:BRB

[162] A. Rupert Hall. Book review: The Correspondence of Isaac Newton. Volume II. 1676–1687 by H. W. Turnbull. Isis, 52(1):115–117, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228354.

Rosenfeld:1961:MNC

[163] L. Rosenfeld. Marginalia to Newton’s correspondence. Isis, 52(1):117– 120, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228355.

VanNiel:1961:BRB

[164] C. B. Van Niel. Book review: Measuring the Invisible World. The Life and Works of Antoni van Leeuwenhoek by A. Schierbeek. Isis, 52(1):120– 122, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228356.

Howells:1961:BRB

[165] W. W. Howells. Book review: The Leopard’s Spots. Scientific Attitudes toward Race in America, 1815–59 by William Stanton. Isis, 52(1):122– 123, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228357.

Anonymous:1961:PRS

[166] Anonymous. Publications received September 1960–February 1961. Isis, 52(1):124–128, March 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228358.

Green:1961:AD

[167] John C. Green. Administrative documents. Isis, 52(1):129–131, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228359.

Anonymous:1961:BMa

[168] Anonymous. Back matter. Isis, 52(1):??, March 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228360.

Woolf:1961:CHQ

[169] Harry Woolf. The conference on the history of quantification in the sciences. Isis, 52(2):133–134, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228675.

Wilks:1961:SAQ

[170] S. S. Wilks. Some aspects of quantification in science. Isis, 52(2):135– 142, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228676.

Crombie:1961:QMP

[171] A. C. Crombie. Quantification in Medieval physics. Isis, 52(2):143– 160, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228677.

Kuhn:1961:FMM

[172] Thomas S. Kuhn. The function of measurement in modern physical science. Isis, 52(2):161–193, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228678.

Guerlac:1961:QC

[173] Henry Guerlac. Quantification in chemistry. Isis, 52(2):194–214, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228679.

Shryock:1961:HQM

[174] Richard H. Shryock. The history of quantification in medical science. Isis, 52(2):215–237, June 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228680.

Boring:1961:BGM

[175] Edwin G. Boring. The beginning and growth of measurement in psychology. Isis, 52(2):238–257, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228681.

Spengler:1961:PQE

[176] Joseph J. Spengler. On the progress of quantification in economics. Isis, 52(2):258–276, June 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228682.

Lazarsfeld:1961:NHQ

[177] Paul F. Lazarsfeld. Notes on the history of quantification in sociology– trends, sources and problems. Isis, 52(2):277–333, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228683.

Gerard:1961:QB

[178] R. W. Gerard. Quantification in biology. Isis, 52(2):334–352, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228684.

Anonymous:1961:BMb

[179] Anonymous. Back matter. Isis, 52(2):353–354, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228685.

Anonymous:1961:FMb

[180] Anonymous. Front matter. Isis, 52(2):??, June 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228674.

Anonymous:1961:FMc

[181] Anonymous. Front matter. Isis, 52(3):353–444, September 1961. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228076.

Hendrickson:1961:NCS

[182] Walter B. Hendrickson. Nineteenth-Century state geological surveys: Early government support of science. Isis, 52(3):357–371, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228077.

Steuer:1961:CPC

[183] Robert O. Steuer. Controversial problems concerning the interpretation of the physiological treatises of papyrus ebers. Isis, 52(3):372–380, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228078.

Dales:1961:RGS

[184] Richard C. Dales. Robert Grosseteste’s scientific works. Isis, 52(3): 381–402, September 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228079.

Africa:1961:CRA

[185] Thomas W. Africa. Copernicus’ relation to Aristarchus and Pythagoras. Isis, 52(3):403–409, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228080.

Herivel:1961:IEN

[186] J. W. Herivel. Interpretation of an early Newton manuscript. Isis, 52 (3):410–416, September 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228081.

Hermelink:1961:NC

[187] H. Hermelink and E. S. Kennedy. Notes & correspondence. Isis, 52(3): 417–419, September 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228082.

Anonymous:1961:NPb

[188] Anonymous. News of the profession. Isis, 52(3):419–420, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228083.

Wilson:1961:BRB

[189] Leonard G. Wilson. Book review: The History of Biology: An Introduction by F. S. Bodenheimer. Isis, 52(3):421–423, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228084.

Zirkle:1961:BRB

[190] Conway Zirkle. Book review: Der Irrweg des Darwinismus by Robert Nachtwey. Isis, 52(3):423–425, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228085.

Spiro:1961:BRB

[191] Melford E. Spiro. Book review: Biennial Review of Anthropology, 1959 by Bernard J. Siegel. Isis, 52(3):425–426, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228086.

Haviland:1961:BRB

[192] James W. Haviland. Book review: Centaur. Essays on the History of Medical Ideas by Felix Marti-Ibanez. Isis, 52(3):426–427, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228087.

Veith:1961:BRB

[193] Ilza Veith. Book review: Parasites and Parasitic Infections in Early Medicine and Science by R. Hoeppli. Isis, 52(3):427–428, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228088.

Grmek:1961:BRB

[194] M. D. Grmek. Book review: Ancient Egyptian and Cnidian Medicine by Robert O. Steuer; J. B. De C. N. Saunders. Isis, 52(3):428, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228089.

Kaminsky:1961:BRB

[195] Howard Kaminsky. Book review: Artes Liberales von der Antiken Bildung zur Wissenschaft des Mittelalters by Josef Koch. Isis, 52(3):429– 430, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228090.

Drake:1961:BRB

[196] Stillman Drake. Book review: Quesiti et Inventioni Diverse by Niccolo Tartaglia. Isis, 52(3):430–431, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228091.

Marsak:1961:BRB

[197] Leonard M. Marsak. Book review: La Metrie’s l’Homme machine by Aram Vartanian. Isis, 52(3):431–432, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228092.

Rath:1961:BRB

[198] Gernot Rath. Book review: Daniel Bernoulli als Physiologe und Statistiker (1700–1782) by Friedrich Huber. Isis, 52(3):432–433, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228093.

Mayr:1961:BRB

[199] Ernst Mayr. Book review: Darwin’s Biological Work: Some Aspects Reconsidered by P. R. Bell. Isis, 52(3):433–435, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228094.

Miller:1961:BRB

[200] Genevieve Miller. Book review: Disease, Life, and Man, Selected Essays by Rudolf Virchow; Lelland J. Rather. Isis, 52(3):436, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228095.

Eaton:1961:BRB

[201] Leonard K. Eaton. Book review: American Building Art–The Nineteenth Century by Carl W. Condit. Isis, 52(3):436–437, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228096.

Klemm:1961:BRB

[202] F. Klemm. Book review: More Studies in Early Petroleum History 1860– 1880 by R. J. Forbes. Isis, 52(3):437–438, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228097.

Hallowell:1961:BRB

[203] A. Irving Hallowell. Book review: The Phenomenon of Man by Pierre Teilhard de Chardin. Isis, 52(3):439–441, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228098.

Woolf:1961:ESC

[204] Harry Woolf, Phyllis Brooks Bosson, and Carol B. Hewitt. Eighty-sixth critical bibliography of the history of science and its cultural influences (to 1 January 1961). Isis, 52(3):445–526, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228099.

Anonymous:1961:BMc

[205] Anonymous. Back matter. Isis, 52(3):527–530, September 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228100.

Anonymous:1961:FMd

[206] Anonymous. Front matter. Isis, 52(4):531–616, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228645.

Middleton:1961:AKB

[207] W. E. Knowles Middleton. Archimedes, Kircher, Buffon, and the burning-mirrors. Isis, 52(4):533–543, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228646.

Guerlac:1961:SDD

[208] Henry Guerlac. Some Daltonian doubts. Isis, 52(4):544–554, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228647. See [494, 826].

Koyre:1961:CMT

[209] Alexandre Koyre and I. Bernard Cohen. The case of the missing tanquam: Leibniz, Newton & Clarke. Isis, 52(4):555–566, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228648.

Kuhn:1961:SCC

[210] Thomas S. Kuhn. Sadi Carnot and the Cagnard engine. Isis, 52(4): 567–574, December 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228649.

See comments [431].

Lemay:1961:FK

[211] J. A. L. Lemay. Franklin and Kinnersley. Isis, 52(4):575–581, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228650.

Gruber:1961:NC

[212] Howard E. Gruber, Øystein Øre, A. Rupert Hall, Marie Boas Hall, and Waclaw Slabczynski. Notes & correspondence. Isis, 52(4):582–586, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228651.

Anonymous:1961:NPc

[213] Anonymous. News of the profession. Isis, 52(4):587–589, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228652.

Multhauf:1961:BRB

[214] Robert P. Multhauf. Book review: The Search for Order: The Development of the Major Ideas in the Physical Sciences from the Earliest Times to the Present by Cecil J. Schneer. Isis, 52(4):589–590, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228653.

Rescher:1961:BRB

[215] Nicholas Rescher. Book review: The Structure of Scientific Thought: An Introduction to Philosophy of Science by Edward H. Madden. Isis, 52 (4):590–591, December 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228654.

Joravsky:1961:BRB

[216] David Joravsky. Book review: Istoriia Akademii Nauk SSSR. Volume I: 1724–1803 by K. V. Ostrovitianov. Isis, 52(4):591–592, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228655.

Scott:1961:BRB

[217] J. F. Scott. Book review: Anthemius of Tralles: A Study in Later Greek Geometry by G. L. Huxley. Isis, 52(4):592–593, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228656.

Greenaway:1961:BRB

[218] Frank Greenaway. Book review: The Story of Chemistry by Georg Lockemann. Isis, 52(4):593–594, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228657.

Dunbar:1961:BRB

[219] M. J. Dunbar. Book review: The Death of Adam: Evolution and Its Impact on Western Thought by John C. Greene. Isis, 52(4):594–596, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228658.

Merton:1961:BRB

[220] Robert K. Merton. Book review: The Idea of a Social Science and Its Relation to Philosophy by Peter Winch. Isis, 52(4):596–599, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228659.

Fleming:1961:BRB

[221] Donald Fleming. Book review: Medicine and Society in America: 1660– 1860 by Richard Harrison Shryock. Isis, 52(4):599–600, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228660.

Saunders:1961:BRB

[222] J. B. deC. M. Saunders. Book review: On the History of Medicine by Henry E. Sigerist; Felix Marti-Ibanez; On the Sociology of Medicine by Henry E. Sigerist; Milton E. Roemer. Isis, 52(4):600–601, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228661.

Hubotter:1961:BRB

[223] Franz Hubotter. Book review: La Medecine Chinoise au Cours des Siècles by Pierre Huard; Ming Wong. Isis, 52(4):601–602, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228662.

Murdoch:1961:BRB

[224] John E. Murdoch. Book review: Ramus, Method, and the Decay of Dialogue: From the Art of Discourse to the Art of Reason by Walter J. Ong. Isis, 52(4):602–606, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228663.

DeSollaPrice:1961:BRB

[225] Derek J. De Solla Price. Book review: The Globe of Martin Bylica of Olkusz and Celestial Maps in the East and in the West by Zofia Ameisenowa; Andrzej Potocki. Isis, 52(4):606, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228664.

Truesdell:1961:BRBb

[226] C. Truesdell. Book review: Novalis and Mathematics: A Study of Friedrich von Hardenberg’s Fragments on Mathematics and Its Relation to Magic, Music, Religion, Philosophy, Language and Literature by Martin Dyck. Isis, 52(4):606–607, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228665.

Stahlman:1961:BRB

[227] William D. Stahlman. Book review: The Transits of Venus: A Study of Eighteenth-Century Science by Harry Woolf. Isis, 52(4):607–608, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228666.

Siegfried:1961:BRB

[228] Robert Siegfried. Book review: Robert Boyle: Father of Chemistry by Roger Pilkington. Isis, 52(4):609, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228667.

Clark:1961:BRB

[229] Thomas D. Clark. Book review: Thomas Gage’s Travels in the New World by J. Eric S. Thompson. Isis, 52(4):609–610, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228668.

Bishop:1961:BRB

[230] Philip W. Bishop. Book review: History of the Milling Machine: A Study in Technological Development by Robert S. Woodbury. Isis, 52 (4):610–611, December 1961. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228669.

Lorch:1961:BRB

[231] Jaacov Lorch. Book review: A Biologist in Israel: A Book of Reminiscences by F. S. Bodenheimer. Isis, 52(4):611–612, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228670.

Anonymous:1961:VI

[232] Anonymous. Volume information. Isis, 52(4):617–620, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228644.

Anonymous:1961:AD

[233] Anonymous. Administrative documents. Isis, 52(4):621–623, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228671.

Anonymous:1961:BMd

[234] Anonymous. Back matter. Isis, 52(4):??, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228673.

Anonymous:1961:IIE

[235] Anonymous. [illustrations]: Interpretation of an early Newton manuscript. Isis, 52(4):??, December 1961. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228672.

Anonymous:1962:FMa

[236] Anonymous. Front matter. Isis, 53(1):1–2, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227822.

Woolf:1962:MHS

[237] Harry Woolf. Manuscripts and the history of science. Isis, 53(1):3– 4, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227823.

Anonymous:1962:BMa

[238] Anonymous. Back matter. Isis, 53(1):5–157, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227846.

Shryock:1962:VHM

[239] Richard H. Shryock. The viewpoint of an historian and a manuscript librarian. Isis, 53(1):7–13, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227824.

Bell:1962:ESP

[240] Whitfield J. Bell, Jr. Editing a Scientist’s papers. Isis, 53(1):14–20, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227825.

Lurie:1962:SOR

[241] Edward Lurie. Some observations on research in Nineteenth-Century American science. Isis, 53(1):21–30, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227826.

Hewlett:1962:PSC

[242] R. G. Hewlett. A pilot study in contemporary scientific history. Isis, 53(1):31–38, March 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227827.

Guerlac:1962:Da

[243] Dr. Guerlac, Dr. Benison, Robert P. Multhauf, Dr. Lurie, Edith M. Fox, Dr. Gruman, Leland Anderson, R. Bruce Lindsay, Karl Honaman, Mr. Shafer, and Dr. Posner. Discussion. Isis, 53(1):39–51, March 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227828.

Grover:1962:RAP

[244] Wayne C. Grover. The role of the archivist in the preservation of scientific records. Isis, 53(1):55–62, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227829.

Dupree:1962:WMH

[245] A. Hunter Dupree. What manuscripts the historian wants saved. Isis, 53(1):63–66, March 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227830.

Alpert:1962:SRV

[246] Harry Alpert. Science records: Viewpoints of the sociology of science. Isis, 53(1):67–71, March 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227831.

Birr:1962:WSW

[247] Kendall Birr. “What shall we save?” An Historian’s view. Isis, 53(1): 72–79, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227832.

Odishaw:1962:WSW

[248] Hugh Odishaw. What shall we save in the geophysical sciences? Isis, 53(1):80–86, March 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227833.

Multhauf:1962:D

[249] Robert P. Multhauf, Ralph H. Gabriel, Nathan Reingold, Dr. Dupree, and Luther Evans. Discussion. Isis, 53(1):87–98, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227834.

Heumann:1962:CSM

[250] Karl F. Heumann. The collection of scientific manuscripts. Isis, 53(1): 99, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227835.

Evans:1962:SPA

[251] Luther H. Evans. Some proposals for action. Isis, 53(1):101–105, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227836.

Reingold:1962:SAD

[252] Nathan Reingold. Subject analysis and description of manuscript collections. Isis, 53(1):106–112, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227837.

Benison:1962:OHM

[253] Saul Benison. Oral history and manuscript collecting. Isis, 53(1):113– 117, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227838.

Fleming:1962:EVP

[254] Donald Fleming. The ends in view of the preservation of the private papers of American scientists. Isis, 53(1):118–121, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227839.

Heumann:1962:D

[255] Karl F. Heumann, Oliver W. Holmes, Dr. Reingold, Dr. Guerlac, Dr. Shryock, Dr. Leake, Dr. Born, and Dr. Hewlett. Discussion. Isis, 53(1):

123–135, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227840.

Glass:1962:SPS

[256] H. Bentley Glass. The scientist and the preservation of science manuscripts. Isis, 53(1):136–142, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227841.

Leake:1962:RSA

[257] Chauncey D. Leake. Responsibility for science archives. Isis, 53(1):143– 146, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227842.

Guerlac:1962:Db

[258] Dr. Guerlac and Gerald J. Gruman. Discussion. Isis, 53(1):146–148, March 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227843.

Anonymous:1962:PCS

[259] Anonymous. Participants in the conference on science manuscripts. Isis, 53(1):149–150, March 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227844.

Quirk:1962:PSM

[260] R. N. Quirk. The problem of scientific manuscripts in Britain. Isis, 53(1):151–154, March 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227845.

Anonymous:1962:FMb

[261] Anonymous. Front matter. Isis, 53(2):159, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228027.

Burstyn:1962:GAP

[262] Harold L. Burstyn. Galileo’s attempt to prove that the Earth moves. Isis, 53(2):161–185, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228028.

Gruber:1962:ERD

[263] Howard E. Gruber and Valmai Gruber. The eye of reason: Darwin’s development during the Beagle voyage. Isis, 53(2):186–200, June 1962.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228029.

Drake:1962:GGX

[264] Stillman Drake and Galileo Galilei. Galileo gleanings XII: An unpublished letter of Galileo to Peiresc. Isis, 53(2):201–211, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228030.

Herivel:1962:NRB

[265] J. W. Herivel and Isaac Newton. Newton on rotating bodies. Isis, 53(2): 212–218, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228031.

Guerlac:1962:NC

[266] Henry Guerlac and Johs. Lohne. Notes & correspondence. Isis, 53(2): 219–226, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228032.

Anonymous:1962:NPa

[267] Anonymous. News of the profession. Isis, 53(2):226–229, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228033.

Hoskin:1962:BRB

[268] M. A. Hoskin. Book review: Critical Problems in the History of Science: Proceedings of the Institute for the History of Science at the University of Wisconsin, September 1–11, 1957 by Marshall Clagett. Isis, 53(2):230– 231, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228034.

Hall:1962:BRBa

[269] A. Rupert Hall. Book review: The Western Intellectual Tradition: From Leonardo to Hegel by J. Bronowski; Bruce Mazlish. Isis, 53(2):231–232, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228035.

Toulmin:1962:BRB

[270] Stephen Toulmin. Book review: Philosophy of Science by Arthur Danto; Sidney Morgenbesser. Isis, 53(2):232–234, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228036.

Salmon:1962:BRB

[271] Wesley C. Salmon. Book review: An Introduction to the Logic of the Sciences by R. Harre; The General Science of Nature by Vincent Edward Smith; The Nature of Science by David Greenwood. Isis, 53(2):234– 235, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228037.

Hall:1962:BRBb

[272] Marie Boas Hall. Book review: A History of Greek Fire and Gunpowder by J. R. Partington. Isis, 53(2):236, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228038.

Kennedy:1962:BRB

[273] E. S. Kennedy. Book review: The Observatory in Islam and Its Place in the General History of the Observatory by Aydin Sayili. Isis, 53(2): 237–239, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228039.

Hiebert:1962:BRBa

[274] Erwin N. Hiebert. Book review: America in the Antarctic to 1840 by Philip I. Mitterling. Isis, 53(2):239–240, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228040.

Zirkle:1962:BRBa

[275] Conway Zirkle. Book review: Classic Papers in Genetics by James A. Peters. Isis, 53(2):240–241, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228041.

Mayr:1962:BRB

[276] Ernst Mayr. Book review: Geschichte der Zoologie und der Zoologischen Anstalten in Iena 1779–1919 by Georg Uschmann. Isis, 53(2):241, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228042.

Poynter:1962:BRB

[277] F. N. L. Poynter. Book review: The Royal College of Surgeons of England: A History by Zachary Cope. Isis, 53(2):241–242, June 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228043.

Mehl:1962:BRB

[278] Robert F. Mehl. Book review: A History of Metals by Leslie Aitchison. Isis, 53(2):242–244, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228044.

Maryon:1962:BRB

[279] Herbert Maryon. Book review: A History of Metallography by Cyril Stanley Smith. Isis, 53(2):244–246, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228045.

Tierney:1962:BRB

[280] James J. Tierney. Book review: The Geographical Fragments of Hipparchus by D. R. Dicks. Isis, 53(2):246–247, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228046.

Edelstein:1962:BRB

[281] Ludwig Edelstein. Book review: Medizin und Arzt bei den Griechischen Kirchenvatern bis Chrysostomos by Hermann Josef Frings. Isis, 53(2): 247–248, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228047.

Peck:1962:BRB

[282] A. L. Peck. Book review: Aristotle by John Herman Randall. Isis, 53(2): 248–249, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228048.

Diamadopoulos:1962:BRB

[283] Peter Diamadopoulos. Book review: Heraclitus by Philip Wheelwright. Isis, 53(2):249–250, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228049.

Clagett:1962:BRB

[284] Marshall Clagett. Book review: Le Systeme du Monde: Histoire des Doctrines Cosmologiques de Platon a Copernic by Pierre Duhem. Isis, 53(2):251–252, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228050.

Carmody:1962:BRB

[285] Francis J. Carmody. Book review: Al-Biruni on Transits by Mohammad Saffouri; Adnan Ifram. Isis, 53(2):252–253, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228051.

Rosen:1962:BRBa

[286] Edward Rosen. Book review: La Science au seizième siècle: Colloque international de Royaumont, 1–4 Juillet 1957. Isis, 53(2):253–254, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228052.

Debus:1962:BRB

[287] Allen G. Debus. Book review: Spiritual and Demonic Magic from Ficino to Campanella by D. F. Walker. Isis, 53(2):254–255, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228053.

Williams:1962:BRBa

[288] L. Pearce Williams. Book review: Henry Cavendish: His Life and Scientific Work by A. J. Berry. Isis, 53(2):255–256, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228054.

Hindle:1962:BRBa

[289] Brooke Hindle. Book review: A Reasonable Rebel: George Christoph Lichtenberg by Carl Brinitzer. Isis, 53(2):256–257, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228055.

Wheatley:1962:BRB

[290] J. M. O. Wheatley. Book review: Francis Bacon: The First Statesman of Science by J. G. Crowther. Isis, 53(2):257–258, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228056.

LeFanu:1962:BRB

[291] W. R. LeFanu. Book review: Doctor Jenner of Berkeley by Dorothy Fisk. Isis, 53(2):259, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228057.

Hellman:1962:BRB

[292] C. Doris Hellman. Book review: The Controversy on the Comets of 1618 by Galileo Galilei; Horatio Grassi; Mario Guiducci; Johann Kepler; Stillman Drake; C. D. O’Malley. Isis, 53(2):259–261, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228058.

Eisele:1962:BRB

[293] Carolyn Eisele. Book review: Classical Mathematics: A Concise History of the Classical Era in Mathematics by Joseph Ehrenfried Hofmann. Isis, 53(2):261–262, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228059.

Marsak:1962:BRB

[294] Leonard M. Marsak. Book review: French Free Thought from Gassendi to Voltaire by J. S. Spink. Isis, 53(2):263, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228060.

Ravetz:1962:BRB

[295] J. R. Ravetz. Book review: The Rational Mechanics of Flexible or Elastic Bodies by C. Truesdell. Isis, 53(2):263–266, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228061.

Emerson:1962:BRB

[296] Donald E. Emerson. Book review: Aufklaerung und Revolution. Historia Mundi, Ein Handbuch der Weltgeschichte by Fritz Valjavec. Isis, 53(2): 266–267, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228062.

Boyer:1962:BRB

[297] Carl B. Boyer. Book review: Mathematiques et Metaphysique chez Descartes by Jules Vuillemin. Isis, 53(2):267–268, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228063.

Leicester:1962:BRB

[298] Henry M. Leicester. Book review: Istoriya Klassicheskoi Teorii Khimicheskogo Stroeniya. [The History of the Classical Theory of Chemical Structure] by G. V. Bykov. Isis, 53(2):268–269, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228064.

Ferguson:1962:BRB

[299] Eugene S. Ferguson. Book review: The American Civil Engineer: Origins and Conflict by Daniel Hovey Calhoun. Isis, 53(2):269–270, June 1962.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228065.

Condit:1962:BRB

[300] Carl W. Condit. Book review: Louis Sullivan as He Lived: The Shaping of American Architecture by Willard Connely. Isis, 53(2):270–272, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228066.

Carlson:1962:BRB

[301] Robert E. Carlson. Book review: From Engines to Autos: Five Pioneers in Engine Development and Their Contributions to the Automotive Industry by Eugen Diesel; Gustav Goldbeck; Friedrich Schildberger. Isis, 53(2):272–273, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228067.

Gillispie:1962:BRB

[302] Charles C. Gillispie. Book review: Naturphilosophie im 19. Jahrhundert by Gerhard Hennemann. Isis, 53(2):273–275, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228068.

Drake:1962:BRBa

[303] Stillman Drake. Book review: The Concepts and Theories of Modern Physics by J. B. Stallo; P. W. Bridgman. Isis, 53(2):275–276, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228069.

Wilson:1962:BRBa

[304] Leonard G. Wilson. Book review: Claude Bernard and His Place in the History of Ideas by Reino Virtanen. Isis, 53(2):276–277, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228070.

Williams:1962:BRBb

[305] L. Pearce Williams. Book review: Elihu Thomson: Beloved Scientist, 1853–1937 by David O. Woodbury. Isis, 53(2):278–279, June 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228071.

Schmidl:1962:BRB

[306] Fritz Schmidl. Book review: Letters of Sigmund Freud by Sigmund Freud; Ernest L. Freud; James Stern; Tania Stern. Isis, 53(2):279–280, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228072.

Oppenheimer:1962:BRB

[307] Jane Oppenheimer. Book review: In and Out of the Ivory Tower: The Autobiography of Richard B. Goldschmidt by Richard B. Goldschmidt. Isis, 53(2):280–281, June 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228073.

Anonymous:1962:BMb

[308] Anonymous. Back matter. Isis, 53(2):282–284, June 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228075.

Anonymous:1962:CFM

[309] Anonymous. Correction: Front matter. Isis, 53(2):284, June 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228074.

Anonymous:1962:FMc

[310] Anonymous. Front matter. Isis, 53(3):285, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227783.

Hoff:1962:BGR

[311] Hebbel E. Hoff and L. A. Geddes. The beginnings of graphic recording. Isis, 53(3):287–324, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227784.

Coleman:1962:LRS

[312] William Coleman. Lyell and the “reality” of species: 1830–1833. Isis, 53 (3):325–338, September 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227785.

Westfall:1962:DNT

[313] Richard S. Westfall. The development of Newton’s theory of color. Isis, 53(3):339–358, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227786.

Hanson:1962:LZN

[314] Norwood Russell Hanson. Leverrier: The zenith and nadir of Newtonian mechanics. Isis, 53(3):359–378, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227787.

Cohen:1962:FEV

[315] I. Bernard Cohen. The first English version of Newton’s hypotheses non fingo. Isis, 53(3):379–388, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227788.

Williams:1962:NC

[316] L. Pearce Williams. Notes & correspondence. Isis, 53(3):389–392, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227789.

Anonymous:1962:NPb

[317] Anonymous. News of the profession. Isis, 53(3):392–393, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227790.

Anonymous:1962:CPS

[318] Anonymous. Correction: The problem of scientific manuscripts in Britain. Isis, 53(3):393, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227791.

Schofield:1962:BRB

[319] Robert E. Schofield. Book review: The Major Achievements of Science by A. E. E. McKenzie. Isis, 53(3):394–395, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227792.

Butterfield:1962:BRB

[320] Herbert Butterfield. Book review: Science Since Babylon by Derek J. De Solla Price. Isis, 53(3):395–396, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227793.

Zirkle:1962:BRBb

[321] Conway Zirkle. Book review: Darwin’s Vision and Christian Perspectives by Walter J. Ong. Isis, 53(3):396–397, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227794.

Williams:1962:BRBc

[322] L. Pearce Williams. Book review: Oersted and the Discovery of Electromagnetism by Bern Dibner. Isis, 53(3):397–398, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227795.

Carozzi:1962:BRB

[323] Albert V. Carozzi. Book review: Geologie und Palaontologie in Texten und ihrer Geschichte by Helmut Holder. Isis, 53(3):398–400, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227796.

Boyd:1962:BRB

[324] Mary Boyd. Book review: The Discovery of the Pacific Islands by Andrew Sharp. Isis, 53(3):400, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227797.

Kilgour:1962:BRB

[325] Frederick G. Kilgour. Book review: Ironworks on the Saugus: The Lynn and Braintree Ventures of the Company of Undertakers of the Ironworks in New England by E. N. Hartley. Isis, 53(3):401, September 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227798.

Klemm:1962:BRB

[326] F. Klemm. Book review: The Story of Engineering by James Kip Finch. Isis, 53(3):401–402, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227799.

OMalley:1962:BRB

[327] C. D. O’Malley. Book review: On Anatomical Procedures by Galen; W. L. H. Duckworth; M. C. Lyons; B. Towers. Isis, 53(3):402–403, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227800.

Kirk:1962:BRB

[328] G. S. Kirk. Book review: Anaximander and the Origins of Greek Cosmology by Charles H. Kahn. Isis, 53(3):403–405, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227801.

Bockstaele:1962:BRB

[329] Paul P. Bockstaele. Book review: De Zestiende-eeuwse Nedelandse Rekenboeken by A. J. E. M. Smeur. Isis, 53(3):405–406, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227802.

Spitz:1962:BRB

[330] Lewis W. Spitz. Book review: Renaissance Concepts of Method by Neal Ward Gilbert. Isis, 53(3):406–408, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227803.

Allen:1962:BRB

[331] Don Cameron Allen. Book review: Les Animaux dans la Poesie Française de la Renaissance by Helene Nais. Isis, 53(3):408, September 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227804.

Drake:1962:BRBb

[332] Stillman Drake. Book review: Galileo in China by Pasquale M. D’Elia; Rufus Suter; Matthew Sciascia. Isis, 53(3):409–410, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227805.

Herivel:1962:BRB

[333] J. W. Herivel. Book review: The Royal Society: Its Origins and Founders by Harold Hartley. Isis, 53(3):410–411, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227806. See correction [358].

Truesdell:1962:BRB

[334] C. Truesdell. Book review: Die Berliner und die Petersburger Akademie der Wissenschaften im Briefwechsel Leonhard Eulers. Teil 2. Der Briefwechsel L. Eulers mit Nartov, Razumovskij, Schumacher, Teplov und der Petersburger Akademie, 1730–1763 by A. P. Juskevic; E. Winter; P. Hoffmann; Ju. Ch. Kopelevic. Isis, 53(3):411–413, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227807.

Popkin:1962:BRB

[335] Richard H. Popkin. Book review: Dissertations en forme de paradoxes contre les Aristoteliciens (Exercitationes Paradoxicae Adversus Aristoteleos) by Pierre Gassendi; Bernard Rochot. Isis, 53(3):413–415, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227808.

Agnew:1962:BRB

[336] L. R. C. Agnew. Book review: De Motu Locali Animalium 1627 by William Harvey; Gweneth Whitteridge. Isis, 53(3):415–416, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227809.

Cope:1962:BRB

[337] Jackson I. Cope. Book review: Abraham Cowley’s World of Order by Robert B. Hinman. Isis, 53(3):416–418, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227810.

Hall:1962:BRBc

[338] Marie Boas Hall. Book review: A Bibliography of the Honourable Robert Boyle Fellow of the Royal Society by John F. Fulton. Isis, 53(3):419–420, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227811.

Hall:1962:BRBd

[339] Marie Boas Hall. Book review: A Bibliography of Dr. Robert Hooke by Geoffrey Keynes. Isis, 53(3):420–421, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227812.

Barber:1962:BRB

[340] Bernard Barber. Book review: Soviet Marxism and Natural Science, 1917–1932 by David Joravsky. Isis, 53(3):421–422, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227813.

Hendrickson:1962:BRB

[341] Walter B. Hendrickson. Book review: A History of the United States Weather Bureau by Donald R. Whitnah. Isis, 53(3):422–423, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227814.

Hindle:1962:BRBb

[342] Brooke Hindle. Book review: Biographical Sketch and Guide to the Writings of Charles Caldwell, M. D. (1772–1853) with Sections on Phrenology and Hypnotism by Emmet Field Horine. Isis, 53(3):423–424, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227815.

Stanton:1962:BRB

[343] William Stanton. Book review: Louis Agassiz: A Life in Science by Edward Lurie. Isis, 53(3):424–426, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227816.

Siegfried:1962:BRB

[344] Robert Siegfried. Book review: Edward Williams Morley: His Influence on Science in America by Howard R. Williams; Michelson and the Speed of Light by Bernard Jaffe. Isis, 53(3):426–428, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227817.

Gruber:1962:BRB

[345] Jacob W. Gruber. Book review: Darwin’s Place in History by C. D. Darlington. Isis, 53(3):428–429, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227818.

Struve:1962:BRB

[346] Otto Struve. Book review: Source Book in Astronomy, 1900–1960 by Harlow Shapley. Isis, 53(3):430–431, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227819.

Anonymous:1962:BMc

[347] Anonymous. Back matter. Isis, 53(3):432–436, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227821.

Greene:1962:AD

[348] John C. Greene. Administrative documents. Isis, 53(3):433–435, September 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227820.

Anonymous:1962:FMd

[349] Anonymous. Front matter. Isis, 53(4):437–544, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227718.

Woodruff:1962:ADN

[350] A. E. Woodruff. Action at a distance in Nineteenth Century electrodynamics. Isis, 53(4):439–459, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227719.

Sambursky:1962:GAC

[351] S. Sambursky. Galileo’s attempt at a cosmogony. Isis, 53(4):460–464, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227720.

Vucinich:1962:NIL

[352] Alexander Vucinich. Nikolai Ivanovich Lobachevskii: The man behind the first non-Euclidean geometry. Isis, 53(4):465–481, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227721.

Hall:1962:WBO

[353] A. Rupert Hall and Marie Boas Hall. Why blame Oldenburg? Isis, 53 (4):482–491, December 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227722.

Abbud:1962:PTI

[354] Fuad Abbud. The planetary theory of Ibn al-Shatir: Reduction of the geometric models to numerical tables. Isis, 53(4):492–499, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227723.

Deischer:1962:EEV

[355] Claude K. Deischer. Eloge: Eva Vivian Armstrong, 1877–1962: Historian and friend of historians. Isis, 53(4):500–501, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227724.

Ballabh:1962:NC

[356] R. Ballabh, Edward Rosen, Thomas W. Africa, and Lynn Thorndike. Notes & correspondence. Isis, 53(4):502–510, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227725.

Anonymous:1962:NPc

[357] Anonymous. News of the profession. Isis, 53(4):511–512, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227726.

Anonymous:1962:CRS

[358] Anonymous. Correction: The Royal Society: Its origins and founders. Isis, 53(4):512, December 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227727.

See [333].

Cohen:1962:BRBa

[359] I. Bernard Cohen. Book review: Smithsonian Treasury of Science by Webster P. True. Isis, 53(4):513–514, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227728.

Roger:1962:BRB

[360] Jacques Roger. Book review: The History of Scepticism from Erasmus to Descartes by Richard H. Popkin. Isis, 53(4):514–515, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227729.

Rescher:1962:BRB

[361] Nicholas Rescher. Book review: Concise History of Logic by Heinrich Scholz; Kurt F. Leidecker. Isis, 53(4):515–516, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227730.

Hiebert:1962:BRBb

[362] Erwin N. Hiebert. Book review: Great Experiments in Physics by Morris H. Shamos. Isis, 53(4):516–517, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227731.

Rosen:1962:BRBb

[363] Edward Rosen. Book review: La Revolution Astronomique: Copernic, Kepler, Borelli by Alexandre Koyre. Isis, 53(4):517–519, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227732.

Veith:1962:BRB

[364] Ilza Veith. Book review: Die Historischen Grundlagen der Leberforschung. Teil 1. Die Vorstellungen über Anatomie, Physiologie und Pathologie der Leber in der Antike by Nikolaus Mani. Isis, 53(4):519– 520, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227733.

Williams:1962:BRBd

[365] L. Pearce Williams. Book review: Ces inconnus ont fait le siècle by Pierre Rousseau. Isis, 53(4):520–521, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227734.

Strauss:1962:BRB

[366] Felix F. Strauss. Book review: Bergbau und Bauleute: Neue Beitrage zur Geschichte des Bergbaus und der Geologie. Isis, 53(4):521–523, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227735.

Bruins:1962:BRB

[367] E. M. Bruins. Book review: Egyptian Astronomical Texts. Volume 1. The Early Decans by O. Neugebauer; R. Parker. Isis, 53(4):523–525, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227736.

Almgren:1962:BRB

[368] Beverly S. Almgren. Book review: Leonardo da Vinchi, 1452–1519 by V. P. Zubov. Isis, 53(4):525–527, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227737.

Pagel:1962:BRB

[369] Walter Pagel. Book review: Samtliche Werke. Zweite Abteilung: Theologische und religions-philosophische Schriften: Band IV: Auslegung des Psalters Davids, Teil I. Kommentar zu den Psalmen 75 (76) bis 102 (103) by [Paracelsus] Theophrast von Hohenheim; Kurt Goldammer; Samtliche Werke. Zweite Abteilung: Theologische und religions-philosophische Schriften: Band V: Auslegung des Psalters Davids, Teil II. Kommentar zu den Psalmen 103 (104) bis 117 (118); Samtliche Werke. Zweite Abteilung: Theologische und religionsphilosophische Schriften: Band VI: Auslegung des Psalters Davids, Teil III. Kommentar zu den Psalmen 118 (119) bis 137 (138); Samtliche Werke. Erste Abteilung: Medizinische, naturwissenschaftliche und philosophische Schristen by [Paracelsus] Theophrast von Hohenheim; Karl Sudhoff; Liber de Nymphis, Sylphis, Pygmaeis et Salamandris et de caeteris Spiritibus by [Paracelsus] Theophrastus von Hohenheim; Robert Blaser; Himmel und Erde machen den Menschen. Weisheit des Paracelsus: Aus dem Gesamtwerk des Theophrast von Hohenheim ausgewahlt und übertragen by Robert Blaser; Paracelsus: Vom Licht der Natur und des Geistes, eine Auswahl by Kurt Goldammer; Karl-Heinz Weimann. Isis, 53(4):527–530, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227738.

Hahn:1962:BRB

[370] Roger Hahn. Book review: Lavoisier et la Revolution française. Vol. II. Le journal de Fougeroux de Bondaroy by Lucien Scheler; W. A. Smeaton. Isis, 53(4):530–532, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227739.

Cohen:1962:BRBb

[371] I. Bernard Cohen. Book review: “Rational Fluid Mechanics, 1687–1765” by Clifford Ambrose Truesdell. Isis, 53(4):532–533, December 1962. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227740.

Belaval:1962:BRB

[372] Yvon Belaval. Book review: Leibniz et la dynamique: Les textes de 1692 by Pierre Costabel. Isis, 53(4):533–535, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227741.

Bodemer:1962:BRB

[373] Charles W. Bodemer. Book review: Istoriia Embriologii v Rossii XIX– XX VV. bespozvonotschnie. (History of Embryology in Russia, 19th and 20th Centuries. Invertebrate) by L. Ia Bliakher. Isis, 53(4):535–537, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227742.

Hiebert:1962:BRBc

[374] Erwin N. Hiebert. Book review: Kekule te Gent (1858–1867): De Geschiedenis van de Benoeming van August Kekule te Gent en de Oprichting van het eerste Onderrichts-laboratorium voor Scheikunde in Belgie by J. Gillis. Isis, 53(4):537–538, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227743.

Loewenberg:1962:BRB

[375] Bert James Loewenberg. Book review: A Conscience in Conflict: The Life of St. George Jackson Mivart by Jacob W. Gruber. Isis, 53(4): 538–540, December 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227744.

Miller:1962:BRB

[376] Genevieve Miller. Book review: The Library of the Medical Institution of Yale College and Its Catalogue of 1865 by Frederick G. Kilgour. Isis, 53 (4):540–541, December 1962. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227745.

Wilson:1962:BRBb

[377] Leonard G. Wilson. Book review: The Growth of Scientific Physiology: Physiological Method and the Mechanist — Vitalist Controversy, Illustrated by the Problems of Respiration and Animal Heat by G. J. Goodfield. Isis, 53(4):541–542, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227746.

Woolf:1962:ESC

[378] Harry Woolf and Aleida Thompson. Eighty-seventh critical bibliography of the history of science and its cultural influences (to 1 January 1962). Isis, 53(4):545–648, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227747.

Anonymous:1962:VI

[379] Anonymous. Volume information. Isis, 53(4):649–654, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227717.

Anonymous:1962:BMd

[380] Anonymous. Back matter. Isis, 53(4):??, December 1962. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227748.

Anonymous:1963:FMa

[381] Anonymous. Front matter. Isis, 54(1):1–3, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228723.

Woolf:1963:B

[382] Harry Woolf. ISIS: 1913–1963. Isis, 54(1):5, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228724.

Sarton:1963:B

[383] George Sarton. Le but d’Isis. Isis, 54(1):7–10, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228725.

Middleton:1963:PTH

[384] W. E. Knowles Middleton. The place of Torricelli in the history of the barometer. Isis, 54(1):11–28, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228726.

Wong:1963:COW

[385] George H. C. Wong. China’s opposition to Western science during Late Ming and Early Ch’ing. Isis, 54(1):29–49, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228727. See comments [700].

Holmes:1963:EAO

[386] Frederic L. Holmes. Elementary analysis and the origins of physiological chemistry. Isis, 54(1):50–81, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228728.

Westfall:1963:NRH

[387] Richard S. Westfall. Newton’s reply to Hooke and the theory of colors. Isis, 54(1):82–96, March 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228729.

Bosworth:1963:PAE

[388] C. E. Bosworth. A pioneer Arabic encyclopedia of the sciences: Al Khwarizmi’s keys of the sciences. Isis, 54(1):97–111, March 1963. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228730.

Lattimore:1963:EVS

[389] Owen Lattimore. Eloge: Vilhjalmur Stefansson (1879–1962). Isis, 54(1): 112–113, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228731.

Millas-Vallicrosa:1963:ADC

[390] J.-M. Millas-Vallicrosa. La autenticidad del comentario a las tablas astronomicas de al-Jwarizmi por Ahmad ibn al-Mutanna’. (spanish) [The authenticity of the commentary on the astronomical tables of al-Khwarizmi by Ahmad ibn al-Muthanna’]. Isis, 54(1):114–119, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228732.

MacKinney:1963:MHA

[391] Loren MacKinney. Moon-happy apes, monkeys and baboons. Isis, 54 (1):120–122, March 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228733.

deCamp:1963:XOG

[392] L. Sprague de Camp. Xerxes’ okapi and Greek geography. Isis, 54(1): 123–125, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228734.

Semmel:1963:PMS

[393] Bernard Semmel. Parliament and the metric system. Isis, 54(1):125–133, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228735. See comments

[836, 837, 986].

Goldberg:1963:NMH

[394] J. Philip Goldberg. The numerical method: How it struck a contemporary. Isis, 54(1):133–135, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228736.

Anonymous:1963:HSP

[395] Anonymous. The Henry Schuman Prize in the History of Science. Isis, 54(1):135, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228737.

Anonymous:1963:NPa

[396] Anonymous. News of the profession. Isis, 54(1):136–140, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228738.

Klemm:1963:BRB

[397] Friedrich Klemm. Book review: Sowjetische Beitrage zur Geschichte der Naturwissenschaft by Gerhard Harig. Isis, 54(1):141–143, March 1963.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228739.

Boyer:1963:BRBa

[398] Carl B. Boyer. Book review: Nuevos Estudios Sobre Historia de la Ciencia Espanola by Jose M. Millas-Vallicrosa. Isis, 54(1):143–144, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228740.

Hall:1963:BRBa

[399] Marie Boas Hall. Book review: The Life and Work of William Higgins, Chemist (1763–1825) by T. S. Wheeler; J. R. Partington; William Higgins. Isis, 54(1):144–145, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228741.

Hiebert:1963:BRBa

[400] Erwin N. Hiebert. Book review: A History of Polar Exploration by L. P. Kirwan. Isis, 54(1):145–147, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228742.

Porter:1963:BRBa

[401] J. R. Porter. Book review: Milestones in Microbiology by Thomas Dale Brock. Isis, 54(1):147, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228743.

Porter:1963:BRBb

[402] J. R. Porter. Book review: Pioneer Microbiologists of America by Paul Franklin Clark. Isis, 54(1):147–148, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228744.

Greene:1963:BRB

[403] John C. Greene. Book review: Evolution and Christians by Philip G. Fothergill. Isis, 54(1):148–150, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228745.

Stannard:1963:BRB

[404] Jerry Stannard. Book review: Recherches sur le traite pseudohippocratique du Regime by Robert Joly. Isis, 54(1):150–151, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228746.

Vlastos:1963:BRB

[405] Gregory Vlastos. Book review: Aristotle’s System of the Physical World by Friedrich Solmsen. Isis, 54(1):151–152, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228747.

DeSollaPrice:1963:BRB

[406] Derek J. De Solla Price. Book review: The Planetary Equatorium of Jamshid Ghiyath al-Din al-Kashi (d. 1429) by Jamshid Guiyath al-Din al-Kashi; E. S. Kennedy. Isis, 54(1):153–154, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228748.

Hummel:1963:BRB

[407] Arthur W. Hummel. Book review: Heavenly Clockwork: The Great Astronomical Clocks of Medieval China by Joseph Needham; Wang Ling; Derek J. de Solla Price. Isis, 54(1):154–155, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228749.

Brown:1963:BRB

[408] Harcourt Brown. Book review: Oeuvres Completes by Bernard Palissy. Isis, 54(1):155–156, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228750.

Marsak:1963:BRB

[409] Leonard M. Marsak. Book review: Fontenelle: Sa vie et son oeuvre, 1657–1757 by Suzanne Delorme. Isis, 54(1):156–157, March 1963. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228751.

Hindle:1963:BRB

[410] Brooke Hindle. Book review: Mark Catesby: The Colonial Audubon by George F. Frick; Raymond P. Stearns. Isis, 54(1):157–158, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228752.

Leicester:1963:BRBa

[411] Henry M. Leicester. Book review: Lavoisier — The Crucial Year: The Background and Origin of His First Experiments on Combustion in 1772 by Henry Guerlac. Isis, 54(1):158–159, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228753.

Westfall:1963:BRBa

[412] Richard S. Westfall. Book review: Unpublished Scientific Papers of Isaac Newton. A Selection from the Portsmouth Collection in the University Library, Cambridge by A. Rupert Hali; Marie Boas Hall. Isis, 54(1):159– 160, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228754.

Doorman:1963:BRB

[413] G. Doorman. Book review: The Two Netherlanders: Humphrey Bradley and Cornelius Drebbel by L. E. Harris. Isis, 54(1):160–163, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228755.

Allen:1963:BRBa

[414] D. C. Allen. Book review: Sir Thomas Browne: A Biographical and Critical Study by Frank Livingstone Huntley. Isis, 54(1):163–164, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228756.

Coleman:1963:BRBa

[415] William Coleman. Book review: Titian Ramsay Peale, 1799–1885, and His Journals of the Wilkes Expedition by Jessie Poesch. Isis, 54(1):164– 165, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228757.

Woolf:1963:BRBa

[416] Harry Woolf. Book review: The Principal Works. Volume III Astronomy and Navigation by Simon Stevin; A. Pannekoek; Ernst Crone. Isis, 54 (1):165–167, March 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228758.

Ospovat:1963:BRB

[417] Alexander M. Ospovat. Book review: On the External Characters of Minerals by A. G. Werner; Albert V. Carozzi. Isis, 54(1):167–168, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228759.

Carlson:1963:BRB

[418] Robert E. Carlson. Book review: The Heroic Age of American Invention by L. Sprague De Camp. Isis, 54(1):168–169, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228760.

Ackerknecht:1963:BRB

[419] Erwin H. Ackerknecht. Book review: Theodor Schwann, Leben und Werk by Rembert Watermann; Lettres de Theodor Schwann by Marcel Florkin. Isis, 54(1):169–170, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228761.

Craven:1963:BRB

[420] Avery Craven. Book review: An Essay on Calcareous Manures by Edmund Ruffin; J. Carlyle Sitterson. Isis, 54(1):170, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228762.

Brozek:1963:BRBa

[421] Josef Brožek. Book review: Psychologist at Large. An Autobiography and Selected Essays by Edwin G. Boring. Isis, 54(1):170–172, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228763.

Anonymous:1963:BMa

[422] Anonymous. Back matter. Isis, 54(1):173–174, March 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228764.

Anonymous:1963:FMb

[423] Anonymous. Front matter. Isis, 54(2):175–177, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228536.

Bochner:1963:SSB

[424] Salomon Bochner. The significance of some basic mathematical conceptions for physics. Isis, 54(2):179–205, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228537.

Bernstein:1963:JCM

[425] Henry T. Bernstein. J. Clerk Maxwell on the history of the kinetic theory of gases, 1871. Isis, 54(2):206–216, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228538.

Nash:1963:CBP

[426] Gerald D. Nash. The conflict between pure and applied science in Nineteenth-Century public policy: The California State Geological Survey, 1860–1874. Isis, 54(2):217–228, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228539.

Pingree:1963:AAI

[427] David Pingree. Astronomy and astrology in India and Iran. Isis, 54(2): 229–246, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228540.

Siegfried:1963:DPS

[428] Robert Siegfried. The discovery of potassium and sodium, and the problem of the chemical elements. Isis, 54(2):247–258, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228541.

Babini:1963:EJR

[429] Jose Babini. Eloge: Julio Rey Pastor, 1888–1962. Isis, 54(2):259, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228542.

Drabkin:1963:TVG

[430] I. E. Drabkin. Two versions of G. B. Benedetti’s Demonstratio Proportionum Motuum Localium. (Latin) [… Demonstration of the capacity of local motion]. Isis, 54(2):259–262, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228543.

Mendoza:1963:SCC

[431] E. Mendoza. Sadi Carnot and the Cagnard engine. Isis, 54(2):262– 263, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228544. See [210].

Hall:1963:SMP

[432] A. Rupert Hall and James A. Ruffner. A strange mistranslation in the Principia. Isis, 54(2):263–264, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228545.

Aiton:1963:GEM

[433] E. J. Aiton. On Galileo and the Earth–Moon system. Isis, 54(2):265– 266, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228546.

Sabra:1963:NBV

[434] A. I. Sabra. Newton and the “bigness” of vibrations. Isis, 54(2):267– 268, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228547.

Thorndike:1963:TOP

[435] Lynn Thorndike. Two other passages de complexionibus. Isis, 54(2): 268–269, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228548.

Thorndike:1963:MMT

[436] Lynn Thorndike. More manuscripts of Thomas of Cantimpre, De Naturis Rerum. Isis, 54(2):269–277, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 228549. See correction [464].

Anonymous:1963:NPb

[437] Anonymous. News of the profession. Isis, 54(2):278–282, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228550.

Poynter:1963:BRB

[438] F. N. L. Poynter. Book review: Serial Publications Containing Medical Classics by Lee Ash. Isis, 54(2):283, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228551.

Reingold:1963:BRB

[439] Nathan Reingold. Book review: A History of Scientific and Technical Periodicals. The Origins and Development of the Scientific and Technological Press, 1665–1790 by David A. Kronick. Isis, 54(2):284–285, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228552.

Woolf:1963:BRBb

[440] Harry Woolf. Book review: Articles on Antiquity in Festschriften. The Ancient Near East, the Old Testament, Greece, Rome, Roman Law, Byzantium. An Index by Dorothy Rounds. Isis, 54(2):285, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228553.

Hesse:1963:BRB

[441] Mary Hesse. Book review: The Structure of Scientific Revolutions by Thomas S. Kuhn. Isis, 54(2):286–287, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228554.

Toulmin:1963:BRB

[442] Stephen Toulmin. Book review: Essay on Atomism: From Democritus to 1960 by Lancelot Law Whyte. Isis, 54(2):287–288, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228555.

Rosen:1963:BRB

[443] Edward Rosen. Book review: Die Tat des Kopernikus: Die Wandlung des Astronomischen Weltbildes im 16. und 17. Jahrhundert by Gerhard Harig. Isis, 54(2):288–289, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228556.

Truesdell:1963:BRB

[444] C. Truesdell. Book review: Concepts of Mass in Classical and Modern Physics by Max Jammer. Isis, 54(2):290–291, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228557.

Boyer:1963:BRBb

[445] Carl B. Boyer. Book review: Historia de la fisica: Desde la antiguedad hasta los umbrales del siglo XX by Desiderio Papp. Isis, 54(2):292– 293, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228558.

Coleman:1963:BRBb

[446] William Coleman. Book review: Histoire de la zoologie des origines a Linne by G. Petit; J. Theodorides. Isis, 54(2):293–295, June 1963.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228559.

Brozek:1963:BRBb

[447] Josef Brožek. Book review: Histoire de la psychologie de l’antiquite a nos jours by F. L. Mueller. Isis, 54(2):295–296, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228560.

Barber:1963:BRB

[448] Bernard Barber. Book review: Diffusion of Innovations by Everett M. Rogers. Isis, 54(2):296–297, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228561.

Solmsen:1963:BRB

[449] Friedrich Solmsen. Book review: Vom Chaos zum Kosmos. Zur Geschichte einer Idee by Franz Lammli. Isis, 54(2):297–299, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228562.

Merlan:1963:BRB

[450] Philip Merlan. Book review: Die Bedeutung der Bewegungslehre des Aristoteles fur seine beiden Losungen der zenonischen paradoxie by Matthias Schramm. Isis, 54(2):299–300, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228563.

deSantillana:1963:BRB

[451] Giorgio de Santillana. Book review: The School of Padua and the Emergence of Modern Science by John Herman Randall. Isis, 54(2):300–302, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228564.

Kish:1963:BRB

[452] George Kish. Book review: Correspondance Mercatorienne by M. Van Durme. Isis, 54(2):302–303, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228565.

Whiteside:1963:BRB

[453] D. T. Whiteside. Book review: Un savant méconnu: Gilles Personne de Roberval (1602–1675): Son activité intellectuelle dans les domaines mathématique, physique, mécanique et philosophique by Leon Auger. Isis, 54(2):303–305, June 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228566.

Williams:1963:BRB

[454] L. Pearce Williams. Book review: Histoire generale des sciences. Tome III: La science contemporaine. Volume I: Le XIXe siècle by Rene Taton. Isis, 54(2):305–306, June 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228567.

Rescher:1963:BRBa

[455] Nicholas Rescher. Book review: The Interpretation of Science: Selected Essays by Alfred North Whitehead; A. H. Johnson. Isis, 54(2):306– 307, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228568.

Anonymous:1963:BMb

[456] Anonymous. Back matter. Isis, 54(2):308–313, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228569.

Anonymous:1963:FMc

[457] Anonymous. Front matter. Isis, 54(3):315–317, September 1963. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228802.

Cohen:1963:PTN

[458] I. Bernard Cohen and H. Pemberton. Pemberton’s translation of Newton’s principia, with notes on Motte’s translation. Isis, 54(3):319–351, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228803.

Clark:1963:PGP

[459] Joseph T. Clark. Pierre Gassendi and the physics of Galileo. Isis, 54 (3):352–370, September 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228804.

Vorzimmer:1963:CDB

[460] Peter Vorzimmer. Charles Darwin and blending inheritance. Isis, 54 (3):371–390, September 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228805.

Jayawardene:1963:UDR

[461] S. A. Jayawardene and Rafael Bombelli. Unpublished documents relating to Rafael Bombelli in the Archives of Bologna. Isis, 54(3):391–395, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228806.

Sinclair:1963:UAA

[462] Keith V. Sinclair. An unnoticed astronomical and astrological manuscript. Isis, 54(3):396–399, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228807.

Burstyn:1963:GEM

[463] Harold L. Burstyn. Galileo and the Earth–Moon system: Reply to dr. aiton. Isis, 54(3):400–401, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228808.

Anonymous:1963:CMM

[464] Anonymous. Correction: More manuscripts of Thomas of Cantimpre, De Naturis Rerum. Isis, 54(3):401, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228809. See [436].

Anonymous:1963:NPc

[465] Anonymous. News of the profession. Isis, 54(3):402–403, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228810.

Drake:1963:BRB

[466] Stillman Drake. Book review: The Achievement of Galileo by James Brophy; Henry Paolucci. Isis, 54(3):404, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228811.

Agassi:1963:BRB

[467] J. Agassi. Book review: Logic, Methodology and Philosophy of Science by Ernest Nagel; Patrick Suppes; Alfred Tarski. Isis, 54(3):405–407, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228812.

Whitrow:1963:BRB

[468] G. J. Whitrow. Book review: Time and the Physical World by Richard Schlegel. Isis, 54(3):407–408, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228813.

Achinstein:1963:BRB

[469] Peter Achinstein. Book review: Foresight and Understanding by Stephen Toulmin. Isis, 54(3):408–410, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228814.

Schlegel:1963:BRB

[470] Richard Schlegel. Book review: The Natural Philosophy of Time by G. J. Whitrow. Isis, 54(3):410–411, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228815.

Kainen:1963:BRB

[471] Jacob Kainen. Book review: Art and the Scientist by Geoffrey Lapage. Isis, 54(3):411–412, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228816.

Allen:1963:BRBb

[472] D. C. Allen. Book review: Seventeenth Century Science and the Arts by Stephen Toulmin; Douglas Bush; James S. Ackerman; Claude V. Palisca; Hedley Howell Rhys. Isis, 54(3):412, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228817.

Hanson:1963:BRB

[473] Norwood Russell Hanson. Book review: The Discovery of Neptune by Morton Grosser. Isis, 54(3):413–414, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228818.

Graff-Petersen:1963:BRB

[474] Poul Graff-Petersen. Book review: Geologiens Historie i Danmark by Axel Garboe. Isis, 54(3):414–415, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228819.

Mendelsohn:1963:BRB

[475] Everett Mendelsohn. Book review: A History of Biology to about the Year 1900. A General Introduction to the Study of Living Things by Charles Singer. Isis, 54(3):415–417, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228820.

Hughes:1963:BRB

[476] Thomas P. Hughes. Book review: A Short History of Technology: From the Earliest Times to A.D. 1900 by T. K. Derry; Trevor I. Williams. Isis, 54(3):417–418, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228821.

Needham:1963:BRB

[477] Joseph Needham. Book review: Medieval Technology and Social Change by Lynn White. Isis, 54(3):418–420, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228822.

Youschkevitch:1963:BRB

[478] A. P. Youschkevitch. Book review: Al-Kashi on Root Extraction by Abdul-Kader Dakhel; Wasfi A. Hijab; E. S. Kennedy. Isis, 54(3):420–421, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228823.

VonGrunebaum:1963:BRB

[479] G. E. Von Grunebaum. Book review: The Travels of Ibn Battuta, A.D. 1325–1354. Volume II by Hamilton Gibb. Isis, 54(3):421, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228824.

Stringa:1963:BRB

[480] Marialuisa Stringa. Book review: Scetticismo ed emptrismo, studio su Gassendi by Tullio Gregory. Isis, 54(3):421–424, September 1963. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228825.

Lorch:1963:BRB

[481] J. Lorch. Book review: Alle de brieven van Antoni van Leeuwenhoek (The Collected Letters). Volume VI. 1686–1687 by Antoni Van Leeuwenhoek; Dutch scientists. Isis, 54(3):424–426, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228826.

Dieke:1963:BRB

[482] G. H. Dieke. Book review: Pierre Lyonet. 1706–1789: Sa vie, ses collections de coquillages et de tableaux, ses recherches entomologiques by W. H. Van Seters. Isis, 54(3):426–427, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228827.

Burstyn:1963:BRB

[483] Harold L. Burstyn. Book review: Roger Joseph Boscovich, S.J., F.R.S., 1711–1787. Studies of His Life and Work on the 250th Anniversary of His Birth by Lancelot Law Whyte. Isis, 54(3):428–430, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228828.

Wassermann:1963:BRB

[484] Felix M. Wassermann. Book review: Alexander von Humboldt. Band II: Vom Reisewerk zum Kosmos by Hanno Beck. Isis, 54(3):430–431, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228829.

Almgren:1963:BRB

[485] Beverly S. Almgren. Book review: Dmitrii Ivanovich Mendeleev, 1834– 1907 by N. A. Figurovskii. Isis, 54(3):431–433, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228830.

Miller:1963:BRBa

[486] Genevieve Miller. Book review: Daniel Drake (1785–1852): Pioneer Physician of the Midwest by Emmet Field Horine. Isis, 54(3):433, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228831.

Leicester:1963:BRBb

[487] Henry M. Leicester. Book review: Centenary of the Theory of Chemical Structure. Collection of Papers by A. M. Butlerov, A. S. Couper, A. Kekule, and V. V. Markovnikov by B. A. Kazansky; G. V. Bykov. Isis, 54 (3):433–434, September 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228832.

Woodruff:1963:BRB

[488] A. E. Woodruff. Book review: Wilhelm Webers Stellung in der Entwichlung der Elektrizitatslehre by Karl Heinrich Wiederkehr. Isis, 54 (3):434–435, September 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228833.

Anonymous:1963:BMc

[489] Anonymous. Back matter. Isis, 54(3):436–437, September 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228834.

Anonymous:1963:FMd

[490] Anonymous. Front matter. Isis, 54(4):439–514, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228149.

Farber:1963:VPT

[491] Eduard Farber. Variants of preformation theory in the history of chemistry. Isis, 54(4):443–460, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228150.

Silliman:1963:WTS

[492] Robert H. Silliman. William Thomson: Smoke rings and Nineteenth-Century atomism. Isis, 54(4):461–474, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228151.

Morgan:1963:TNL

[493] John Morgan and Antonio Pace. Two new letters from John Morgan’s Italian travels. Isis, 54(4):475–479, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228152.

Siegfried:1963:FDD

[494] Robert Siegfried. Further Daltonian doubts. Isis, 54(4):480–481, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228153. See [208, 826].

Kemp:1963:CLR

[495] D. Alasdair Kemp. The Crawford Library of the Royal Observatory, Edinburgh. Isis, 54(4):481–483, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228154.

Thorndike:1963:WB

[496] Lynn Thorndike. William Bonkis. Isis, 54(4):483–484, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228155.

Anonymous:1963:NPd

[497] Anonymous. News of the profession. Isis, 54(4):485–486, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228156.

Feyerabend:1963:BRB

[498] Paul K. Feyerabend. Book review: The Origins of Science. An Inquiry into the Foundations of Western Thought by Ernest H. Hutten. Isis, 54 (4):487–488, December 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228157.

Vekerdi:1963:BRB

[499] L. Vekerdi. Book review: A magyarorszagi fizika tortenete 1711-ig by M. Zemplen Jolan. Isis, 54(4):488–490, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228158.

Boyer:1963:BRBc

[500] Carl B. Boyer. Book review: Les grands courants de la pensée mathematique by F. Le Lionnais. Isis, 54(4):490–492, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228159.

Leicester:1963:BRBc

[501] Henry M. Leicester. Book review: Historical Studies in the Language of Chemistry by M. P. Crosland. Isis, 54(4):492, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228160.

Papp:1963:BRB

[502] Desiderio Papp. Book review: Storia della fisica by Mario Gliozzi. Isis, 54 (4):492–494, December 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228161.

Grosser:1963:BRB

[503] Morton Grosser. Book review: A History of Astronomy by A. Pannekoek. Isis, 54(4):494–495, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228162.

Hall:1963:BRBb

[504] Marie Boas Hall. Book review: A History of Chemistry. Volume II by J. R. Partington. Isis, 54(4):495–497, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228163.

Miller:1963:BRBb

[505] Genevieve Miller. Book review: Early American Medical Imprints: A Guide to Works Printed in the United States. 1668–1820 by Robert B. Austin. Isis, 54(4):497, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228164.

Pascal:1963:BRB

[506] Paul Pascal. Book review: The Halieutica Ascribed to Ovid by Ovid; J. A. Richmond. Isis, 54(4):497–498, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228165.

Rescher:1963:BRBb

[507] Nicholas Rescher. Book review: The “Apeiron” of Anaximander. A Study in the Origin and Function of Metaphysical Ideas by Paul Seligman. Isis, 54(4):498–499, December 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228166.

Pagel:1963:BRB

[508] Walter Pagel. Book review: A History of Medicine. Volume II. Early Greek, Hindu, and Persian Medicine by Henry E. Sigerist. Isis, 54(4): 499–501, December 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228167.

Hermelink:1963:BRB

[509] Heinrich Hermelink. Book review: The Astronomical Works of Thabit B. Qurra by Francis J. Carmody. Isis, 54(4):501–503, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228168.

Pines:1963:BRB

[510] S. Pines. Book review: The Travels of Ibn Battuta, A.D. 1325–1354. Vol. I by Hamilton Gibb; C. Defremery; B. R. Sanguinetti. Isis, 54 (4):503–507, December 1963. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228169.

May:1963:BRB

[511] Margaret T. May. Book review: Preface d’Andre Vesale a ses livres sur l’anatomie, suivie d’une lettre a Jean Oporinus, son imprimeur by Andreas Vesalius; Louis Bakelants. Isis, 54(4):507, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228170.

Schofield:1963:BRB

[512] Robert Schofield. Book review: Joseph Priestley: Selections from His Writings by Ira V. Brown. Isis, 54(4):508–509, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228171.

Westfall:1963:BRBb

[513] Richard S. Westfall. Book review: The Correspondence of Isaac Newton. Vol. 3. 1688–1694 by Isaag Newton; H. W. Turnbull. Isis, 54(4):509–511, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228172.

Coleman:1963:BRBc

[514] William Coleman. Book review: Essay on Classification by Louis Agassiz; Edward Lurie. Isis, 54(4):511–512, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228173.

Neville:1963:BRB

[515] Roy G. Neville. Book review: Michel Eugene Chevreul: Pioneer of Organic Chemistry by Albert B. Costa. Isis, 54(4):512–513, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228174.

Hiebert:1963:BRBb

[516] Erwin N. Hiebert. Book review: The Russians in the Arctic: Aspects of Soviet Exploration and Exploitation of the Far North 1937–1957 by Terence Armstrong. Isis, 54(4):513, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228175.

Woolf:1963:EEC

[517] Harry Woolf and Aleida Cattell. Eighty-eighth critical bibliography of the history of science and its cultural influences (to 1 January 1963). Isis, 54(4):515–667, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228176.

Anonymous:1963:VI

[518] Anonymous. Volume information. Isis, 54(4):668–674, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228148.

Anonymous:1963:BMd

[519] Anonymous. Back matter. Isis, 54(4):??, December 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228177.

Anonymous:1964:FMa

[520] Anonymous. Front matter. Isis, 55(1):1–3, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227749.

Anonymous:1964:GSM

[521] Anonymous. The George Sarton Medal for 1963: Award of the George Sarton Medal to Vassili P. Zoubov (1899–1963). Isis, 55(1):4–6, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227750.

Finn:1964:LSS

[522] Bernard S. Finn. Laplace and the speed of sound. Isis, 55(1):7–19, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227751.

Graham:1964:SMV

[523] Loren R. Graham. A Soviet Marxist view of structural chemistry: The Theory of Resonance controversy. Isis, 55(1):20–31, March 1964. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227752.

Susskind:1964:OEW

[524] Charles Susskind. Observations of electromagnetic-wave radiation before Hertz. Isis, 55(1):32–42, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227753.

Debus:1964:PAN

[525] Allen G. Debus. The Paracelsian aerial niter. Isis, 55(1):43–61, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227754.

Rauschenberg:1964:LSJ

[526] Roy Rauschenberg and Joseph Banks. A letter of Sir Joseph Banks describing the life of Daniel Solander. Isis, 55(1):62–67, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227755.

Boyer:1964:CCT

[527] Carl B. Boyer. Clairaut le Cadet and a theorem of Thabit ibn Qurra. Isis, 55(1):68–70, March 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227756.

Dales:1964:ABP

[528] Richard C. Dales. The authorship of the Summa in Physica attributed to Robert Grosseteste. Isis, 55(1):70–74, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227757.

Kelly:1964:RDS

[529] Michael Kelly. Robert Darwin’s splendid character. Isis, 55(1):74–79, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227758.

Suter:1964:LXG

[530] Rufus Suter. Leonardo Ximenes and the gnomon at the Cathedral of Florence. Isis, 55(1):79–82, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227759.

Babini:1964:VBP

[531] Jose Babini. Valentin Balbin y la primera revista matematica Argentina. (spanish) [Valentin Balbin and the first journal Mathematics Argentina].

Isis, 55(1):82–85, March 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227760.

Roofe:1964:HPU

[532] Paul G. Roofe. The Herrick Papers, an unusual scientific archive. Isis, 55(1):85–87, March 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227761.

Barr:1964:BMP

[533] E. Scott Barr. Biographical material in the philosophical magazine to 1900. Isis, 55(1):88–90, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227762.

Rancke-Madsen:1964:RQN

[534] E. Rancke-Madsen. Reply to query no. 155. Isis, 55(1):91, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227763.

Anonymous:1964:NP

[535] Anonymous. News of the profession. Isis, 55(1):92–96, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227764.

Gillispie:1964:BRB

[536] Charles C. Gillispie. Book review: Towards an Historiography of Science by Joseph Agassi. Isis, 55(1):97–99, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227765.

Hujer:1964:BRB

[537] Karel Hujer. Book review: History of Exact Sciences in Bohemian Lands Up to the End of the Nineteenth Century by J. Folta; Z. Horsky; L. Novy; I. Seidlerova; J. Smolka; M. Teich. Isis, 55(1):99–101, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227766.

Briggs:1964:BRB

[538] J. Morton Briggs, Jr. Book review: A History of Science and Technology by R. J. Forbes; E. J. Dijksterhuis. Isis, 55(1):101–102, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227767.

Williams:1964:BRB

[539] L. Pearce Williams. Book review: The Architecture of Matter by Stephen Toulmin; June Goodfield. Isis, 55(1):102–104, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227768.

Partridge:1964:BRB

[540] M. F. Partridge. Book review: Method in the Physical Sciences by G. Schlesinger. Isis, 55(1):104–106, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227769.

Leicester:1964:BRBa

[541] Henry M. Leicester. Book review: A History of Chemistry. Vol. III by J. R. Partington. Isis, 55(1):106–107, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227770.

Pagel:1964:BRB

[542] Walter Pagel. Book review: Catalogue of the Wellcome Historical Medical Library. I. Books Printed before 1641. Isis, 55(1):107–109, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227771.

Cowen:1964:BRB

[543] David L. Cowen. Book review: History of Pharmacy in Britain by Leslie G. Matthews. Isis, 55(1):109–110, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227772.

Grant:1964:BRB

[544] Edward Grant. Book review: The Origins of Scientific Thought from Anaximander to Proclus, 600 B.C.–500 A.D. by Giorgio de Santillana. Isis, 55(1):110–111, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227773.

Sambursky:1964:BRB

[545] S. Sambursky. Book review: Roman Science. Origins, Development, and Influence to the Later Middle Ages by William H. Stahl. Isis, 55(1):111– 113, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227774.

Condit:1964:BRB

[546] Carl W. Condit. Book review: The Construction of Gothic Cathedrals: A Study of Medieval Vault Erection by John Fitchen. Isis, 55(1):113– 115, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227775.

Alexander:1964:BRB

[547] Paul J. Alexander. Book review: La Topographie Chretienne de Cosmas Indicopleustes. Theologie et science au VIe siècle by Wanda Wolska. Isis, 55(1):115–117, March 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227776.

Rudwick:1964:BRB

[548] M. J. S. Rudwick. Book review: The Lying Stones of Dr. Johann Bartholomew Adam Beringer Being His “Lithographiae Wirceburgensis” by Johann Bartholomew Adam Beringer; Melvin E. Jahn; Daniel J. Woolf. Isis, 55(1):117–118, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227777.

Cope:1964:BRBa

[549] Jackson I. Cope. Book review: Henry More. The Rational Theology of a Cambridge Platonist by Aharon Lichtenstein. Isis, 55(1):118–119, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227778.

Wightman:1964:BRB

[550] William P. D. Wightman. Book review: Isaac Newton, Historian by Frank E. Manuel. Isis, 55(1):119–120, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227779.

Oppenheimer:1964:BRB

[551] Jane Oppenheimer. Book review: Les Sciences de la vie dans la pensée française du XVIIIe siècle. La generation des animaux de Descartes a l’Encyclopedie by Jacques Roger. Isis, 55(1):121–122, March 1964. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227780.

Holmes:1964:BRBa

[552] Frederic L. Holmes. Book review: Theodor Boveri. Leben und Werk eines grossen Biologen, 1862–1915 by Fritz Baltzer. Isis, 55(1):122–123, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227781.

Anonymous:1964:BMa

[553] Anonymous. Back matter. Isis, 55(1):124–125, March 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227782.

Anonymous:1964:FMb

[554] Anonymous. Front matter. Isis, 55(2):127–129, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228178.

Multhauf:1964:E

[555] Robert P. Multhauf. Editorial. Isis, 55(2):131–133, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228179.

Stocking:1964:FA

[556] George W. Stocking, Jr. French anthropology in 1800. Isis, 55(2):134– 150, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228180.

Buchdahl:1964:TCW

[557] Gerd Buchdahl. Theory construction: The work of Norman Robert Campbell. Isis, 55(2):151–162, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228181.

Grosser:1964:SPB

[558] Morton Grosser. The search for a Planet beyond Neptune. Isis, 55(2): 163–183, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228182.

Kargon:1964:WCR

[559] Robert Kargon. Walter Charleton, Robert Boyle, and the acceptance of epicurean atomism in England. Isis, 55(2):184–192, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228183.

Goodrich:1964:ECC

[560] L. Carrington Goodrich. Early cannon in China. Isis, 55(2):193–195, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228184.

Geus:1964:LBS

[561] Armin Geus. Leonhard Baldner, a Strasbourg fisherman. Isis, 55(2): 195–199, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228185.

Poulle:1964:BVT

[562] Emmanuel Poulle and Bernard de Verdun. Bernard de Verdun et le turquet. (French) [Bernard de Verdun and the ????]. Isis, 55(2):200– 208, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228186.

Anonymous:1964:Na

[563] Anonymous. News. Isis, 55(2):209–212, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228187.

Glass:1964:BRB

[564] Bentley Glass. Book review: Classics in Biology. A Course of Selected Reading by Authorities by S. Zuckerman; Classics in Science. A Course of Selected Reading by Authorities by E. N. Da C. Andrade. Isis, 55(2): 213–214, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228188.

Temkin:1964:BRB

[565] Owsei Temkin. Book review: Catalogue of Western Manuscripts on Medicine and Science in the Wellcome Historical Medical Library. Volume I. Mss. Written before 1650 A.D. by S. A. J. Moorat. Isis, 55(2): 214–215, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228189.

Salmon:1964:BRB

[566] Wesley C. Salmon. Book review: Frontiers of Science and Philosophy by Robert G. Colodny. Isis, 55(2):215–216, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228190.

Scriba:1964:BRB

[567] Christoph J. Scriba. Book review: Games, Gods and Gambling by F. N. David. Isis, 55(2):216–218, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228191.

Diamandopoulos:1964:BRB

[568] Peter Diamandopoulos. Book review: The Rainbow: From Myth to Mathematics by Carl B. Boyer. Isis, 55(2):218–220, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228192.

Watanabe:1964:BRBa

[569] Masao Watanabe. Book review: Count Rumford, Physicist Extraordinary by Sanborn C. Brown. Isis, 55(2):220–221, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228193.

Kennedy:1964:BRB

[570] E. S. Kennedy. Book review: Solar and Planetary Longitudes for Years −2500 to +2000 by William D. Stahlman; Owen Gingerich. Isis, 55(2): 221–222, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228194.

Burke:1964:BRB

[571] John G. Burke. Book review: The Mercurial Chemist. A Life of Sir Humphry Davy by Anne Treneer. Isis, 55(2):222–223, June 1964. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228195.

Wilson:1964:BRB

[572] Leonard G. Wilson. Book review: Georges Cuvier, Zoologist. A Study in the History of Evolution Theory by William Coleman. Isis, 55(2):223– 224, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228196.

Berkeley:1964:BRBa

[573] Edmund Berkeley. Book review: “John Lyon, Nurseryman and Plant Hunter, and His Journal, 1799–1814” by Joseph Ewan; Nesta Ewan. Isis, 55(2):224–225, June 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228197.

Dunn:1964:BRB

[574] L. C. Dunn. Book review: Animal Species and Evolution by Ernst Mayr. Isis, 55(2):225–227, June 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228198.

Holmes:1964:BRBb

[575] Frederic L. Holmes. Book review: The Growth of Medical Thought by Lester S. King. Isis, 55(2):227–228, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228199.

White:1964:BRB

[576] Lynn White, Jr. Book review: Histoire generale des techniques. Tome I. Les origines de la civilisation technique by Maurice Daumas. Isis, 55(2): 228–230, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228200.

Hanson:1964:BRB

[577] Norwood Russell Hanson. Book review: The American Heritage History of Flight by Alvin M. Josephy. Isis, 55(2):230–233, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228201.

Watanabe:1964:BRBb

[578] Masao Watanabe. Book review: Historical Development of Science and Technology in Japan by Hideomi Tuge. Isis, 55(2):233–234, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228202.

Aitchison:1964:BRB

[579] Leslie Aitchison. Book review: The Coming of the Age of Steel by Theodore A. Wertime. Isis, 55(2):234–235, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228203.

Coghlan:1964:BRB

[580] H. H. Coghlan. Book review: Metallurgy in Archaeology. A Prehistory of Metallurgy in the British Isles by R. F. Tylecote. Isis, 55(2):235– 236, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228204.

Stahl:1964:BRB

[581] William H. Stahl. Book review: The Mechanical Technology of Greek and Roman Antiquity by A. G. Drachmann. Isis, 55(2):236–237, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228205.

Philip:1964:BRB

[582] J. A. Philip. Book review: Parmenides, Melissus, Gorgias. A Reinterpretation of Eleatic Philosophy by J. H. M. M. Loenen. Isis, 55(2):237–238, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228206.

Beaujouan:1964:BRB

[583] Guy Beaujouan. Book review: Las tablas astronomicas del Rey Don Pedro el Ceremonioso by Jose M. Millas-Vallicrosa. Isis, 55(2):238–239, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228207.

Rath:1964:BRB

[584] Gernot Rath. Book review: Das medizinische Weltbild des Paracelsus. Seine Zusammen hange mit Neuplatonismus und Gnosis by Walter Pagel. Isis, 55(2):239–240, June 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228208.

Cope:1964:BRBb

[585] Jackson I. Cope. Book review: Theodore Haak (1605–1690): The First German Translator of Paradise Lost by Pamela R. Barnett. Isis, 55(2): 241, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228209.

Hall:1964:BRB

[586] Marie Boas Hall. Book review: Francis Bacon and the Modern Dilemma by Loren Eiseley. Isis, 55(2):241–242, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228210.

Hellman:1964:BRBa

[587] C. Doris Hellman. Book review: Gesammelte Werke. Band 9: Mathematische Schriften by Johannes Kepler. Isis, 55(2):242–243, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228211.

Hellman:1964:BRBb

[588] C. Doris Hellman. Book review: An Account of the Astronomical Discoveries of Kepler by Robert Small. Isis, 55(2):243–245, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228212.

Boyer:1964:BRB

[589] Carl B. Boyer. Book review: Algebre et logique d’après les textes originaux de G. Boole et W. S. Jevons avec les plans de la machine logique by G. Boole; W. S. Jevons; Frederic Gillot. Isis, 55(2):245–246, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228213.

Schofield:1964:BRB

[590] Robert Schofield. Book review: American and British Technology in the Nineteenth Century. The Search for Labour-Saving Inventions by H. J. Habakkuk. Isis, 55(2):246–247, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228214.

Graham:1964:BRBa

[591] Loren Graham. Book review: Istoriia kristallografii v Rossii by I. I. Shafranovskii. Isis, 55(2):248–249, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228215.

McIlwain:1964:BRB

[592] H. McIlwain. Book review: A Treatise on the Chemical Constitution of the Brain by J. L. W. Thudichum. Isis, 55(2):249–250, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228216.

Ward:1964:BRB

[593] J. C. Ward. Book review: The Concept of the Positron. A Philosophical Analysis by Norwood Russell Hanson. Isis, 55(2):250–251, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228217.

Graham:1964:BRBb

[594] Loren Graham. Book review: Einshtein (Einstein) by B. G. Kuznetsov. Isis, 55(2):251–252, June 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228218.

Anonymous:1964:BMb

[595] Anonymous. Back matter. Isis, 55(2):253–259, June 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228219.

Anonymous:1964:FMc

[596] Anonymous. Front matter. Isis, 55(3):261–263, September 1964. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228570.

Grant:1964:MVP

[597] Edward Grant. Motion in the void and the principle of inertia in the middle ages. Isis, 55(3):265–292, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228571.

Carozzi:1964:LTE

[598] Albert V. Carozzi. Lamarck’s theory of the Earth: Hydrogéologie. Isis, 55 (3):293–307, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228572.

See comments [751].

Hanson:1964:CED

[599] Norwood Russell Hanson. Contra-equivalence a defense of the originality of Copernicus. Isis, 55(3):308–325, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228573.

Bork:1964:FDN

[600] Alfred M. Bork. The fourth dimension in Nineteenth-Century physics. Isis, 55(3):326–338, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228574.

Goodfield:1964:QPD

[601] June Goodfield and Stephen Toulmin. The qattara: A primitive distillation and extraction apparatus still in use. Isis, 55(3):339–342, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228575.

Taylor:1964:RGO

[602] E. G. R. Taylor. Robert Grosseteste as an observer. Isis, 55(3):342, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228576.

Pines:1964:SDB

[603] S. Pines. The semantic distinction between the terms astronomy and astrology according to al-Biruni. Isis, 55(3):343–349, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228577.

Goodfield:1964:GSP

[604] June Goodfield and Leonard G. Wilson. The growth of scientific physiology. Isis, 55(3):349–351, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228578.

Anonymous:1964:QDB

[605] Anonymous. Query 171 Darwin–Bates correspondence. Isis, 55(3):352, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228579.

Stannard:1964:FCB

[606] Jerry Stannard. A Fifteenth-Century botanical glossary. Isis, 55(3): 353–367, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228580.

Anonymous:1964:Nb

[607] Anonymous. News. Isis, 55(3):368–369, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228581.

Anonymous:1964:CNP

[608] Anonymous. Correction: News of the profession. Isis, 55(3):369, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228582.

Reingold:1964:BRB

[609] Nathan Reingold. Book review: The Physical Geography of the Sea and Its Meteorology. Isis, 55(3):370–372, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228583.

Graham:1964:BRBc

[610] Loren R. Graham. Book review: Science in Russian Culture. A History to 1860 by Alexander Vucinich. Isis, 55(3):373–374, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228584.

Daub:1964:BRB

[611] Edward E. Daub. Book review: Science in the History of Modern Culture by Masao Watanabe. Isis, 55(3):374–375, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228585.

Wassermann:1964:BRB

[612] Felix M. Wassermann. Book review: Science and Education by Thomas Huxley. Isis, 55(3):375–376, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228586.

Hellman:1964:BRBc

[613] C. Doris Hellman. Book review: Watchers of the Skies. An Informal History of Astronomy from Babylon to the Space Age by Willy Ley. Isis, 55 (3):376–377, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228587.

Stern:1964:BRB

[614] Curt Stern. Book review: Kurze Geschichte der Genetik bis zur Wiederentdeckung der Vererbungsregeln Gregor Mendels. Beitrag 1. Genetik. Grundlagen, Ergebnisse und Probleme in Einzeldarstellungen. by Hans Stubbe. Isis, 55(3):377–379, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228588.

Brieger:1964:BRB

[615] Gert H. Brieger. Book review: Discoverers of Blood Circulation: From Aristotle to the Times of da Vinci and Harvey by Tibor Doby. Isis, 55(3):379, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228589.

Dibner:1964:BRBa

[616] Bern Dibner. Book review: The Ancient Engineers by L. Sprague de Camp. Isis, 55(3):380–382, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228590.

Dibner:1964:BRBb

[617] Bern Dibner. Book review: The Making of the Electrical Age. From the Telegraph to Automation by Harold I. Sharlin. Isis, 55(3):382–383, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228591.

Solmsen:1964:BRB

[618] Friedrich Solmsen. Book review: A Portrait of Aristotle by Marjorie Grene. Isis, 55(3):383–385, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228592.

Drabkin:1964:BRB

[619] I. E. Drabkin. Book review: The Physical World of Late Antiquity by S. Sambursky. Isis, 55(3):385–386, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228593.

Bullough:1964:BRB

[620] Vern L. Bullough. Book review: The Salernitan Questions. An Introduction to the History of Medieval and Renaissance Problem Literature by Brian Lawn. Isis, 55(3):386–388, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228594.

Zirkle:1964:BRBa

[621] Conway Zirkle. Book review: Aldrovandi on Chickens. The Ornithology of Ulisse Aldrovandi (1600). Vol. II, Book XIV by L. R. Lind. Isis, 55 (3):388–389, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228595.

Debus:1964:BRB

[622] Allen G. Debus. Book review: Giordano Bruno and the Hermetic Tradition by Frances A. Yates. Isis, 55(3):389–391, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228596.

Stannard:1964:BRB

[623] Jerry Stannard. Book review: John Clayton, Pioneer of American Botany by Edmund Berkeley; Dorothy Smith Berkeley. Isis, 55(3):391– 392, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228597.

Rappaport:1964:BRB

[624] Rhoda Rappaport. Book review: Memoires du Museum national d’Histoire naturelle. Serie C: Sciences de la terre. Tome X: Buffon, Les Epoques de la nature by Jacques Roger. Isis, 55(3):392–394, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228598.

Hoch:1964:BRB

[625] J. Hampton Hoch. Book review: Flowers for the King. Ruiz and Pavon and the Flora of Peru by Arthur Robert Steele. Isis, 55(3):394, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228599.

Farber:1964:BRB

[626] Eduard Farber. Book review: Lomonosov, Schlozer, Pallas, Deutsch– russiche Wissenschaftsbeziehungen im 18. Jahrhundert by E. Winter. Isis, 55(3):394–396, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228600.

Leicester:1964:BRBb

[627] Henry M. Leicester. Book review: Justus von Liebig in eigenen Zeugnissen und solchen seiner Zeitgenossen by Hertha von Dechend. Isis, 55(3):396, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228601.

Berkeley:1964:BRBb

[628] Edmund Berkeley. Book review: Correspondence between Spencer Fullerton Baird and Louis Agassiz–Two Pioneer American Naturalists by Elmer Charles Herber. Isis, 55(3):397, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228602.

Jackson:1964:BRB

[629] W. Turrentine Jackson. Book review: The Western Journals of Dr. George Hunter, 1796–1805 by John Francis McDermott. Isis, 55(3): 398–399, September 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228603.

Zirkle:1964:BRBb

[630] Conway Zirkle. Book review: A Professor’s Story of Hybrid Corn by Herbert K. Hayes. Isis, 55(3):399–400, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228604.

Bork:1964:BRB

[631] Alfred M. Bork. Book review: The Electron. Its Isolation and Measurement and the Determination of Some of Its Properties by Robert Andrews Millikan; Jesse W. M. DuMond. Isis, 55(3):400–401, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228605.

Anonymous:1964:BMc

[632] Anonymous. Back matter. Isis, 55(3):402–403, September 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228606.

Anonymous:1964:FMd

[633] Anonymous. Front matter. Isis, 55(4):405–474, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228310.

Gorman:1964:GA

[634] Mel Gorman. Gassendi in America. Isis, 55(4):409–417, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228311.

Brampton:1964:OHA

[635] C. K. Brampton. Ockham and his authorship of the summulae in libros physicorum. Isis, 55(4):418–426, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228312.

Apmann:1964:CHT

[636] Robert P. Apmann. A case history in theory and experiment: Fluid flow in bends. Isis, 55(4):427–434, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228313.

Kennedy:1964:CUC

[637] E. S. Kennedy. The Chinese–Uighur calendar as described in the Islamic sources. Isis, 55(4):435–443, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228314.

Anonymous:1964:Nc

[638] Anonymous. News. Isis, 55(4):444–445, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228315.

Cannon:1964:BRB

[639] Walter F. Cannon. Book review: Charles Darwin, by Gavin de Beer. Evolution by Natural Selection. Isis, 55(4):446–447, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228316.

Achinstein:1964:BRB

[640] Peter Achinstein. Book review: Sensationalism and Scientific Explanation by Peter Alexander. Isis, 55(4):448–449, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228317.

Post:1964:BRB

[641] Heinz R. Post. Book review: Philosophical Problems of Space and Time by Adolf Grunbaum. Isis, 55(4):449–450, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228318.

Harrison:1964:BRB

[642] Louis P. Harrison. Book review: The History of the Barometer by W. E. Knowles Middleton. Isis, 55(4):450–451, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228319.

Heyden:1964:BRB

[643] Francis J. Heyden. Book review: William Herschel by Angus Armitage. Isis, 55(4):452, December 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228320.

Buttmann:1964:BRB

[644] Gunther Buttmann. Book review: William Herschel and the Construction of the Heavens by Michael A. Hoskin. Isis, 55(4):452–454, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228321.

Stocking:1964:BRB

[645] George W. Stocking, Jr. Book review: Early Anthropology in the Sixteenth and Seventeenth Centuries by Margaret T. Hodgen. Isis, 55(4): 454–455, December 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228322.

Huard:1964:BRB

[646] P. Huard. Book review: The Scientific Evolution of Psychology. Vol. I by J. R. Kantor. Isis, 55(4):455–457, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228323.

Ackerknecht:1964:BRB

[647] Erwin H. Ackerknecht. Book review: A History of American Pathology by Esmond R. Long. Isis, 55(4):457–458, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228324.

Chipman:1964:BRB

[648] Robert Chipman. Book review: Popov and the Beginnings of Radiotelegraphy by Charles Susskind; The Invention of the Traveling-Wave Tube by Rudolf Kompfner; On the Early History of Radio Guidance by Benjamin Franklin Miessner. Isis, 55(4):458–459, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228325.

vonFritz:1964:BRB

[649] Kurt von Fritz. Book review: Weisheit und Wissenschaft. Studien zu Pythagoras, Philolaos und Platon by Walter Burkert. Isis, 55(4):459– 461, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228326.

Durling:1964:BRB

[650] Richard J. Durling. Book review: On the Passions and Errors of the Soul by Galen; Paul W. Harkins. Isis, 55(4):461–463, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228327.

Toomer:1964:BRB

[651] G. J. Toomer. Book review: Ibn al-Haythams Weg zur Physik by Matthias Schramm. Isis, 55(4):463–465, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228328.

Grew:1964:BRB

[652] Raymond Grew. Book review: Tradition and Enlightenment in the Tuscan Academies, 1690–1800 by Eric W. Cochrane. Isis, 55(4):465–466, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228329.

Schmidl:1964:BRB

[653] Fritz Schmidl. Book review: Studies in the Prehistory of Psychoanalysis. The Etiology of Psychoneuroses and Some Related Themes in Sigmund Freud’s Scientific Writings and Letters, 1886–1896 by Ola Andersson. Isis, 55(4):467, December 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228330.

Coleman:1964:BRB

[654] William Coleman. Book review: Hydrogeology by J. B. Lamarck; A. V. Carozzi. Isis, 55(4):468, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228331.

Ritterbush:1964:BRB

[655] Philip C. Ritterbush. Book review: The Mechanistic Conception of Life by Jacques Loeb; Donald Fleming. Isis, 55(4):468–469, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228332.

Spitzer:1964:BRB

[656] Alan B. Spitzer. Book review: European Positivism in the Nineteenth Century. An Essay in Intellectual History by W. M. Simon. Isis, 55 (4):469–470, December 1964. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228333.

Mendelsohn:1964:BRB

[657] Everett Mendelsohn. Book review: The Impact of Atomic Energy. A History of Responses of Government, Scientists and Religious Groups by Erwin N. Hiebert. Isis, 55(4):470–472, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228334.

Multhauf:1964:ENC

[658] Robert P. Multhauf, Bernard S. Finn, Aleida v. H. Cattell, and Diana D. Menkes. Eighty-ninth critical bibliography of the history of science and its cultural influences (to 1 January 1964). Isis, 55(4):475–605, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228335.

Anonymous:1964:VI

[659] Anonymous. Volume information. Isis, 55(4):606–611, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228309.

Anonymous:1964:BMd

[660] Anonymous. Back matter. Isis, 55(4):??, December 1964. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228336.

Anonymous:1965:FMa

[661] Anonymous. Front matter. Isis, 56(1):1–3, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228454.

Brock:1965:ADM

[662] W. H. Brock and D. M. Knight. The atomic debates: “Memorable and interesting evenings in the life of the Chemical Society”. Isis, 56(1):5– 25, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228455.

Weisheipl:1965:POQ

[663] James A. Weisheipl. The principle Omne quod movetur ab alio movetur in Medieval physics. (Latin) [… It is necessary for any thing that is to be moved to be moved by someone…]. Isis, 56(1):26–45, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228456.

Goodfield:1965:HWT

[664] June Goodfield and Stephen Toulmin. How was the tunnel of Eupalinus aligned? Isis, 56(1):46–55, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228457.

Aiton:1965:GTT

[665] E. J. Aiton and Harold L. Burstyn. Galileo and the theory of the tides. Isis, 56(1):56–63, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228458.

Kargon:1965:WPM

[666] Robert Kargon. William Petty’s mechanical philosophy. Isis, 56(1):63– 66, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228459.

Smith:1965:SOF

[667] Cyril Stanley Smith. A speculation on the origin of Fahrenheit’s temperature scale. Isis, 56(1):66–69, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228460.

Gingerich:1965:LVA

[668] Owen Gingerich. Lunar visibilities in Ancient Babylon. Isis, 56(1):69, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228461.

Hahn:1965:BAB

[669] Roger Hahn and Roger Boscovich. The Boscovich Archives at Berkeley. Isis, 56(1):70–78, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228462.

Anonymous:1965:Na

[670] Anonymous. News. Isis, 56(1):79–82, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228463.

Klein:1965:BRQ

[671] Martin J. Klein. Book review: Quantum comments: Schrödinger — Planck — Einstein — Lorentz. Briefe zur Wellenmechanik, edited by Karl Przibram. Isis, 56(1):83–84, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228464.

Williams:1965:BRB

[672] L. Pearce Williams. Book review: Scientific Change. Historical Studies in the Intellectual, Social and Technical Conditions for Scientific Discovery and Technical Invention, from Antiquity to the Present by Alistair C. Crombie. Isis, 56(1):84–86, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228465.

Ravetz:1965:BRB

[673] Jerome R. Ravetz. Book review: The Nature of the Natural Sciences by Leonard K. Nash. Isis, 56(1):86–87, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228466.

Feyerabend:1965:BRB

[674] Paul K. Feyerabend. Book review: Conjectures and Refutations. The Growth of Scientific Knowledge by Karl R. Popper. Isis, 56(1):88, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228467.

Theodorides:1965:BRB

[675] Jean Théodoridès. Book review: The Evolution of Biology by M. J. Sirks; Conway Zirkle. Isis, 56(1):88–90, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228468.

Harris:1965:BRB

[676] J. R. Harris. Book review: Studies in Ancient Technology. Vol. 7 by R. J. Forbes. Isis, 56(1):90–92, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228469.

Furley:1965:BRB

[677] D. J. Furley. Book review: A History of Greek Philosophy. Volume I: The Earlier Presocratics and the Pythagoreans by W. K. C. Guthrie. Isis, 56(1):92–93, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228470.

Saunders:1965:BRB

[678] J. B. De C. M. Saunders. Book review: The Medical Background of Anglo-Saxon England. A Study in History, Psychology and Folklore by Wilfred Bonser. Isis, 56(1):93–95, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228471.

Clagett:1965:BRB

[679] Marshall Clagett. Book review: A Catalogue of Incipits of Mediaeval Scientific Writings in Latin by Lynn Thorndike; Pearl Kibre. Isis, 56(1):

96–99, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228472.

Grant:1965:BRBa

[680] Edward Grant. Book review: The Discarded Image. An Introduction to Medieval and Renaissance Literature by C. S. Lewis. Isis, 56(1):99– 100, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228473.

Marsak:1965:BRB

[681] Leonard M. Marsak. Book review: Reason and Authority in the Eighteenth Century by Gerald R. Cragg. Isis, 56(1):100–101, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228474.

Hahn:1965:BRB

[682] Roger Hahn. Book review: Fourcroy, Chemist and Revolutionary, 1755– 1809 by William A. Smeaton. Isis, 56(1):101–103, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228475.

Jackson:1965:BRB

[683] Stanley W. Jackson. Book review: An Inquiry concerning the Indications of Insanity with Suggestions for the Better Protection and Care of the Insane by John Conolly; Description of the Retreat. An Institution near York for Insane Persons of the Society of Friends by Samuel Tuke. Isis, 56(1):103–105, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228476.

Stauffer:1965:BRBa

[684] Robert C. Stauffer. Book review: On the Origin of Species by Charles Darwin. Isis, 56(1):105–107, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228477.

DeVaucouleurs:1965:BRB

[685] G. De Vaucouleurs. Book review: Max Wolf: der Bahnbrecher der Himmelsphotographie, 1863–1932 by Hans Christian Freiesleben. Isis, 56(1): 107–108, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228478.

Vorzimmer:1965:BRB

[686] Peter J. Vorzimmer. Book review: Darwiniana: Essays and Reviews Pertaining to Darwinism by Asa Gray; A. Hunter Dupree. Isis, 56(1): 108–110, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228479.

Canguilhem:1965:BRBa

[687] G. Canguilhem. Book review: Das Leben des Biologen Johannes Muller, 1801–1858 by Gottfried Koller. Isis, 56(1):110, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228480.

Barber:1965:BRB

[688] Bernard Barber. Book review: America’s New Policy Makers: The Scientists’ Rise to Power by Donald W. Cox; Cambridge U. S. A.: Hub of a New World by Christopher Rand; The Management of Scientists by Karl Hill. Isis, 56(1):110–111, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228481.

Anderson:1965:BRB

[689] Oscar E. Anderson, Jr. Book review: Britain and Atomic Energy, 1939– 1945 by Margaret Gowing. Isis, 56(1):111–113, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228482.

Hewlett:1965:BRB

[690] Richard G. Hewlett. Book review: The Dispersion of Nuclear Weapons. Strategy and Politics by R. N. Rosecrance. Isis, 56(1):113–114, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228483.

Anonymous:1965:BMa

[691] Anonymous. Back matter. Isis, 56(1):115–116, Spring 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228484.

Anonymous:1965:FMb

[692] Anonymous. Front matter. Isis, 56(2):117–119, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227908.

Vlastos:1965:MPE

[693] Gregory Vlastos. Minimal parts in epicurean atomism. Isis, 56(2):121– 147, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227909.

Vorzimmer:1965:DEI

[694] Peter Vorzimmer. Darwin’s ecology and its influence upon his theory. Isis, 56(2):148–155, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227910.

Pfeifer:1965:GAN

[695] Edward J. Pfeifer. The Genesis of American neo-Lamarckism. Isis, 56 (2):156–167, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227911.

Vinogradov:1965:CKU

[696] S. N. Vinogradov. Chemistry at Kazan University in the Nineteenth Century: A case history of intellectual lineage. Isis, 56(2):168–173, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227912.

Dales:1965:BET

[697] Richard C. Dales. Anonymi De elementis: From a Twelfth-Century collection of scientific works in British Museum MS Cotton Galba E. IV. Isis, 56(2):174–189, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227913.

Khatchadourian:1965:AKE

[698] Haig Khatchadourian, Nicholas Rescher, and Ya’qub ibn Ishaq al Kindi. Al-Kindi’s epistle on the concentric structure of the Universe. Isis, 56 (2):190–195, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227914.

Phillips:1965:KBL

[699] J. P. Phillips. Kepler’s Echinus. (Latin) [Hedgehog]. Isis, 56(2):196– 200, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227915.

Sivin:1965:COW

[700] N. Sivin. On “China’s Opposition to Western Science during Late Ming and Early Ch’ing”. Isis, 56(2):201–205, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227916. See [385].

Guerlac:1965:WSL

[701] Henry Guerlac. The word spectrum: A lexicographic note with a query. Isis, 56(2):206–207, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227917.

Thorndike:1965:ILT

[702] Lynn Thorndike. Iohannes de Luna Theutonicus, about A.D. 1305. Isis, 56(2):207–208, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227918.

Suter:1965:MSC

[703] Rufus Suter. Manzoni and Seventeenth-Century science. Isis, 56(2):208– 209, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227919.

Smith:1965:FTS

[704] Cyril Stanley Smith. Fahrenheit’s temperature scale: Postcript to a speculation. Isis, 56(2):209–210, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227920.

Snelders:1965:RQN

[705] H. A. M. Snelders. Reply to query no. 155. Isis, 56(2):210–211, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227921.

Anonymous:1965:Nb

[706] Anonymous. News. Isis, 56(2):212–213, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227922.

Nielsen:1965:BRO

[707] J. Rud Nielsen. Book review: The occasional Bohr: The unity of knowledge: Niels Bohr: Essays 1958/1962 on Atomic Physics and Human Knowledge. Isis, 56(2):214–216, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227923.

Drennan:1965:BRB

[708] Ollin J. Drennan. Book review: A Brief History of Science by A. Rupert Hall; Marie Boas Hall. Isis, 56(2):216–217, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227924.

Africa:1965:BRB

[709] Thomas W. Africa. Book review: Four Hundred Years of the Copernican Heritage by Eugeniusz Rybka; Edward Jozef Czerwinski. Isis, 56(2):217– 218, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227925.

Scriven:1965:BRB

[710] Michael Scriven. Book review: The Philosophy of Science. Part Two: A Study of the Division and Nature of Various Groups of Sciences by P. Henry van Laer. Isis, 56(2):218–220, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227926.

Durling:1965:BRB

[711] Richard J. Durling. Book review: A History of the Royal College of Physicians of London. Volume 1 by George Clark. Isis, 56(2):220–221, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227927.

Feibleman:1965:BRB

[712] James K. Feibleman. Book review: The Scientific Age. The Impact of Science on Society by L. V. Berkner; Science as a Cultural Force by Harry Woolf. Isis, 56(2):221–222, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227928.

Boyer:1965:BRB

[713] Carl B. Boyer. Book review: The Treasury of Mathematics by Henrietta O. Midonick. Isis, 56(2):222–223, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227929.

Brown:1965:BRB

[714] Lloyd A. Brown. Book review: History of Cartography by Leo Bagrow; R. A. Skelton. Isis, 56(2):224, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227930.

Page:1965:BRB

[715] Leroy E. Page. Book review: The History of the Study of Landforms, or the Development of Geomorphology. Volume 1. Geomorphology before Davis by Richard J. Chorley; Antony J. Dunn; Robert P. Beckinsale. Isis, 56(2):225–226, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227931.

Skelton:1965:BRB

[716] R. A. Skelton. Book review: Kunstler des Kartenbildes. Biographien und Portrats by Hans Harms. Isis, 56(2):226, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227932.

Deacon:1965:BRB

[717] G. E. R. Deacon. Book review: Marine Cartography in Britain: A History of the Sea Chart to 1855 by A. H. W. Robinson. Isis, 56(2):226–228, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227933.

Holmes:1965:BRB

[718] Frederic L. Holmes. Book review: A Short History of Biology by Isaac Asimov. Isis, 56(2):228–229, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227934.

Kertesz:1965:BRB

[719] Z. I. Kertesz. Book review: A magyar talajtani kutalas torlenete 1944ig. (The History of Hungarian Soil Research up to 1944.) by Robert Ballenegger; Istvan Finaly. Isis, 56(2):229–230, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227935.

Krech:1965:BRB

[720] David Krech. Book review: Visual Perception by William Dember; Thinking: From Association to Gestalt by Jean Matter Mandler; George Mandler; Mathematics and Psychology by George A. Miller. Isis, 56 (2):230–232, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227936.

Brozek:1965:BRB

[721] J. Brožek. Book review: A Short History of British Psychology. 1840– 1940 by L. S. Hearnshaw. Isis, 56(2):232–233, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227937.

Benedict:1965:BRB

[722] Burton Benedict. Book review: Human Nature and the Study of Society: The Papers of Robert Redfield. Volume 1 by Robert Redfield; Margaret Park Redfield. Isis, 56(2):234, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227938.

Lefanu:1965:BRB

[723] W. R. Lefanu. Book review: Doctors Monro. A Medical Saga by Rex E. Wright-St. Clair. Isis, 56(2):234–235, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227939.

Bedini:1965:BRB

[724] Silvio A. Bedini. Book review: Dictionnaire archéologique des techniques. Isis, 56(2):235–236, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227940.

Evans:1965:BRB

[725] Clifford Evans. Book review: Ancient Ruins and Archaeology by L. Sprague De Camp; Catherine C. De Camp. Isis, 56(2):236, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227941.

Palisca:1965:BRB

[726] Claude V. Palisca. Book review: Musical Thought in Ancient Greece by Edward A. Lippman. Isis, 56(2):237, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227942.

vonFritz:1965:BRB

[727] Kurt von Fritz. Book review: Pitagorici. Testimonianze e Frammenti by Maria Timpanaro-Cardini. Isis, 56(2):238–239, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227943.

Grant:1965:BRBb

[728] Edward Grant. Book review: Manuscrits scientifiques médiévaux de l’université de Salamanque et de ses “colegios mayores” by Guy Beaujouan. Isis, 56(2):239–240, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227944.

Cohen:1965:BRB

[729] I. Bernard Cohen. Book review: The Scientific Renaissance, 1450–1630 by Marie Boas. Isis, 56(2):240–242, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227945.

Hankins:1965:BRB

[730] Thomas L. Hankins. Book review: Jean d’Alembert, 1717–83 by Ronald Grimsley. Isis, 56(2):242–243, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227946.

Ihde:1965:BRB

[731] Aaron J. Ihde. Book review: Iconographia Boerhaavii (Volume 4 of Analecta Boerhaaviana) by G. A. Lindeboom. Isis, 56(2):243, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227947.

Multhauf:1965:BRBa

[732] Robert P. Multhauf. Book review: Essayes of Natural Experiments. Made in the Academie del Cimento, Under the Protection of the Most Serene Prince Leopold of Tuscany by Richard Waller; Experiments and Considerations Touching Colours by Robert Boyle. Isis, 56(2):243–244, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227948.

Canguilhem:1965:BRBb

[733] G. Canguilhem. Book review: La Vie et l’oeuvre d’Etienne Geoffroy Saint-Hilaire by Theophile Cahn. Isis, 56(2):244–246, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227949.

Sharpe:1965:BRBa

[734] William D. Sharpe. Book review: Malaria in the Interior Valley of North America by Daniel Drake; Norman D. Levine. Isis, 56(2):246, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227950.

Rosenberg:1965:BRB

[735] Charles E. Rosenberg. Book review: Science in Nineteenth-Century America. A Documentary History by Nathan Reingold. Isis, 56(2):246– 247, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227951.

Shryock:1965:BRB

[736] Richard H. Shryock. Book review: The Cholera Years: The United States in 1832, 1849, and 1866 by Charles E. Rosenberg. Isis, 56(2):248–249, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227952.

Cannon:1965:BRBa

[737] Walter Cannon. Book review: Of Men and Galaxies by Fred Hoyle. Isis, 56(2):249, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227953.

Anonymous:1965:BMb

[738] Anonymous. Back matter. Isis, 56(2):250–253, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227954.

Anonymous:1965:FMc

[739] Anonymous. Front matter. Isis, 56(3):255–257, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228101.

MacLeod:1965:ERO

[740] Roy M. MacLeod. Evolutionism and Richard Owen, 1830–1868: An episode in Darwin’s century. Isis, 56(3):259–280, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228102.

Hankins:1965:ECA

[741] Thomas L. Hankins. Eighteenth-Century attempts to resolve the Vis viva controversy. Isis, 56(3):281–297, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228103.

Jayawardene:1965:RBE

[742] S. A. Jayawardene. Rafael Bombelli, engineer-architect: Some unpublished documents of the Apostolic Camera. Isis, 56(3):298–306, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228104.

Reti:1965:PGS

[743] Ladislao Reti. Parting of gold and silver with nitric acid in a page of the Codex Atlanticus of Leonardo da Vinci. Isis, 56(3):307–319, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228105.

Middleton:1965:PHM

[744] W. E. Knowles Middleton. P. H. Maille, a forgotten pioneer in meteorology. Isis, 56(3):320–326, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228106.

Grant:1965:PNO

[745] Edward Grant and Nicole Oresme. Part I of Nicole Oresme’s algorismus proportionum. Isis, 56(3):327–341, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228107.

Susskind:1965:HTS

[746] Charles Süsskind. Hertz and the technological significance of electromagnetic waves. Isis, 56(3):342–345, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228108.

May:1965:OFC

[747] Kenneth O. May. The origin of the four-color conjecture. Isis, 56(3):346– 348, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228109.

Chambers:1965:DMK

[748] David Wade Chambers. Did the Maya know the metonic cycle? Isis, 56 (3):348–351, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228110.

Lindberg:1965:GEF

[749] David C. Lindberg. Galileo’s experiments on falling bodies. Isis, 56 (3):352–354, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228111.

Barr:1965:NSW

[750] E. Scott Barr. Nature’s “scientific worthies”. Isis, 56(3):354–356, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228112.

Snelders:1965:LTE

[751] H. A. M. Snelders and Albert V. Carozzi. Lamarck’s theory of the Earth: Comments. Isis, 56(3):356–357, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228113. See [598].

Zirkle:1965:N

[752] Conway Zirkle. News. Isis, 56(3):358–361, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228114.

Leibowitz:1965:BRB

[753] Joshua O. Leibowitz. Book review: Andreas Vesalius of Brussels, 1514– 1564. Isis, 56(3):362–365, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228115.

Pease:1965:BRB

[754] Otis Pease. Book review: Man and Nature in America by Arthur A. Ekirch. Isis, 56(3):365–367, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228116.

Herivel:1965:BRB

[755] J. W. Herivel. Book review: From Galileo to Newton, 1630–1720 by A. Rupert Hall. Isis, 56(3):367–368, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228117.

Mayerhofer:1965:BRB

[756] Josef Mayerhöfer. Book review: Asimov’s Biographical Encyclopedia of Science and Technology by Isaac Asimov. Isis, 56(3):368–369, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228118.

Fleming:1965:BRB

[757] Donald Fleming. Book review: The Scientific Intellectual. The Psychological and Sociological Origins of Modern Science by Lewis S. Feuer. Isis, 56(3):369–370, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228119.

Laurie:1965:BRB

[758] P. S. Laurie. Book review: The Planet Saturn. A History of Observation, Theory and Discovery by A. F. O’D. Alexander. Isis, 56(3):370–371, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228120.

Leicester:1965:BRB

[759] Henry M. Leicester. Book review: Classics in the Theory of Chemical Combination by O. Theodor Benfey. Isis, 56(3):371, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228121.

Multhauf:1965:BRBb

[760] Lettie Multhauf. Book review: Bibliographie générale de l’astronomie jusqu’en 1880 by J. C. Houzeau; A. Lancaster. Isis, 56(3):371–372, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228122.

Farber:1965:BRBa

[761] Eduard Farber. Book review: The Development of Modern Chemistry by Aaron J. Ihde. Isis, 56(3):372–373, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228123.

Mills:1965:BRB

[762] Deborah J. Mills. Book review: Exploration of the Universe. The Story of Astronomy by H. C. King. Isis, 56(3):373, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228124.

Hammond:1965:BRB

[763] Lincoln Davis Hammond. Book review: My Voyage Around the World. The Chronicles of a 16th-Century Florentine Merchant by Francesco Carletti; Herbert Weinstock; Byron’s Journal of His Circumnavigation, 1764–1766 by John Byron; Robert E. Gallagher. Isis, 56(3):373–375, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228125.

Freeman:1965:BRB

[764] John F. Freeman. Book review: The Lewis and Clark Expedition by Meriwether Lewis; The Natural History of the Lewis and Clark Expedition by Raymond Darwin Burroughs; Letters of the Lewis and Clark Expedition with Related Documents, 1783–1854 by Donald Jackson. Isis, 56(3):375–377, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228126.

Mayr:1965:BRB

[765] Ernst Mayr. Book review: Die Ornithologen Mittel-Europas by Ludwig Gebhardt. Isis, 56(3):377, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228127.

Mani:1965:BRB

[766] Nikolaus Mani. Book review: A Short History of the Gout and the Rheumatic Diseases by W. S. C. Copeman. Isis, 56(3):377–379, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228128.

Blake:1965:BRB

[767] John B. Blake. Book review: Two Centuries of Medicine: A History of the School of Medicine, University of Pennsylvania by George W. Corner. Isis, 56(3):379, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228129.

McDaniel:1965:BRB

[768] W. B. McDaniel. Book review: Medicine and Health in New Jersey: A History by David L. Cowen. Isis, 56(3):380, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228130.

Gruman:1965:BRB

[769] Gerald J. Gruman. Book review: A History of the Care and Study of the Mentally Retarded by Leo Kanner. Isis, 56(3):380–381, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228131.

Sharpe:1965:BRBb

[770] William D. Sharpe. Book review: Les Commentaires de Martin de Saint-Gille sur les Amphorismes Ypocras by Germaine Lafeuille. Isis, 56(3): 381–382, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228132.

Stannard:1965:BRB

[771] Jerry Stannard. Book review: Pliny’s Natural History. A Selection from Philemon Holland’s Translation by J. Newsome. Isis, 56(3):382–383, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228133.

Eisele:1965:BRB

[772] Carolyn Eisele. Book review: Ein byzantinisches Rechenbuch des 15. Jahrhunderts by Herbert Hunger; Kurt Vogel. Isis, 56(3):383–384, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228134.

Ferguson:1965:BRB

[773] Eugene S. Ferguson. Book review: John Rennie (1761–1821). The Life and Work of a Great Engineer by Cyril T. G. Boucher. Isis, 56(3):384– 386, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228135.

Almgren:1965:BRB

[774] Beverly S. Almgren. Book review: M. V. Lomonosov, Astronom i Astrofizik by P. G. Kulikovskii. Isis, 56(3):386–387, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228136.

Heilbron:1965:BRB

[775] John L. Heilbron. Book review: Ebenezer Kinnersley, Franklin’s Friend by J. A. Leo Lemay. Isis, 56(3):387–388, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228137.

Beer:1965:BRB

[776] John J. Beer. Book review: The Rise of the American Chemistry Profession, 1850–1900 by Edward H. Beardsley. Isis, 56(3):388–389, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228138.

Follesdal:1965:BRB

[777] Dagfinn Føllesdal. Book review: Bolzano’s Logic by Jan Berg. Isis, 56 (3):390–391, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228139.

Cannon:1965:BRBb

[778] Walter F. Cannon. Book review: Charles Darwin and His World by Julian Huxley; H. B. D. Kettlewell. Isis, 56(3):391–392, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228140.

Spencer:1965:BRB

[779] J. Brookes Spencer. Book review: Faraday, Maxwell, and Kelvin by D. K. C. MacDonald. Isis, 56(3):392–393, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228141.

Badash:1965:BRB

[780] Lawrence Badash. Book review: The Discovery of Radioactivity and Transmutation by Alfred Romer. Isis, 56(3):393–394, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228142.

Farber:1965:BRBb

[781] Eduard Farber. Book review: Vie de Berthelot by Léon Velluz. Isis, 56 (3):394–395, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228143.

Pendergrass:1965:BRB

[782] Eugene P. Pendergrass. Book review: The Life and Times of Emil H. Grubbe by Paul C. Hodges. Isis, 56(3):395, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228144.

Clements:1965:BRBa

[783] Raymond D. Clements. Book review: Bibliography of the History of Medicine of the United States and Canada (1939–1960) by Genevieve Miller. Isis, 56(3):395–397, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228145.

Durant:1965:BRB

[784] F. C. Durant, III. Book review: Astronautics and Aeronautics, 1963. Chronology on Science, Technology, and Policy by National Aeronautics and Space Administration Historical Staff. Isis, 56(3):397, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228146.

Anonymous:1965:BMc

[785] Anonymous. Back matter. Isis, 56(3):398–400, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228147.

Hooker:1965:OVC

[786] Marjorie Hooker. The origin of the volcanological concept nuee ardente. Isis, 56(4):401–407, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228767.

Howe:1965:RVH

[787] Herbert M. Howe. A root of van Helmont’s tree. Isis, 56(4):408–419, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228768.

Stannard:1965:PRB

[788] Jerry Stannard. Pliny and Roman botany. Isis, 56(4):420–425, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228769.

Rescher:1965:AKE

[789] Nicholas Rescher, Haig Khatchadourian, and Ya’qub ibn Ishaq al Kindi. Al-Kindi’s epistle on the finitude of the Universe. Isis, 56(4):426–433, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228770.

Rosen:1965:EIE

[790] Edward Rosen. Eloge: Israel Edward Drabkin (7 February 1905–27 March 1965). Isis, 56(4):434–437, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228771.

Hartner:1965:NP

[791] Willy Hartner. Notes on picatrix. Isis, 56(4):438–451, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228772.

Suter:1965:NIA

[792] Rufus Suter. A note on the identity of Ascanio Piccolomini, Galileo’s host at Siena. Isis, 56(4):452, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228773.

Hindle:1965:PTI

[793] Brooke Hindle and Henry Guerlac. Proceedings of the tenth international Congress of the history of science. Isis, 56(4):453–454, Winter 1965.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228774.

Aiton:1965:N

[794] E. J. Aiton. News. Isis, 56(4):455–457, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228775.

White:1965:BRB

[795] Lynn White, Jr. Book review: Saturn and Melancholy, by Raymond Klibansky; Erwin Panofsky; Fritz Saxl. Studies in the History of Natural Philosophy, Religion and Art. Isis, 56(4):458–459, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228776.

Scott:1965:BRB

[796] J. F. Scott. Book review: Bibliographia Cartesiana. A Critical Guide to the Descartes Literature, 1800–1960 by Gregor Sebba. Isis, 56(4): 459, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228777.

Loemker:1965:BRB

[797] L. E. Loemker. Book review: Philosophy, Science, and Sense Perception: Historical and Critical Studies by Maurice Mandelbaum. Isis, 56(4):459– 461, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228778.

Goodfield:1965:BRB

[798] June Goodfield. Book review: Heat and Life. The Development of the Theory of Animal Heat by Everett Mendelsohn. Isis, 56(4):461–465, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228779.

Straus:1965:BRB

[799] William L. Straus, Jr. Book review: British Anatomy, 1525–1800: A Bibliography by K. F. Russell. Isis, 56(4):465, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228780.

Hall:1965:BRB

[800] Thomas S. Hall. Book review: Foundations of Experimental Embryology by Benjamin H. Willier; Jane M. Oppenheimer. Isis, 56(4):465–466, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228781.

Clements:1965:BRBb

[801] Raymond D. Clements. Book review: A Hundred Years of Psychology, 1833–1933. With an Additional Part, 1933–1963 by J. C. Flugel; Donald J. West. Isis, 56(4):466–467, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228782.

Leake:1965:BRB

[802] Chauncey D. Leake. Book review: Readings in Pharmacology by B. Holmstedt; G. Liljestrand. Isis, 56(4):467–469, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228783.

Wertime:1965:BRB

[803] Theodore A. Wertime. Book review: Tin in Social and Economic History by Ernest S. Hedges. Isis, 56(4):469–470, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228784.

Battison:1965:BRB

[804] Edwin A. Battison. Book review: A Short History of Machine Tools by L. T. C. Rolt. Isis, 56(4):470–471, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228785.

Mau:1965:BRB

[805] Jurgen H. Mau. Book review: Institutio logica by Galen; John S. Kiefer. Isis, 56(4):471–473, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228786.

Strang:1965:BRB

[806] Colin Strang. Book review: Anaxagoras and the Birth of Physics by Daniel E. Gershenson; Daniel A. Greenberg. Isis, 56(4):473–474, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228787.

Schullian:1965:BRB

[807] Dorothy M. Schullian. Book review: Opere di Ippocrate, a cura di Mario Vegetti by Hippocrates. Isis, 56(4):474, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228788.

Grant:1965:BRBc

[808] Edward Grant. Book review: Commentary on Aristotle’s Physics by St. Thomas Aquinas; Richard J. Blackwell; Richard J. Spath; W. Edmund Thirlkel. Isis, 56(4):474–475, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228789.

Daumas:1965:BRB

[809] Maurice Daumas. Book review: Early American Scientific Instruments and Their Makers by Silvio A. Bedini. Isis, 56(4):476–478, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228790.

Hellman:1965:BRB

[810] C. Doris Hellman. Book review: Gesammelte Werke. Band 8: Mysterium cosmographicum. Editio altera cum notis; De cometis; Hyperaspistes by Johannes Kepler. Isis, 56(4):478–479, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228791.

Gingerich:1965:BRB

[811] Owen Gingerich. Book review: Kepler’s Dream by John Lear; Patricia Frueh Kirkwood. Isis, 56(4):479–480, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228792.

Stauffer:1965:BRBb

[812] Robert C. Stauffer. Book review: The Works of Charles Darwin: An Annotated Bibliographical Handlist by R. B. Freeman. Isis, 56(4):480– 481, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228793.

Loewenberg:1965:BRB

[813] Bert James Loewenberg. Book review: Biologist Philosopher. A Study of the Life and Writings of Alfred Russel Wallace by Wilma George. Isis, 56(4):481–482, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228794.

Thiele:1965:BRB

[814] Joachim Thiele. Book review: Ernst Mach, Wegbereiter der modernen Physik. Mit ausgewahlten Kapiteln aus seinem Werk by K. D. Heller. Isis, 56(4):482, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228795.

Bishop:1965:BRB

[815] P. W. Bishop. Book review: Iron and Steel in Nineteenth-Century America. An Economic Enquiry by Peter Temin. Isis, 56(4):483–484, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228796.

Finn:1965:BRB

[816] Bernard S. Finn. Book review: Michael Faraday. A Biography by L. Pearce Williams. Isis, 56(4):484–485, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228797.

Clarke:1965:BRB

[817] E. J. S. Clarke. Book review: Tizard by Ronald W. Clark. Isis, 56 (4):485–486, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228798.

Brush:1965:BRB

[818] Stephen G. Brush. Book review: The Use of Citation Data in Writing the History of Science by Eugene Garfield; Irving H. Sher; Richard J. Torpie. Isis, 56(4):487, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228799.

Anonymous:1965:FMd

[819] Anonymous. Front matter. Isis, 56(4):488–489, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228766.

Multhauf:1965:NCB

[820] Robert P. Multhauf, Bernard S. Finn, Aleida V. H. Cattell, Gertrude B. Calvert, and Diana D. Menkes. Ninetieth critical bibliography of the history of science and its cultural influences (to 1 January 1965). Isis, 56(4):490–657, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228800.

Anonymous:1965:VI

[821] Anonymous. Volume information. Isis, 56(4):658–665, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228765.

Anonymous:1965:BMd

[822] Anonymous. Back matter. Isis, 56(4):666–669, Winter 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228801.

Anonymous:1966:FMa

[823] Anonymous. Front matter. Isis, 57(1):1–3, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228686.

Webster:1966:RPS

[824] Charles Webster. The recognition of plant sensitivity by English botanists in the Seventeeth Century. Isis, 57(1):5–23, Spring 1966. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228687.

Pav:1966:GSP

[825] Peter Anton Pav. Gassendi’s statement of the principle of inertia. Isis, 57(1):24–34, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228688.

Thackray:1966:ODC

[826] Arnold W. Thackray. The origin of Dalton’s chemical atomic theory: Daltonian doubts resolved. Isis, 57(1):35–55, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228689. See [208, 494].

Scriba:1966:JWT

[827] Christoph J. Scriba. John Wallis’ treatise of angular sections and Thabit ibn Qurra’s generalization of the Pythagorean Theorem. Isis, 57(1):56– 66, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228690.

Hesse:1966:HPA

[828] Mary B. Hesse. Hooke’s philosophical algebra. Isis, 57(1):67–83, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228691.

Struik:1966:EJA

[829] Dirk J. Struik. Eloge: Johan Adriaan Vollgraff (1877–1965). Isis, 57(1): 84, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228692.

Clagett:1966:ELT

[830] Marshall Clagett. Eloge: Lynn Thorndike (1882–1965). Isis, 57(1):85– 89, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228693.

Thorndike:1966:DWR

[831] Lynn Thorndike. I. a daily weather record continued from 1 September 1400 to 25 June 1401. Isis, 57(1):90–99, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228694.

Frederick:1966:ICA

[832] Ralph H. Frederick, Helmut E. Landsberg, and Walter Lenke. II. A climatological analysis of the Basel weather manuscript. Isis, 57(1):99– 101, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228695.

Westfall:1966:NOP

[833] Richard S. Westfall. Newton’s optics: The present state of research. Isis, 57(1):102–107, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228696.

Dolby:1966:NDC

[834] R. G. A. Dolby. A note on Dijksterhuis’ criticism of Newton’s axiomatization of mechanics. Isis, 57(1):108–115, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228697.

Goldstein:1966:NMC

[835] Bernard R. Goldstein. A note on the metonic cycle. Isis, 57(1):115– 116, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228698.

Mayer:1966:PMS

[836] Joseph Mayer. “Parliament and the metric system” — comments. Isis, 57(1):117–119, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228699.

See [393, 837, 986].

Semmel:1966:PMS

[837] Bernard Semmel. “Parliament and the metric system” — a reply. Isis, 57(1):119–120, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228700.

See [393, 836, 986].

Easton:1966:DRR

[838] Joy B. Easton. On the date of Robert Recorde’s birth. Isis, 57(1): 121, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228701.

Anonymous:1966:Na

[839] Anonymous. News. Isis, 57(1):122–124, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228702.

Debus:1966:BRBa

[840] Allen G. Debus. Book review: Robert Boyle on Natural Philosophy, by Marie Boas Hall. An Essay with Selections from His Writings. Isis, 57 (1):125–126, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228703.

Simpson:1966:BRB

[841] George Gaylord Simpson. Book review: The Act of Creation by Arthur Koestler. Isis, 57(1):126–127, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228704.

Watanabe:1966:BRB

[842] Satosi Watanabe. Book review: The Discovery of Time by Stephen Toulmin; June Goodfield. Isis, 57(1):127–129, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228705.

Vogel:1966:BRB

[843] Kurt Vogel. Book review: Principles of Hindu Reckoning by Kushyar ibn Labban; Martin Levey; Marvin Petruck. Isis, 57(1):129–131, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228706.

Gillis:1966:BRB

[844] Jean Gillis. Book review: A History of the Concept of Valency to 1930 by W. G. Palmer. Isis, 57(1):131–132, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228707.

Rowe:1966:BRB

[845] John H. Rowe. Book review: Readings in Early Anthropology by James Sydney Slotkin. Isis, 57(1):132–133, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228708.

Cowen:1966:BRB

[846] David L. Cowen. Book review: David Hosack. Citizen of New York by Christine C. Robbins. Isis, 57(1):133–134, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228709.

OMalley:1966:BRB

[847] C. D. O’Malley. Book review: Surgery and Ambroise Pare by J. F. Malgaigne; Wallace B. Hamby; Grundlicher Bericht vom Heissen und kalten Brand by Wilhelm Fabry von Hilden; Prefazione alla “Fabrica” e lettera a G. Oporino (Testo e versione) by Andreas Vesalius. Isis, 57(1):134–135, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228710.

Multhauf:1966:BRB

[848] Robert P. Multhauf. Book review: The Conquest of the Material World by John U. Nef; The Rise of Industrial Society in England, 1815–1885 by S. G. Checkland. Isis, 57(1):136, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228711.

Pingree:1966:BRB

[849] David Pingree. Book review: Egyptian Astronomical Texts. Vol. 2: The Ramesside Star Clocks by O. Neugebauer; R. Parker. Isis, 57(1):136– 137, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228712.

Crowe:1966:BRB

[850] Michael J. Crowe. Book review: The Birth of Mathematics in the Age of Plato by Francois Lasserre. Isis, 57(1):137–138, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228713.

Durling:1966:BRBa

[851] Richard J. Durling. Book review: Medical Illustrations in Medieval Manuscripts by Loren MacKinney. Isis, 57(1):139–140, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228714.

Kelly:1966:BRB

[852] Suzanne Kelly. Book review: Kepler’s Conversation with Galileo’s Sidereal Messenger by Edward Rosen. Isis, 57(1):141–142, Spring 1966. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228715.

Stimson:1966:BRB

[853] Dorothy Stimson. Book review: Intellectual Origins of the English Revolution by Christopher Hill. Isis, 57(1):142–143, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228716.

Siegfried:1966:BRB

[854] Robert Siegfried. Book review: Supplement to a Bibliography of the Works of Antoine Laurent Lavoisier 1743–1794 by Dennis I. Duveen. Isis, 57(1):143–144, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228717.

Pursell:1966:BRB

[855] Carroll W. Pursell, Jr. Book review: Brooklyn Bridge: Fact and Symbol by Alan Trachtenberg. Isis, 57(1):144–145, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228718.

Zarechnak:1966:BRB

[856] Galina Zarechnak. Book review: Essays on the History of Physiology in Russia by Khachatur Sedrakovich Koshtoiants; David Boder; Kristan Hanes; Natalie O’Brien; Donald B. Lindsley. Isis, 57(1):145–146, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228719.

Major:1966:BRBa

[857] Ralph H. Major. Book review: Die Wiener medizinische Schule im 19. Jahrhundert by Erna Lesky. Isis, 57(1):146–148, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228720.

Major:1966:BRBb

[858] Ralph H. Major. Book review: Ignaz Philipp Semmelweis und die Wiener medizinische Schule by Erna Lesky. Isis, 57(1):148–149, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228721.

Anonymous:1966:BMa

[859] Anonymous. Back matter. Isis, 57(1):150–152, Spring 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228722.

Anonymous:1966:FMb

[860] Anonymous. Front matter. Isis, 57(2):153–155, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227955.

Cohen:1966:AKC

[861] I. Bernard Cohen and Marshall Clagett. Alexandre Koyre (1892–1964): Commemoration. Isis, 57(2):157–166, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227956.

Burstyn:1966:EER

[862] Harold L. Burstyn. Early explanations of the role of the Earth’s rotation in the circulation of the atmosphere and the ocean. Isis, 57(2):167– 187, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227957.

Woodruff:1966:WCR

[863] A. E. Woodruff. William Crookes and the radiometer. Isis, 57(2):188– 198, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227958.

Bork:1966:FC

[864] Alfred M. Bork. The “FitzGerald” contraction. Isis, 57(2):199–207, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227959.

Roberts:1966:PTI

[865] Victor Roberts. The planetary theory of Ibn al-Shatir: Latitudes of the planets. Isis, 57(2):208–219, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227960.

Blackwell:1966:DLM

[866] Richard J. Blackwell. Descartes’ laws of motion. Isis, 57(2):220–234, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227961.

Lindberg:1966:RBT

[867] David C. Lindberg. Roger Bacon’s theory of the rainbow: Progress or regress? Isis, 57(2):235–248, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227962.

Merton:1966:ONP

[868] Stephen Merton. Old and new physiology in Sir Thomas Browne: Digestion and some other functions. Isis, 57(2):249–259, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227963.

Hahn:1966:ETF

[869] Roger Hahn. Eloges: Three French historians of science. Isis, 57(2):260– 261, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227964.

Rosen:1966:GKT

[870] Edward Rosen. Galileo and Kepler: Their first two contacts. Isis, 57 (2):262–264, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227965.

Heilbron:1966:GMB

[871] John L. Heilbron. G. M. Bose: The prime mover in the invention of the Leyden jar? Isis, 57(2):264–267, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227966.

Badash:1966:BUP

[872] Lawrence Badash. Becquerel’s “unexposed” photographic plates. Isis, 57(2):267–269, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227967.

Drake:1966:FFG

[873] Stillman Drake. Free fall in Galileo’s dialogue. Isis, 57(2):269–271, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227968.

Anonymous:1966:Nb

[874] Anonymous. News. Isis, 57(2):272–273, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227969.

Heilbron:1966:BRB

[875] John L. Heilbron. Book review: Guilielmi Gilberti Colcestrensis, medici regii, De mundo nostro sublunari philosophia nova. Isis, 57(2):274–275, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227970.

Reingold:1966:BRBa

[876] Nathan Reingold. Book review: Science: U.S.A. by William Gilman. Isis, 57(2):275–276, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227971.

Sherwood:1966:BRB

[877] Morgan B. Sherwood. Book review: The French in the Mississippi Valley by John Francis McDermott. Isis, 57(2):276–277, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227972.

Carozzi:1966:BRB

[878] Albert V. Carozzi. Book review: L’esplorazione naturalistica dell ’Appennino by Francesco Rodolico. Isis, 57(2):277–278, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227973.

Zirkle:1966:BRBa

[879] Conway Zirkle. Book review: A History of Genetics by A. H. Sturtevant. Isis, 57(2):278–279, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227974.

Clagett:1966:BRB

[880] Marshall Clagett. Book review: Codex Constantinopolitanus. Palatii Veteris No. 1 by E. M. Bruins. Isis, 57(2):280–281, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227975.

Durling:1966:BRBb

[881] Richard J. Durling. Book review: The Greek and Latin Literary Texts from Greco-Roman Egypt by Roger Ambrose Pack. Isis, 57(2):281, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227976.

Caley:1966:BRB

[882] Earle R. Caley. Book review: De lapidibus by Theophrastus; D. E. Eichholz. Isis, 57(2):281–282, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227977.

Struik:1966:BRB

[883] Dirk J. Struik. Book review: David Rittenhouse by Brooke Hindle. Isis, 57(2):282–284, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227978.

McDaniel:1966:BRB

[884] W. B. McDaniel. Book review: John Shaw Billings, Selected Papers of by Frank Bradway Rogers; John Shaw Billings. Isis, 57(2):284–285, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227979.

Duffy:1966:BRB

[885] John Duffy. Book review: Doctors and the State. The British Medical Profession and Government Action in Public Health, 1870–1912 by Jeanne L. Brand. Isis, 57(2):285–286, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227980.

Egerton:1966:BRB

[886] Frank N. Egerton, III. Book review: The History of Botany in Finland, 1828–1918 by Runar Collander; David Barrett. Isis, 57(2):286– 287, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227981.

Washburn:1966:BRBa

[887] Wilcomb E. Washburn. Book review: Meriwether Lewis: A Biography by Richard Dillon. Isis, 57(2):287–288, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227982.

Farber:1966:BRB

[888] Eduard Farber. Book review: Collected Scientific Papers by Wolfgang Pauli; R. Kronig; V. F. Weisskopf. Isis, 57(2):288–289, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227983.

Reingold:1966:BRBb

[889] Nathan Reingold. Book review: The Politics of American Science: 1939 to the Present by James L. Penick; Carroll W. Pursell; Morgan B. Sherwood; Donald C. Swain. Isis, 57(2):289–290, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227984.

Blake:1966:BRB

[890] William D. Blake. Book review: Homer William Smith, Sc.D. His Scientific and Literary Achievements by Herbert Chasis; William Goldring; Homer William Smith. Isis, 57(2):290–291, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227985.

Anonymous:1966:BMb

[891] Anonymous. Back matter. Isis, 57(2):292–293, Summer 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227986.

Anonymous:1966:FMc

[892] Anonymous. Front matter. Isis, 57(3):295–297, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228361.

Westfall:1966:NDH

[893] Richard S. Westfall. Newton defends his first publication: The Newton– Lucas correspondence. Isis, 57(3):299–314, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228362.

Hodgen:1966:EPV

[894] Margaret T. Hodgen. Ethnology in 1500: Polydore Vergil’s collection of customs. Isis, 57(3):315–324, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228363.

Siegfried:1966:SHD

[895] Robert Siegfried. Sir Humphry Davy on the nature of the diamond. Isis, 57(3):325–335, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228364.

Heilbron:1966:WHG

[896] John L. Heilbron. The work of H. G. J. Moseley. Isis, 57(3):336–364, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228365. See [985].

Kennedy:1966:LMP

[897] E. S. Kennedy. Late medieval planetary theory. Isis, 57(3):365–378, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228366.

Stannard:1966:ELE

[898] Jerry Stannard. Eloge: Ludwig Edelstein (1902–1965). Isis, 57(3):379– 384, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228367.

Celoria:1966:DGC

[899] Francis Celoria. Delta as a geographical concept in Greek literature. Isis, 57(3):385–388, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228368.

Salie:1966:PP

[900] Hans Salie and Ralph E. Oesper. Poggendorff and Poggendorff. Isis, 57 (3):389–392, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228369.

Rosen:1966:KDT

[901] Edward Rosen and Patricia Frueh Kirkwood. Kepler’s dream in translation. Isis, 57(3):392–394, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228370.

Anonymous:1966:Nc

[902] Anonymous. News. Isis, 57(3):395, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228371.

Bullough:1966:BRB

[903] Vern L. Bullough. Book review: The Medical Practitioners in Medieval England: A Biographical Register. Isis, 57(3):396–397, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228372.

Greenaway:1966:BRB

[904] Frank Greenaway. Book review: Selected Readings in the History of Chemistry by Aaron J. Ihde; William F. Kieffer. Isis, 57(3):397–398, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228373.

Zirkle:1966:BRBb

[905] Conway Zirkle. Book review: Botanical Latin: History, Grammar, Syntax, Terminology and Vocabulary by William T. Stearn. Isis, 57(3):398– 399, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228374.

Gershenson:1966:BRB

[906] Daniel E. Gershenson. Book review: Aristotle’s Vision of Nature by Frederick J. E. Woodbridge; John Herman Randall; Charles H. Kahn; Harold Larrabee. Isis, 57(3):399–400, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228375.

Debus:1966:BRBb

[907] Allen G. Debus. Book review: Michael Scot by Lynn Thorndike. Isis, 57(3):400–402, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228376.

Nebbia:1966:BRB

[908] Giorgio Nebbia. Book review: A Pharmaceutical View of Abulcasis al-Zahrawi in Moorish Spain with Special Reference to the “Adhan” by Sami Khalaf Hamarneh; Glenn Sonnedecker. Isis, 57(3):402–403, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228377.

Kargon:1966:BRB

[909] Robert Kargon. Book review: The Background to Newton’s Principia by John Herivel. Isis, 57(3):403–404, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228378.

Diethelm:1966:BRB

[910] Oskar Diethelm. Book review: Mind and Body in Eighteenth-Century Medicine. A Study Based on Jerome Gaub’s De regimine mentis by L. J. Rather. Isis, 57(3):404–405, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228379.

Costa:1966:BRB

[911] Albert B. Costa. Book review: Animal Chemistry or Organic Chemistry in Its Application to Physiology and Pathology by Justus Liebig; William Gregory. Isis, 57(3):405–406, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228380.

Anonymous:1966:BMc

[912] Anonymous. Back matter. Isis, 57(3):407–408, Autumn 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228381.

Anonymous:1966:FMd

[913] Anonymous. Front matter. Isis, 57(4):409–515, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228513.

Simpson:1966:MDE

[914] Thomas K. Simpson. Maxwell and the direct experimental test of his electromagnetic theory. Isis, 57(4):411–432, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228514.

Hesse:1966:HVT

[915] Mary Hesse. Hooke’s vibration theory and the isochrony of springs. Isis, 57(4):433–441, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228515.

Nakayama:1966:CCA

[916] Shigeru Nakayama. Characteristics of Chinese astrology. Isis, 57(4):442– 454, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228516.

Dales:1966:TRG

[917] Richard C. Dales and Robert Grosseteste. The text of Robert Grosseteste’s questio de fluxu et refluxu maris with an English translation.

Isis, 57(4):455–474, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228517.

Saidan:1966:EEA

[918] A. S. Saidan. The earliest extant Arabic arithmetic: Kitab al-Fusul fi al Hisab al-Hindi of Abu al-Hasan, Ahmad ibn Ibrahim al-Uqlidisi. Isis, 57(4):475–490, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228518.

Anonymous:1966:Nd

[919] Anonymous. News. Isis, 57(4):491–493, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228519.

Anonymous:1966:CN

[920] Anonymous. Correction: News. Isis, 57(4):493, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228520.

Chambers:1966:HSS

[921] David Wade Chambers. History of science on the silver screen. Isis, 57(4):494–497, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228521.

Finn:1966:BRB

[922] Bernard S. Finn. Book review: The New Treasury of Science by Harlow Shapley; Samuel Rapport; Helen Wright. Isis, 57(4):497–498, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228522.

Harre:1966:BRB

[923] R. Harre. Book review: Science, Technology, and Human Values by A. Cornelius Benjamin. Isis, 57(4):498–499, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228523.

Kantor:1966:BRB

[924] J. R. Kantor. Book review: Psychoanalytic Pioneers by Franz Alexander; Samuel Eisenstein; Martin Grotjahn. Isis, 57(4):499–500, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228524.

Stannard:1966:BRBa

[925] Jerry Stannard. Book review: The Plants of Virgil’s by Elfriede Abbe. Isis, 57(4):500–501, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228525.

Cohen:1966:BRB

[926] I. Bernard Cohen. Book review: On Motion and On Mechanics by Galileo Galilei; I. E. Drabkin; Le Meccaniche; Stillman Drake. Isis, 57(4):501– 504, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228526.

Washburn:1966:BRBb

[927] Wilcomb E. Washburn. Book review: The Principall Navigations Voiages & Discoveries of the English Nation by Richard Hakluyt; David B. Quinn; Raleigh A. Skelton. Isis, 57(4):504–505, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228527.

Stannard:1966:BRBb

[928] Jerry Stannard. Book review: The Anatomy of Plants: With an Idea of a Philosophical History of Plants and Several Other Lectures Read before the Royal Society by Nehemiah Grew. Isis, 57(4):505–506, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228528.

Ore:1966:BRB

[929] Øystein Øre. Book review: Leonhard Euler und Christian Goldbach Briefwechsel 1729–1764 by A. P. Juskevic; E. Winter. Isis, 57(4):506– 507, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228529.

Clarke:1966:BRB

[930] Edwin Clarke. Book review: Lecture on the Anatomy of the Brain by Nicolaus Steno. Isis, 57(4):507–508, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228530.

Sivin:1966:BRB

[931] N. Sivin. Book review: T’ien-Kung K’ai-Wu. Chinese Technology in the Seventeenth Century by Sung Ying-hsing; E-tu Zen Sun; Shiou-chuan Sun. Isis, 57(4):508–509, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228531.

Bibby:1966:BRB

[932] Cyril Bibby. Book review: South Kensington to Robbins: An Account of English Scientific and Technical Education since 1851 by Michael Argles. Isis, 57(4):510–511, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228532.

Palter:1966:BRB

[933] Robert Palter. Book review: The Mystery of Matter by Louise B. Young. Isis, 57(4):511–513, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228533.

Multhauf:1966:NFC

[934] Robert P. Multhauf, Bernard S. Finn, Aleida Cattell Renwick, and Diana Davis Menkes. Ninety-first critical bibliography of the history of science and its cultural influences (to 1 January 1966). Isis, 57(4):516– 644, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228534.

Anonymous:1966:VI

[935] Anonymous. Volume information. Isis, 57(4):645–650, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228512.

Anonymous:1966:BMd

[936] Anonymous. Back matter. Isis, 57(4):651–654, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228535.

Anonymous:1967:FMa

[937] Anonymous. Front matter. Isis, 58(1):1–3, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228382.

Odom:1967:GRN

[938] Herbert H. Odom. Generalizations on race in Nineteenth-Century physical anthropology. Isis, 58(1):4–18, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228383.

Thoren:1967:EID

[939] Victor E. Thoren. An early instance of deductive discovery: Tycho Brahe’s lunar theory. Isis, 58(1):19–36, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228384.

McCormmach:1967:HPQ

[940] Russell McCormmach. Henri Poincaré and the quantum theory. Isis, 58(1):37–55, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228385.

McVaugh:1967:AVB

[941] Michael McVaugh. Arnald of Villanova and Bradwardine’s Law. Isis, 58(1):56–64, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228386.

Toulmin:1967:ABC

[942] Stephen Toulmin. The astrophysics of Berossos the Chaldean. Isis, 58 (1):65–76, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228387. See

comments [1110, 1329].

Sabra:1967:ABC

[943] A. I. Sabra. The authorship of the Liber de crepusculis, an Eleventh-Century work on atmospheric refraction. Isis, 58(1):77–85, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228388. See correction [1070].

Shankland:1967:RM

[944] R. S. Shankland, Lord Rayleigh, and Albert A. Michelson. Rayleigh and Michelson. Isis, 58(1):86–88, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228389.

Howard:1967:MRC

[945] John N. Howard, Lord Rayleigh, and Michelson. The Michelson–Rayleigh correspondence of the AFCRL Rayleigh Archives. Isis, 58(1):88–89, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228390.

Ospovat:1967:PKK

[946] Alexander M. Ospovat. The place of the Kureze Klassifikation in the work of A. G. Werner. Isis, 58(1):90–95, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228391.

Gorman:1967:FLR

[947] Mel Gorman. Faraday on lightning rods. Isis, 58(1):96–98, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228392.

Pelseneer:1967:GOR

[948] Jean Pelseneer. Galilee ou le recours au protestantisme. Isis, 58(1):99– 102, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228393.

Boring:1967:PLP

[949] Edwin G. Boring. Psychologists’ letters and papers. Isis, 58(1):103– 107, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228394.

Anonymous:1967:Na

[950] Anonymous. News. Isis, 58(1):108–109, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228395.

Hill:1967:BRB

[951] Edward L. Hill. Book review: Philosophical Foundations of Physics. Isis, 58(1):110–111, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228396.

Reingold:1967:BRBa

[952] Nathan Reingold. Book review: The Rise of the Technocrats. A Social History by W. H. G. Armytage. Isis, 58(1):111–113, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228397.

Multhauf:1967:BRB

[953] Robert P. Multhauf. Book review: Chemical, Medical and Pharmaceutical Books Printed before 1800 by John Neu. Isis, 58(1):113, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228398.

Voise:1967:BRB

[954] W. Voise. Book review: The Voices of Time. A Cooperative Survey of Man’s Views of Time as Expressed by the Sciences and by the Humanities by J. T. Fraser. Isis, 58(1):113–115, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228399.

Reingold:1967:BRBb

[955] Nathan Reingold. Book review: The Excitement and Fascination of Science: A Collection of Autobiographical and Philosophical Essays by J. Murray Luck. Isis, 58(1):115–116, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228400.

Agassi:1967:BRB

[956] Joseph Agassi. Book review: Science and Culture. A Study of Cohesive and Disjunctive Forces by Gerald Holton. Isis, 58(1):116–118, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228401.

Beer:1967:BRBa

[957] John J. Beer. Book review: Science in Canada. Selections from the Speeches of E. W. R. Steacie by J. D. Babbitt. Isis, 58(1):118–119, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228402.

Cannon:1967:BRBa

[958] Walter F. Cannon. Book review: The Role of Mathematics in the Rise of Science by Salomon Bochner. Isis, 58(1):119–120, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228403.

Bedini:1967:BRBa

[959] Silvio A. Bedini. Book review: The Mathematical Practitioners of Hanoverian England, 1714–1840 by E. G. R. Taylor. Isis, 58(1):120– 121, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228404.

Kargon:1967:BRB

[960] Robert Kargon. Book review: The World of the Atom by Henry A. Boorse; Lloyd Motz. Isis, 58(1):122–123, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228405.

Hesse:1967:BRB

[961] Mary Hesse. Book review: The Origins of Field Theory by L. Pearce Williams. Isis, 58(1):123–124, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228406.

Warner:1967:BRB

[962] Deborah J. Warner. Book review: Explorer of the Universe: A Biography of George Ellery Hale by Helen Wright. Isis, 58(1):124–125, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228407.

Stern:1967:BRB

[963] Curt Stern. Book review: The Cell in Development and Inheritance by Edmund B. Wilson. Isis, 58(1):125–127, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228408.

Bloom:1967:BRB

[964] Leon Bloom. Book review: Freud’s Neurological Education and Its Influence on Psychoanalytic Theory by Peter Amacher. Isis, 58(1):127–128, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228409.

Youschkevitch:1967:BRB

[965] A. P. Youschkevitch. Book review: Archimedes in the Middle Ages. Volume 1: The Arabo-Latin Tradition by Marshall Clagett. Isis, 58(1): 128–131, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228410.

North:1967:BRBa

[966] J. D. North. Book review: De proportionibus proportionum and Ad pauca respicientes by Nicole Oresme; Edward Grant. Isis, 58(1):131– 132, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228411.

Rosenfeld:1967:BRBa

[967] B. A. Rosenfeld. Book review: The Algebra of Abu Kamil, Kitab fi aljabr wa’ l-mugabala, in a Commentary by Mordecai Finzi by Martin Levey. Isis, 58(1):132–134, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228412.

Schneer:1967:BRB

[968] Cecil J. Schneer. Book review: The Six-Cornered Snowflake by Johannes Kepler; Colin Hardie. Isis, 58(1):134–136, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228413.

Rosenfeld:1967:BRBb

[969] L. Rosenfeld. Book review: Descartes i Uppsala by Rolf Lindborg. Isis, 58(1):136–138, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228414.

Hendrickson:1967:BRB

[970] Walter B. Hendrickson. Book review: Exploration and Empire. The Explorer and the Scientist in the Winning of the American West by William H. Goetzmann. Isis, 58(1):138–139, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228415.

Friis:1967:BRB

[971] Herman R. Friis. Book review: Exploration of Alaska, 1865–1900 by Morgan B. Sherwood. Isis, 58(1):139–140, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228416.

Durant:1967:BRBa

[972] F. C. Durant, III. Book review: The Birth of the Missile by Ernest Klee; Otto Merk; T. Schoeters. Isis, 58(1):140–141, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228417.

Scott:1967:BRE

[973] William T. Scott. Book review: Erwin Schrödinger: My View of the World. Translated by Cecily Hastings from the German Meine Weltansicht (1961). Isis, 58(1):141–143, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228418.

Anonymous:1967:BMa

[974] Anonymous. Back matter. Isis, 58(1):144–146, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228419.

Anonymous:1967:FMb

[975] Anonymous. Front matter. Isis, 58(2):147–149, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228220.

Daniels:1967:PPA

[976] George H. Daniels. The process of professionalization in American science: The emergent period, 1820–1860. Isis, 58(2):150–166, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228221.

Lipman:1967:VRL

[977] Timothy O. Lipman. Vitalism and reductionism in Liebig’s physiological thought. Isis, 58(2):167–185, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228222.

Suchting:1967:BCN

[978] W. A. Suchting. Berkeley’s criticism of Newton on space and motion. Isis, 58(2):186–197, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228223.

Christianson:1967:TBU

[979] John Christianson. Tycho Brahe at the University of Copenhagen, 1559–1562. Isis, 58(2):198–203, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228224.

Vlastos:1967:PST

[980] Gregory Vlastos. Plato’s supposed theory of irregular atomic figures. Isis, 58(2):204–209, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228225.

Bork:1967:MVP

[981] Alfred M. Bork. Maxwell and the vector potential. Isis, 58(2):210– 222, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228226.

Hooykaas:1967:EEJ

[982] R. Hooykaas. Eloge: Eduard Jan Dijksterhuis (1892–1965). Isis, 58 (2):223–225, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228227.

Cohen:1967:NUF

[983] I. Bernard Cohen. Newton’s use of “force,” or, Cajori versus Newton: A note on translations of the Principia. Isis, 58(2):226–230, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228228.

Brush:1967:NHF

[984] Stephen G. Brush, H. A. Lorentz, and George Francis FitzGerald. Note on the history of the FitzGerald–Lorentz contraction. Isis, 58(2):230– 232, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228229.

Heimann:1967:WHG

[985] P. M. Heimann and John L. Heilbron. “The work of H. G. J. Moseley”. Isis, 58(2):232–234, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 228230. See [896].

Cannon:1967:PMS

[986] Walter F. Cannon. “Parliament and the metric system”– a further comment. Isis, 58(2):235, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 228231. See [393, 836, 837].

Siegfried:1967:BDS

[987] Robert Siegfried. Boscovich and Davy: Some cautionary remarks. Isis, 58(2):236–238, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228232.

Suter:1967:WIJ

[988] Rufus Suter. The word Indiscerpible and Jonathan Edwards. Isis, 58 (2):238–239, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228233.

Wood:1967:HW

[989] James O. Wood. A horned woman. Isis, 58(2):239–240, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228234.

Leicester:1967:BLL

[990] Henry M. Leicester. Boyle, Lomonosov, Lavoisier, and the corpuscular theory of matter. Isis, 58(2):240–244, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228235.

Rousseau:1967:SAV

[991] G. S. Rousseau. Smollett’s acidum vagum. Isis, 58(2):244–245, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228236.

Barr:1967:BMF

[992] E. Scott Barr. Biographical material in the first series of the physical review. Isis, 58(2):245–246, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228237.

Anonymous:1967:Nb

[993] Anonymous. News. Isis, 58(2):247, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228238.

White:1967:BRB

[994] Lynn White, Jr. Book review: Science and Civilisation in China, by Joseph Needham; Wang Ling. Vol. 4: Physics and Physical Technology. Part II: Mechanical Engineering. Isis, 58(2):248–251, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228239.

Brush:1967:BRB

[995] Stephen G. Brush. Book review: The Wellesley Index to Victorian Periodicals 1824–1900 by Walter E. Houghton; Josef L. Altholz; Eileen Curran; Harold E. Dailey; Esther Roads Houghton; John A. Lester. Isis, 58(2):251–253, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228240.

Ballabh:1967:BRB

[996] R. Ballabh. Book review: A Bibliography of Sanskrit Works on Astronomy and Mathematics. Part I: Manuscripts, Texts, Translations, and Studies by S. N. Sen. Isis, 58(2):253–254, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228241.

Dolby:1967:BRB

[997] R. G. A. Dolby. Book review: Mind, Matter and Method. Essays in Philosophy and Science in Honor of Herbert Feigl by Paul K. Feverabend; Grover Maxwell. Isis, 58(2):254–255, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228242.

Ravetz:1967:BRB

[998] Jerome R. Ravetz. Book review: Science and Society by Norman Kaplan. Isis, 58(2):255–257, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228243.

Reingold:1967:BRBc

[999] Nathan Reingold. Book review: Science and Society in the United States by David D. Van Tassel; Michael G. Hall. Isis, 58(2):257–258, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228244.

Scott:1967:BRB

[1000] Wilson L. Scott. Book review: The Chemical Bond by J. J. Lagowski; Harold Hart. Isis, 58(2):258–259, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228245.

Williams:1967:BRB

[1001] L. Pearce Williams. Book review: The Concept of Energy by D. W. Theobald. Isis, 58(2):259–260, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228246.

Stannard:1967:BRBa

[1002] Jerry Stannard. Book review: Men with Green Pens. Lives of the Great Writers on Plants in Early Times by Louise Bush-Brown. Isis, 58(2): 260–261, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228247.

Coleman:1967:BRB

[1003] William Coleman. Book review: The Gene. A Critical History by Elof Axel Carlson. Isis, 58(2):261–263, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228248.

Blake:1967:BRB

[1004] John B. Blake. Book review: Medicine in America. Historical Essays by Richard Harrison Shryock. Isis, 58(2):263, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228249.

Hoch:1967:BRBa

[1005] J. Hampton Hoch. Book review: Drug Adulteration. Detection and Control in Nineteenth-Century Britain by Ernst W. Stieb; Glenn Sonnedecker. Isis, 58(2):263–264, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228250.

Calvert:1967:BRB

[1006] Monte A. Calvert. Book review: Men, Machines, and Modern Times by Elting E. Morison. Isis, 58(2):264–265, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228251.

Bedini:1967:BRBb

[1007] Silvio A. Bedini. Book review: Pioneers of Precision Timekeeping. Isis, 58(2):265–266, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228252.

Cannon:1967:BRBb

[1008] Walter F. Cannon. Book review: English Land Measuring to 1800: Instruments and Practices by A. W. Richeson. Isis, 58(2):266–267, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228253.

Durant:1967:BRBb

[1009] F. C. Durant, III. Book review: Collected Works: Vol I: Aerodynamics; Collected Works: Vol II: Reactive Flying Machines; Collected Works: Vol III: Dirigibles by K. E. Tsiolkovsky; Works on Rocket Technology by K. E. Tsiolkovsky; Problems of Flight by Jet Propulsion by F. A. Tsander; Pioneers of Rocket Technology by T. M. Mel’kumov; Rocket Flight Engineering by Eugen Sanger; Rockets in Planetary Space by Herman Oberth. Isis, 58(2):267–268, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228254.

Long:1967:BRB

[1010] E. Croft Long. Book review: The Midwife and the Witch by Thomas R. Forbes. Isis, 58(2):268–269, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228255.

Weisheipl:1967:BRB

[1011] James A. Weisheipl. Book review: Commentarius in VIII libros physicorum Aristotelis by Robert Grosseteste; Richard C. Dales. Isis, 58(2): 269–270, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228256.

Horne:1967:BRB

[1012] R. A. Horne. Book review: Physiologia. Epicuro-Gassendo-Charltoniana: Or a Fabrick of Science Natural, upon the Hypothesis of Atoms, Founded by Epicurus, Repaired by Petrus Gessendus, Augmented by Walter Charleton by Walter Charleton. Isis, 58(2):270–271, Summer 1967. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228257.

Truesdell:1967:BRBa

[1013] C. Truesdell. Book review: Manuscripta Euleriana Archivi Academiae Scientiarum URSS. Tomus I: Descriptio scientifica by Leonhard Euler; J. Ch. Kopelevic; M. V. Krutikova; G. K. Mikhailov; N. M. Raskin; G. A. Kn’asev; V. I. Smirnov; A. P. Juskevic. Isis, 58(2):271–273, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228258.

Truesdell:1967:BRBb

[1014] C. Truesdell. Book review: Manuscripta Euleriana Archivi Academiae Scientiarum URSS. Tomus II: Opera mechanica, Volumen I by Leonhard Euler; G. K. Mikhailov; I. A. Perelmuter; G. A. Knjazev; N. M. Raskin; V. I. Smirnov; A. P. Juskevic. Isis, 58(2):273–274, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228259.

Aiton:1967:BRB

[1015] E. J. Aiton. Book review: Galileo Reappraised by Carlo L. Golino; Homage to Galileo by Morton F. Kaplon. Isis, 58(2):274–275, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228260.

Roger:1967:BRB

[1016] Jacques Roger. Book review: Buffon avant l’“Histoire naturelle” by Lesley Hanks. Isis, 58(2):275–276, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228261.

Dunmore:1967:BRB

[1017] John Dunmore. Book review: La Austrialia del Espiritu Santo. The Journal of Fray Martin de Munilla O.F.M. and Other Documents Relating to the Voyage of Pedro Fernandez de Quiros to the South Sea (1605–1606) and the Franciscan Missionary Plan (1617–1627) by Celsus Kelly. Isis, 58(2):276–277, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228262.

Veith:1967:BRBa

[1018] Ilza Veith. Book review: The State and the Mentally Ill. A History of Worcester State Hospital in Massachusetts, 1830–1920 by Gerald N. Grob. Isis, 58(2):277–278, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228263.

Crosland:1967:BRB

[1019] Maurice Crosland. Book review: Jac. Berzelius. His Life and Work by J. Erik Jorpes; Barbara Steele. Isis, 58(2):278–279, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228264.

North:1967:BRBb

[1020] J. D. North. Book review: Frege’s Logical Theory by Robert Sternfeld. Isis, 58(2):279–280, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228265.

Grotjahn:1967:BRBa

[1021] Martin Grotjahn. Book review: A Psycho-Analytic Dialogue: The Letters of Sigmund Freud and Karl Abraham, 1907–1926 by Sigmund Freud; Karl Abraham; Hilda C. Abraham; Ernst L. Freud; Bernard Marsh. Isis, 58(2):280–282, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228266.

Pursell:1967:BRB

[1022] Carroll W. Pursell, Jr. Book review: An Administrative History of the National Aeronautics and Space Administration, 1958–1963 by Robert L. Rosholt. Isis, 58(2):282–283, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228267.

Heilbron:1967:BRB

[1023] John L. Heilbron. Book review: History of Science. Science in the Twentieth Century by Rene Taton; A. J. Pomerans. Isis, 58(2):283–284, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228268.

Anonymous:1967:BMb

[1024] Anonymous. Back matter. Isis, 58(2):285–287, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228269.

Anonymous:1967:FMc

[1025] Anonymous. Front matter. Isis, 58(3):289–291, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227987.

Daub:1967:AT

[1026] Edward E. Daub. Atomism and thermodynamics. Isis, 58(3):292–303, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227988.

Larson:1967:LNM

[1027] James L. Larson. Linnaeus and the natural method. Isis, 58(3):304– 320, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227989.

Lindberg:1967:ATV

[1028] David C. Lindberg. Alhazen’s theory of vision and its reception in the west. Isis, 58(3):321–341, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227990.

Vucinich:1967:MVS

[1029] Alexander Vucinich. Mendeleev’s views on science and society. Isis, 58 (3):342–351, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227991.

Schmitt:1967:EEA

[1030] Charles B. Schmitt. Experimental evidence for and against a void: The Sixteenth-Century arguments. Isis, 58(3):352–366, Autumn 1967. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227992.

Rauschenberg:1967:DCS

[1031] Roy A. Rauschenberg. Daniel Carl Solander, the naturalist on the endeavour voyage. Isis, 58(3):367–374, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227993.

Rappaport:1967:LGA

[1032] Rhoda Rappaport. Lavoisier’s geologic activities, 1763–1792. Isis, 58 (3):375–384, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227994.

DeSollaPrice:1967:GGS

[1033] Derek J. De Solla Price. A guide to graduate study and research in the history of science and medicine. Isis, 58(3):385–395, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227995.

Treloar:1967:ROI

[1034] F. E. Treloar. Ritual objects illustrating Indian alchemy and tantric religious practice. Isis, 58(3):396–397, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227996.

Geison:1967:DCS

[1035] Gerald L. Geison. Did Conon of Samos transmit Babylonian observations? Isis, 58(3):398–401, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227997.

Drake:1967:SCM

[1036] Stillman Drake. A Seventeenth-Century Malthusian. Isis, 58(3):401– 402, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227998.

Hall:1967:DHC

[1037] A. Rupert Hall and Richard S. Westfall. Did Hooke concede to Newton? Isis, 58(3):402–405, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227999.

Kopal:1967:HGJ

[1038] Zdenek Kopal. H. G. J. Moseley (1887–1915). Isis, 58(3):405–407, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228000.

Anonymous:1967:Nc

[1039] Anonymous. News. Isis, 58(3):408, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228001.

Kuhn:1967:RTR

[1040] Thomas S. Kuhn. Review: The turn to recent science: The Questioners: Physicists and the Quantum Theory by Barbara Lovett Cline Thirty Years That Shook Physics: The Story of Quantum Theory by George Gamow; The Conceptual Development of Quantum Mechanics by Max Jammer; Korrespondenz, Individualitat, und Komplementaritat: Eine Studie zur Geistesgeschichte der Quantentheorie in den Beitragen Niels Bohrs by Klaus Michael Meyer-Abich; Niels Bohr: The Man, His Science, and the World They Changed by Ruth Moore; Sources of Quantum Mechanics by B. L. Van der Waerden. Isis, 58(3):409–419, ???? 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228002;   http://www.jstor.
org/stable/pdfplus/228002.pdf.

McDaniel:1967:BRB

[1041] W. B. McDaniel. Book review: Index zur Geschichte der Medizin und Biologie.Bd. 2: 1949–1951/1952 by Johannes Steudel; Wilfried Ricker; Claus Nissen. Isis, 58(3):419–420, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228003.

Maier:1967:BRB

[1042] Clifford L. Maier. Book review: Physical Science, Its Structure and Development. Vol. I: From Geometric Astronomy to the Mechanical Theory of Heat by Edwin C. Kemble. Isis, 58(3):420–422, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228004.

Triolo:1967:BRB

[1043] Victor A. Triolo. Book review: A History of Parasitology by W. D. Foster. Isis, 58(3):422–423, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228005.

Needham:1967:BRB

[1044] Dorothy Needham. Book review: The History of Cell Respiration and Cytochrome by David Keilin. Isis, 58(3):423–424, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228006.

Mayr:1967:BRB

[1045] Ernst Mayr. Book review: Perspectives de la Zoologie Europeene by Jean Leclercq; Pierre Dagnelie. Isis, 58(3):424–425, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228007.

Huard:1967:BRB

[1046] P. Huard. Book review: Arzte und Medizin in Africa by Ludwig Brandl. Isis, 58(3):425–426, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228008.

Multhacy:1967:BRB

[1047] Robert P. Multhacy. Book review: Lexicon Technicum or an Universal English Dictionary of Arts and Sciences by John Harris. Isis, 58(3):426– 427, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228009.

Rousseau:1967:BRB

[1048] G. S. Rousseau. Book review: The Rhetoric of Science: A Study of Scientific Ideas and Imagery in Eighteenth-Century English Poetry by William Powell Jones. Isis, 58(3):427–429, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228010.

Layton:1967:BRB

[1049] Edwin Layton. Book review: The Mechanical Engineer in America, 1830–1910, Professional Cultures in Conflict by Monte A. Calvert. Isis, 58(3):429–431, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228011.

Reingold:1967:BRBd

[1050] Nathan Reingold. Book review: Statesmen of Science by J. G. Crowther. Isis, 58(3):431, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228012.

Almgren:1967:BRB

[1051] Beverly S. Almgren. Book review: Milestones of Modern Chemistry: Original Reports of the Discoveries by Eduard Farber. Isis, 58(3):432– 433, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228013.

Garratt:1967:BRB

[1052] G. R. M. Garratt. Book review: Le Telegraphe Optique de Claude Chappe by Henri Gachot. Isis, 58(3):433–434, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228014.

Altschuller:1967:BRB

[1053] G. Altschuller. Book review: Opty nad Rastitel’nymi Gibridami by Gregor Mendel. Isis, 58(3):434–436, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228015.

Carlson:1967:BRB

[1054] Elof Axel Carlson. Book review: The Origins of Genetics: A Mendel Source Book by Curt Stern; Eva R. Sherwood. Isis, 58(3):436–437, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228016.

Grotjahn:1967:BRBb

[1055] Martin Grotjahn. Book review: Briefwechsel by Sigmund Freud; Lou Andreas-Salome; Ernst Pfeiffer. Isis, 58(3):437–438, Autumn 1967. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228017.

Ackerknecht:1967:BRB

[1056] Erwin H. Ackerknecht. Book review: Contributions to the History of Medicine from the Bulletin of the New York Academy of Medicine, 1925– 1935 by Fielding H. Garrison. Isis, 58(3):438, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228018.

Wilson:1967:BRB

[1057] Robert Rathbun Wilson. Book review: The Development of High-Energy Accelerators by M. Stanley Livingston. Isis, 58(3):438–439, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228019.

Kaplan:1967:BRB

[1058] Norman Kaplan. Book review: In the Name of Science by H. L. Nieburg. Isis, 58(3):439–440, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228020.

Nielsen:1967:BRB

[1059] J. Rud Nielsen. Book review: Men of Physics: The Quintessence of Irving Langmuir by Albert Rosenfeld. Isis, 58(3):440–441, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228021.

Higham:1967:BRB

[1060] Robin Higham. Book review: This New Ocean: A History of Project Mercury by Loyd S. Swenson; James M. Grimwood; Charles C. Alexander. Isis, 58(3):441–443, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228022.

Anonymous:1967:BMc

[1061] Anonymous. Back matter. Isis, 58(3):444–446, Autumn 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228023.

Anonymous:1967:FMd

[1062] Anonymous. Front matter. Isis, 58(4):447–449, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228421.

Heilbron:1967:KST

[1063] John L. Heilbron. The Kossel–Sommerfeld theory and the ring atom. Isis, 58(4):450–485, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228422.

Miyasita:1967:LWT

[1064] Saburo Miyasita. A link in the westward transmission of Chinese anatomy in the Later Middle Ages. Isis, 58(4):486–490, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228423.

Briggs:1967:AEE

[1065] J. Morton Briggs, Jr. Aurora and Enlightenment Eighteenth-Century explanations of the Aurora Borealis. Isis, 58(4):491–503, Winter 1967.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228424.

Kibre:1967:GGB

[1066] Pearl Kibre. Giovanni Garzoni of Bologna (1419–1505), professor of medicine and defender of astrology. Isis, 58(4):504–514, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228425.

Easton:1967:EER

[1067] Joy B. Easton. The early editions of Robert Recorde’s Ground of Artes. Isis, 58(4):515–532, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228426.

Scott:1967:JBB

[1068] Frederick Scott, Herman Shapiro, and Jean Buridan. Jean Buridan’s De motibus animalium. Isis, 58(4):533–552, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228427.

Zirkle:1967:DGP

[1069] Conway Zirkle. The death of Gaius Plinius Secundus (23–79 a.d.). Isis, 58(4):553–559, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228428.

Anonymous:1967:CAB

[1070] Anonymous. Correction: The authorship of the Liber de crepusculis, an Eleventh-Century work on atmospheric refraction. Isis, 58(4):560, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228430. See [943].

Anonymous:1967:Nd

[1071] Anonymous. News. Isis, 58(4):560, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228429.

Hellman:1967:BRB

[1072] C. Doris Hellman. Book review: Kepler’s Somnium: The Dream, or Posthumous Work on Lunar Astronomy. Isis, 58(4):561–563, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228431.

Goetzmann:1967:BRB

[1073] William H. Goetzmann. Book review: The Bureau of American Ethnology. A Partial History by Neil M. Judd. Isis, 58(4):563, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228432.

Battison:1967:BRB

[1074] Edwin A. Battison. Book review: Les cadrans solaires de Max Elskamp by Henri Michel. Isis, 58(4):563–564, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228433.

Rothblatt:1967:BRB

[1075] Sheldon Rothblatt. Book review: Engineering at Cambridge University, 1783–1965 by T. J. N. Hilken. Isis, 58(4):564–565, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228434.

Jones:1967:BRB

[1076] W. Powell Jones. Book review: Flesh of Steel: Literature and the Machine in American Culture by Thomas Reed West. Isis, 58(4):565–566, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228435.

Baker:1967:BRB

[1077] A. Albert Baker, Jr. Book review: The Structure of Atoms by J. J. Lagowski; From Vital Force to Structural Formulas by O. Theodor Benfey. Isis, 58(4):566–568, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228436.

Hooykaas:1967:BRB

[1078] R. Hooykaas. Book review: Origins of the Science of Crystals by John G. Burke. Isis, 58(4):568–570, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228437.

Burstyn:1967:BRB

[1079] Harold L. Burstyn. Book review: Down to the Sea. A Century of Oceanography by J. R. Dean. Isis, 58(4):570, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228438.

Hoch:1967:BRBb

[1080] J. Hampton Hoch. Book review: Pharmaceutical Historiography by Alex Berman. Isis, 58(4):570–571, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228439.

Veith:1967:BRBb

[1081] Ilza Veith. Book review: Mille ans de chirurgie en occident: Ve–XVe siècles by Pierre Huard; M. D. Grmek. Isis, 58(4):571–572, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228440.

Ferguson:1967:BRB

[1082] Eugene S. Ferguson. Book review: Technology in Early America: Needs and Opportunities for Study by Brooke Hindle. Isis, 58(4):572–574, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228441.

Harris:1967:BRB

[1083] J. R. Harris. Book review: Architecture in Ancient Egypt and the Near East by A. Badawy. Isis, 58(4):574–575, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228442.

Talbot:1967:BRB

[1084] C. H. Talbot. Book review: Medicine. Humaine et veterinaire. A la fin du Moyen Age by Guy Beaujouan; Yvonne Poulle-Drieux; Jeanne-Marie Dureau-Lapeysonnie. Isis, 58(4):575–576, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228443.

Wallis:1967:BRB

[1085] Helen Wallis. Book review: Monumenta cartographica vetustiorls aevi. A.D. 1200–1500. Vol. I: Mappemondes by Marcel Destombes. Isis, 58 (4):576–578, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228444.

Gille:1967:BRB

[1086] Bertrand Gille. Book review: Images of the Universe. Leonardo da Vinci: The Artist as Scientist by Richard McLanathan. Isis, 58(4):578–579, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228445.

Stannard:1967:BRBb

[1087] Jerry Stannard. Book review: Libellus de Re Herbaria, 1538; The Names of Herbes, 1548 by William Turner. Isis, 58(4):579, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228446.

Berman:1967:BRB

[1088] Alex Berman. Book review: Medicine at the Paris Hospital, 1794–1848 by Erwin H. Ackerknecht. Isis, 58(4):580–581, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228447.

Goodman:1967:BRB

[1089] D. C. Goodman. Book review: The Atomic Debates by W. H. Brock. Isis, 58(4):581–582, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228448.

Snelders:1967:BRB

[1090] H. A. M. Snelders. Book review: Briefe eines romantischen Physikers: Johann Wilhelm Ritter an Gotthilf Heinrich Schubert und an Karl von Hardenberg by Friedrich Klemm; Armin Hermann. Isis, 58(4):582–583, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228449.

Maddock:1967:BRB

[1091] A. G. Maddock. Book review: A Scientific Autobiography by Otto Hahn; Willy Ley. Isis, 58(4):583–584, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228450.

Ahlers:1967:BRB

[1092] W. C. Ahlers. Book review: George Sarton. De Mens en Zijn Werk uit Brieven aan Vrienden en Kennissen by Paul van Oye. Isis, 58(4):584– 585, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228451.

Beer:1967:BRBb

[1093] John J. Beer. Book review: The Inner Ring: The Early History of the National Research Council of Canada by Mel Thistle. Isis, 58(4):585– 586, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228452.

Anonymous:1967:BMd

[1094] Anonymous. Back matter. Isis, 58(4):587–598, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228453.

Anonymous:1967:VI

[1095] Anonymous. Volume information. Isis, 58(4):589–595, Winter 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228420.

Multhauf:1967:NSC

[1096] Robert P. Multhauf, Bernard S. Finn, Aleida Cattell Renwick, Diana D. Menkes, and Virginia Skidmore Rutledge. Ninety-second critical bibliography of the history of science and its cultural influences (to 1 January 1967). Isis, 58(5):1–143, December 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228025.

Anonymous:1967:BMe

[1097] Anonymous. Back matter. Isis, 58(5):??, December 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228026.

Anonymous:1967:FMe

[1098] Anonymous. Front matter. Isis, 58(5):??, December 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228024.

Anonymous:1968:FMa

[1099] Anonymous. Front matter. Isis, 59(1):1–3, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227847.

Wilson:1968:KDE

[1100] Curtis Wilson. Kepler’s derivation of the elliptical path. Isis, 59(1):4– 25, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227848.

Gerstner:1968:JHT

[1101] Patsy A. Gerstner. James Hutton’s theory of the Earth and his theory of matter. Isis, 59(1):26–31, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227849.

Hewsen:1968:SSC

[1102] Robert H. Hewsen. Science in Seventh-Century Armenia: Ananias of Sirak. Isis, 59(1):32–45, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227850.

Lovell:1968:HDI

[1103] D. J. Lovell. Herschel’s dilemma in the interpretation of thermal radiation. Isis, 59(1):46–60, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227851.

Elkana:1968:HPC

[1104] Yehuda Elkana and June Goodfield. Harvey and the problem of the “capillaries”. Isis, 59(1):61–73, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227852.

Larder:1968:ECB

[1105] David F. Larder. The editions of Cardanus’ De rerum varietate. Isis, 59(1):74–77, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227853.

Vorzimmer:1968:DMH

[1106] Peter J. Vorzimmer. Darwin and Mendel: The historical connection. Isis, 59(1):77–82, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227854.

vanderWaerden:1968:EHT

[1107] B. L. van der Waerden. Eupalinos and his tunnel. Isis, 59(1):82–83, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227855.

Thiele:1968:BRH

[1108] Joachim Thiele. Briefe Robert h. lowies an Ernst Mach. Isis, 59(1):84– 87, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227856.

Bicknell:1968:DAO

[1109] P. J. Bicknell. Did Anaxagoras observe a sunspot in 467 B.C.? Isis, 59(1):87–90, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227857.

Forman:1968:ABC

[1110] Paul Forman, J. J. O. Palgen, Asger Aaboe, and Stephen Toulmin. “The astrophysics of Berossos the Chaldean”. Isis, 59(1):91–94, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227858. See [942, 1329].

Greene:1968:N

[1111] John C. Greene. News. Isis, 59(1):95–98, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227859.

Anonymous:1968:CRO

[1112] Anonymous. Corrections: Ritual objects illustrating Indian alchemy and tantric religious practice. Isis, 59(1):98, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227860.

Whitteridge:1968:BRB

[1113] Gweneth Whitteridge. Book review: The Life of William Harvey. Isis, 59(1):99–101, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227861.

Clarke:1968:BRB

[1114] Edwin Clarke. Book review: William Harvey’s Biological Ideas: Selected Aspects and Historical Background by Walter Pagel. Isis, 59(1):101– 102, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227862.

Bedini:1968:BRBa

[1115] Silvio A. Bedini. Book review: Alt-Augsburger Kompassmacher, Sonnen- , Mond- und Sternuhren, astronomische und mathematische Gerate, Raderuhren by Maximilian Bobinger. Isis, 59(1):102–104, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227863.

Leicester:1968:BRB

[1116] Henry M. Leicester. Book review: The Origins of Chemistry by Robert P. Multhauf. Isis, 59(1):104–105, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227864.

Schiller:1968:BRBa

[1117] Joseph Schiller. Book review: Essays in the History of Embryology and Biology by Jane M. Oppenheimer. Isis, 59(1):105–107, Spring 1968. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227865.

Poynter:1968:BRB

[1118] F. N. L. Poynter. Book review: Bibliography of the History of Medicine by National Library of Medicine. Isis, 59(1):107–108, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227866.

Meyer:1968:BRB

[1119] K. F. Meyer. Book review: A History of Immunization by H. J. Parish. Isis, 59(1):108–110, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227867.

Forbes:1968:BRBa

[1120] R. J. Forbes. Book review: Archeological Chemistry. A Symposium by Martin Levey. Isis, 59(1):110–111, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227868.

Stannard:1968:BRBa

[1121] Jerry Stannard. Book review: Hortulus: The Little Garden by Walahfrid Strabo; Raef Payne. Isis, 59(1):111–112, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227869.

Wertime:1968:BRB

[1122] Theodore A. Wertime. Book review: The Traditional Crafts of Persia. Their Development, Technology, and Influence on Eastern and Western Civilizations by Hans E. Wulff. Isis, 59(1):112–113, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227870.

Huard:1968:BRBa

[1123] P. Huard and Ming Wong. Book review: Alchemy, Medicine, and Religion in the China of A.D. 320: The Nei P’ien of Ko Hung (Pao-p’u tzu) by James R. Ware. Isis, 59(1):113–114, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227871.

Westfall:1968:BRB

[1124] Richard S. Westfall. Book review: His Autobiographical and Historical Notes. His Correspondence, and Other Contemporary Sources Relating to His Life and Work by Elias Ashmole; C. H. Josten. Isis, 59(1):114– 116, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227872.

Schofield:1968:BRB

[1125] Robert E. Schofield. Book review: Joseph Priestley: Revolutions of the Eighteenth Century by F. W. Gibbs. Isis, 59(1):116–117, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227873.

Forbes:1968:BRBb

[1126] Eric G. Forbes. Book review: John Harrison: The Man Who Found Longitude by Humphrey Quill. Isis, 59(1):117–118, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227874.

Reingold:1968:BRBa

[1127] Nathan Reingold. Book review: American Science in the Age of Jackson by George H. Daniels. Isis, 59(1):118–120, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227875.

Hirosige:1968:BRB

[1128] Tetu Hirosige. Book review: Sources for History of Quantum Physics. An Inventory and Report by Thomas S. Kuhn; John L. Heilbron; Paul Forman; Lini Allen. Isis, 59(1):120–121, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227876.

Nielsen:1968:BRB

[1129] J. Rud Nielsen. Book review: Niels Bohr. His Life and Work as Seen by His Friends and Colleagues by S. Rozental. Isis, 59(1):121–123, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227877.

Anonymous:1968:BMa

[1130] Anonymous. Back matter. Isis, 59(1):124–126, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227878.

Anonymous:1968:FMb

[1131] Anonymous. Front matter. Isis, 59(2):127–129, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228273.

Laudan:1968:VVC

[1132] L. L. Laudan. The vis viva controversy, a post-mortem. Isis, 59(2):130– 143, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228274.

Ikle:1968:SAS

[1133] Frank W. Ikle. Sir Aurel Stein. A Victorian geographer in the tracks of Alexander. Isis, 59(2):144–155, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228275.

Forman:1968:DRA

[1134] Paul Forman. The doublet riddle and atomic physics circa 1924. Isis, 59(2):156–174, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228276.

Egerton:1968:ASA

[1135] Frank N. Egerton, III. Ancient sources for animal demography. Isis, 59 (2):175–189, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228277.

Pottage:1968:VVI

[1136] John Pottage. The vitruvian value of II. Isis, 59(2):190–197, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228278.

Fulford:1968:PHO

[1137] George D. Fulford. Pouring holy oil on troubled waters. Isis, 59(2):198– 199, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228279.

Watson:1968:RDH

[1138] Robert I. Watson. Recent developments in the historiography of American psychology. Isis, 59(2):199–205, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228280.

Anonymous:1968:Na

[1139] Anonymous. News. Isis, 59(2):206, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228281.

Hughes:1968:BRB

[1140] Thomas P. Hughes. Book review: Technology in Western Civilisation. Vol. I: The Emergence of Modern Industrial Society, Earliest Times to 1900 by Melvin Kranzberg; Carroll W. Pursell. Isis, 59(2):207–208, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228282.

Williams:1968:BRB

[1141] Trevor I. Williams. Book review: Technology in Western Civilization. Vol. II: Technology in the Twentieth Century by Melvin Kranzberg; Carroll W. Pursell. Isis, 59(2):209–210, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228283.

Goerke:1968:BRBa

[1142] H. Goerke. Book review: Lychnos. Lardomshistoriska Samfundets Arsbok. 1965–1966. Isis, 59(2):210–211, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228284.

Pagel:1968:BRB

[1143] Walter Pagel. Book review: Catalogue of Printed Books in the Wellcome Historical Medical Library. Vol. II: Books Printed from 1641 to 1850, A-E. Isis, 59(2):211, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228285.

Rousseau:1968:BRB

[1144] G. S. Rousseau. Book review: The Royal Society: Concept and Creation by Margery Purver. Isis, 59(2):211–213, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228286.

Bedini:1968:BRBb

[1145] Silvio A. Bedini. Book review: Scientific Instruments in Art and History by Henri Michel; R. E. W. Maddison; Francis R. Maddison. Isis, 59 (2):213–214, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228287.

Korber:1968:BRB

[1146] Hans-Gunther Korber. Book review: A History of the Thermometer and Its Uses in Meteorology by W. E. Knowles Middleton. Isis, 59(2):214– 216, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228288.

Reingold:1968:BRBb

[1147] Nathan Reingold. Book review: Science, Technology, and American Foreign Policy by Eugene B. Skolnikoff; Reviews of National Science Policy: United States by Organization for Economic Co-operation and Development; The Politics of Pure Science by Daniel S. Greenberg. Isis, 59 (2):216–220, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228289.

Buttmann:1968:BRB

[1148] Gunther Buttmann. Book review: Bau und Bildung des Weltalls. Kosmologische Vorstellungen in Dokumenten aus zwei Jahrtausenden by Bernhard Sticker. Isis, 59(2):220–221, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228290.

Carozzi:1968:BRB

[1149] Albert V. Carozzi. Book review: An Annotated Bibliography of Observations on Illinois Water Resources, 1673–1850 by Thomas P. Schlunz; Robert M. Sutton; George W. White. Isis, 59(2):222–223, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228291.

Vorzimmer:1968:BRBa

[1150] Peter J. Vorzimmer. Book review: Origins of Mendelism by R. C. Olby. Isis, 59(2):223–224, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228292.

Jackson:1968:BRB

[1151] Stanley W. Jackson. Book review: Origins of Modern Psychiatry by Ernest Harms. Isis, 59(2):224–225, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228293.

Stahl:1968:BRB

[1152] W. H. Stahl. Book review: Agricultural Implements of the Roman World by K. D. White. Isis, 59(2):225–226, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228294.

McVaugh:1968:BRB

[1153] Michael R. McVaugh. Book review: Medicine in Medieval England by C. H. Talbot. Isis, 59(2):227–228, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228295.

Durling:1968:BRB

[1154] Richard J. Durling. Book review: Les manuscrits latins de medecine du haut Moyen Age dans les bibliothèques de France by Ernest Wickersheimer. Isis, 59(2):228–229, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228296.

Sivin:1968:BRB

[1155] N. Sivin. Book review: Huang Ti Nei Ching Su Wen. The Yellow Emperor’s Classic of Internal Medicine by Ilza Veith. Isis, 59(2):229–231, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228297.

Rosen:1968:BRB

[1156] Edward Rosen. Book review: Georg Joachim Rhetikus, 1514–1574. Eine Bio-Bibliographie. Band I: Humanist und Wegbereiter der modernen Naturwissenschaften by Karl Heinz Burmeister. Isis, 59(2):231–233, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228298.

Bullough:1968:BRB

[1157] Vern L. Bullough. Book review: A Catalogue of Sixteenth Century Printed Books in the National Library of Medicine by Richard J. Durling. Isis, 59(2):233–234, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228299.

Hamarneh:1968:BRB

[1158] Sami Hamarneh. Book review: Dlicht d’Apotekers by Quiricus de Augustis. Isis, 59(2):234–235, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228300.

Aiton:1968:BRBa

[1159] E. J. Aiton. Book review: Dialogue concerning the Two Chief World Systems by Galileo Galilei; Stillman Drake. Isis, 59(2):235–236, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228301.

Smeaton:1968:BRB

[1160] W. A. Smeaton. Book review: A Scientific Autobiography of Joseph Priestley, 1733–1804. Selected Scientific Correspondence, Edited with a Commentary by Robert E. Schofield. Isis, 59(2):236–237, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228302.

Kantor:1968:BRB

[1161] J. R. Kantor. Book review: Briefe–Correspondance, 1864–1927 by August Forel; H. H. Walser. Isis, 59(2):237–238, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228303.

Schiller:1968:BRBb

[1162] J. Schiller. Book review: Claude Bernard and Experimental Medicine by Francisco Grande; Maurice B. Visscher. Isis, 59(2):238–239, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228304.

Forman:1968:BRB

[1163] Paul Forman. Book review: My Life and My Views, by Max Born. Isis, 59(2):239–240, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228305;
 http://www.jstor.org/stable/pdfplus/228305.pdf.

Corrsin:1968:BRB

[1164] S. Corrsin. Book review: The Wind and Beyond: Theodore von Karman, Pioneer in Aviation and Pathfinder in Space by Theodore von Karman; Lee Edson. Isis, 59(2):240–242, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228306.

Heilbron:1968:BRBa

[1165] John L. Heilbron. Book review: Erwin Schrödinger. An Introduction to His Writings by William T. Scott. Isis, 59(2):242–243, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228307.

Anonymous:1968:BMb

[1166] Anonymous. Back matter. Isis, 59(2):244–246, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228308.

Anonymous:1968:FMc

[1167] Anonymous. Front matter. Isis, 59(3):247–248, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227879.

Young:1968:FBG

[1168] Robert M. Young. The functions of the brain: Gall to Ferrier (1808– 1886). Isis, 59(3):250–268, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227880.

Kren:1968:HAL

[1169] Claudia Kren. Homocentric astronomy in the Latin West. the De reprobatione ecentricorum et epiciclorum of Henry of Hesse. Isis, 59(3):269– 281, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227881.

Mellor:1968:MAS

[1170] D. H. Mellor. Models and analogies in science: Duhem versus Campbell? Isis, 59(3):282–290, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227882.

Larson:1968:SCL

[1171] James L. Larson. The species concept of Linnaeus. Isis, 59(3):291– 299, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227883.

Woodruff:1968:CHH

[1172] A. E. Woodruff. The contributions of Hermann von Helmholtz to electrodynamics. Isis, 59(3):300–311, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227884.

Christianson:1968:TBC

[1173] John Christianson and Tycho Brahe. Tycho Brahe’s cosmology from the astrologia of 1591. Isis, 59(3):312–318, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227885.

Robinson:1968:EDM

[1174] Eric H. Robinson. Eloge: Douglas McKie (15 July 1896–28 August 1967). Isis, 59(3):319–327, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227886.

Anonymous:1968:Nb

[1175] Anonymous. News. Isis, 59(3):328, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227887.

Egerton:1968:BRB

[1176] Frank N. Egerton, III. Book review: The Komarov Botanical Institute, by Stanwyn G. Shetler. 250 Years of Russian Research. Isis, 59(3):329– 332, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227888.

Brumbaugh:1968:BRB

[1177] Robert S. Brumbaugh. Book review: Traces on the Rhodian Shore: Nature and Culture in Western Thought from Ancient Times to the End of the Eighteenth Century by Clarence J. Glacken. Isis, 59(3):332–333, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227889.

Kevles:1968:BRB

[1178] Daniel J. Kevles. Book review: National Academy of Sciences, Biographical Memoirs. Vol. XXXIX. Isis, 59(3):333–334, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227890.

Stannard:1968:BRBb

[1179] Jerry Stannard. Book review: History of Four-Footed Beasts and Serpents and Insects by Edward Topsell; T. Muffet. Isis, 59(3):334–335, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227891.

Meisel:1968:BRB

[1180] Max Meisel. Book review: Medical Reference Works, 1679–1966. A Selected Bibliography by John B. Blake; Charles Roos. Isis, 59(3):335–336, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227892.

Vorzimmer:1968:BRBb

[1181] Peter Vorzimmer. Book review: Bulletin de la Société Française d’Histoire de la Medecine. Tome Ier, 1902. Isis, 59(3):336–338, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227893.

Huard:1968:BRBb

[1182] P. Huard. Book review: Magie, medecine et possession en Ethiopie by Maxime Rodinson. Isis, 59(3):338–339, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227894.

Skramstad:1968:BRB

[1183] Harold K. Skramstad. Book review: American Building. 1: The Historical Forces That Shaped It by James Marston Fitch. Isis, 59(3):339–340, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227895.

Grmek:1968:BRBa

[1184] M. D. Grmek. Book review: Pythagoras and Early Pythagoreanism. An Interpretation of Neglected Evidence on the Philosopher Pythagoras by C. J. De Vogel. Isis, 59(3):340–341, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227896.

Grant:1968:BRB

[1185] Edward Grant. Book review: Hugonis de Sancto Victore Opera Propaedeutica: Practica geometriae, De grammatica, Epitome Dindimi in philosophiam by Roger Baron. Isis, 59(3):342–343, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227897.

Heilbron:1968:BRBb

[1186] John L. Heilbron. Book review: Le origini del metodo analitico. Il Cinquecento by Angelo Crescini. Isis, 59(3):343–344, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227898.

Jayawardene:1968:BRB

[1187] S. A. Jayawardene. Book review: Regiomontanus on Triangles. De triangulis omnimodis by Johann Muller; Barnabas Hughes. Isis, 59(3):344– 345, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227899.

Davis:1968:BRB

[1188] Audrey Davis. Book review: Thomas Willis, 1621–1675. Doctor and Scientist by Hansruedi Isler. Isis, 59(3):345–347, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227900.

Ospovat:1968:BRB

[1189] Alexander M. Ospovat. Book review: On the Origin of Springs by Pierre Perrault; Aurele LaRocque. Isis, 59(3):347–348, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227901.

Goerke:1968:BRBb

[1190] H. Goerke. Book review: A Naturalist in Russia. Letters from Peter Simon Pallas to Thomas Pennant by Carol Urness. Isis, 59(3):348–349, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227902.

Holmes:1968:BRBa

[1191] Frederic L. Holmes. Book review: Lecons sur les Phenomenes de la vie communs aux animaux et aux vegetaux by Claude Bernard. Isis, 59(3): 349–350, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227903.

Holmes:1968:BRBb

[1192] Frederic L. Holmes. Book review: Claude Bernard et les problemes scientifiques de son temps by Joseph Schiller. Isis, 59(3):350–352, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227904.

Woodruff:1968:BRB

[1193] A. E. Woodruff. Book review: Wilhelm Eduard Weber: Erforscher der Wellenbewegung und der Elektrizitat 1804–1891 by K. H. Wiederkehr. Isis, 59(3):352–353, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227905.

Stuewer:1968:BRBa

[1194] Roger H. Stuewer. Book review: The Cosmos of Arthur Holly Compton by Marjorie Johnston. Isis, 59(3):354–355, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227906.

Anonymous:1968:BMc

[1195] Anonymous. Back matter. Isis, 59(3):356–357, Autumn 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227907.

Anonymous:1968:FMd

[1196] Anonymous. Front matter. Isis, 59(4):359–362, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228486.

May:1968:GQM

[1197] Kenneth O. May. Growth and quality of the mathematical literature. Isis, 59(4):363–371, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228487.

Gray:1968:EAB

[1198] Philip Howard Gray. The early animal behaviorists: Prolegomenon to ethology. Isis, 59(4):372–383, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228488.

Wallace:1968:EDS

[1199] William A. Wallace. The enigma of Domingo de Soto: Uniformiter difformis and falling bodies in Late Medieval physics. Isis, 59(4):384–401, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228489.

Corson:1968:PPF

[1200] David W. Corson. Pierre Polinière, Francis Hauksbee, and electroluminescence: A case of simultaneous discovery. Isis, 59(4):402–413, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228490.

Ariotti:1968:GIP

[1201] Piero Ariotti. Galileo on the isochrony of the pendulum. Isis, 59(4):414– 426, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228491.

Kevles:1968:GEH

[1202] Daniel J. Kevles. George Ellery Hale, the First World War, and the advancement of science in America. Isis, 59(4):427–437, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228492.

Pelseneer:1968:CAC

[1203] Jean Pelseneer. Centenaire de l’affaire Chasles. (French) [centenary of the Chasles Affair]. Isis, 59(4):438–439, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228493.

Anonymous:1968:Nc

[1204] Anonymous. News. Isis, 59(4):440–441, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228494.

Needham:1968:BRB

[1205] Joseph Needham. Book review: Marcello Malpighi and the Evolution of Embryology. Isis, 59(4):442–443, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228495.

Crowe:1968:BRB

[1206] Michael J. Crowe. Book review: Science and Culture. Volume I: Time, Space and Motion by Noah Edward Fehl. Isis, 59(4):444–445, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228496.

Stuewer:1968:BRBb

[1207] Roger H. Stuewer. Book review: Philosophy of Science Today by Sidney Morgenbesser. Isis, 59(4):445–446, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228497.

Schaffner:1968:BRB

[1208] Kenneth F. Schaffner. Book review: The Foundations of Scientific Inference by Wesley C. Salmon. Isis, 59(4):447–448, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228498.

Rogers:1968:BRB

[1209] Frank B. Rogers. Book review: Bibliography of the History of Medicine, No. 2, 1966 by National Library of Medicine. Isis, 59(4):448–449, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228499.

Heer:1968:BRB

[1210] Nicholas Heer. Book review: Science and Civilization in Islam by Seyyed Hossein Nasr. Isis, 59(4):449–451, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228500.

Aiton:1968:BRBb

[1211] E. J. Aiton. Book review: Galilee. Aspects de sa vie et de son oeuvre by Centre International de Synthese; Galileo, Man of Science by Ernan McMullin. Isis, 59(4):451–452, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228501.

Drake:1968:BRB

[1212] Stillman Drake. Book review: Galilee: Dialogues by Paul-Henri Michel; Lettres choisies by Paul-Henri Michel. Isis, 59(4):452–453, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228502.

Kreiling:1968:BRB

[1213] Frederick C. Kreiling. Book review: Leibniz. Sein Leben, Sein Wirken, Seine Welt by Wilhelm Totok; Carl Haase. Isis, 59(4):453–454, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228503.

Kauffman:1968:BRB

[1214] George B. Kauffman. Book review: Oxygen and Oxidation — Theories and Techniques in the 19th Century and the First Part of the 20th by Eduard Farber. Isis, 59(4):454–455, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228504.

Wilson:1968:BRB

[1215] Leonard G. Wilson. Book review: Tom Rivers. Reflections on a Life in Medicine and Science. An Oral History Memoir by Saul Benison. Isis, 59(4):455–458, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228505.

Badash:1968:BRB

[1216] Lawrence Badash. Book review: An American Genius: The Life of Ernest Orlando Lawrence, Father of the Cyclotron by Herbert Childs. Isis, 59(4):458–460, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228506.

Kruta:1968:BRB

[1217] Vladislav Kruta. Book review: Charles Scott Sherrington. A Biography of the Neurophysiologist by Ragnar Granit. Isis, 59(4):460–462, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228507.

Duncan:1968:BRB

[1218] Francis Duncan. Book review: The German Atomic Bomb. The History of Nuclear Research in Nazi Germany by David Irving. Isis, 59(4):462– 463, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228508.

Grmek:1968:BRBb

[1219] M. D. Grmek. Book review: Autobiographical Writings by Henry E. Sigerist; Nora Sigerist Beeson; A Bibliography of the Writings of Henry E. Sigerist by Genevieve Miller. Isis, 59(4):463–464, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228509.

Allen:1968:BRB

[1220] Garland E. Allen. Book review: The Double Helix by James D. Watson. Isis, 59(4):464–466, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228510.

Anonymous:1968:BMd

[1221] Anonymous. Back matter. Isis, 59(4):467–478, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228511.

Anonymous:1968:VI

[1222] Anonymous. Volume information. Isis, 59(4):469–474, Winter 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228485.

Anonymous:1968:FMe

[1223] Anonymous. Front matter. Isis, 59(5):1–4, ???? 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228270.

Neu:1968:NTC

[1224] John Neu. Ninety-third critical bibliography of the history of science and its cultural influences (to January 1968). Isis, 59(5):5–241, ???? 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228271.

Anonymous:1968:BMe

[1225] Anonymous. Back matter. Isis, 59(5):242, ???? 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228272.

Anonymous:1969:FMa

[1226] Anonymous. Front matter. Isis, 60(1):1–3, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229019.

MacLeod:1969:SGV

[1227] Roy M. MacLeod. Science and government in Victorian England: Lighthouse illumination and the Board of Trade, 1866–1886. Isis, 60(1):4–38, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229020.

Grant:1969:MSC

[1228] Edward Grant. Medieval and Seventeenth-Century conceptions of an infinite void space beyond the Cosmos. Isis, 60(1):39–60, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229021.

Mauskopf:1969:AST

[1229] Seymour H. Mauskopf. The atomic structural theories of Ampère and Gaudin: Molecular speculation and Avogadro’s hypothesis. Isis, 60(1):

61–74, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229022.

Aiton:1969:KSL

[1230] E. J. Aiton. Kepler’s second law of planetary motion. Isis, 60(1):75– 90, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229023.

Olson:1969:RBI

[1231] Richard Olson. The reception of Boscovich’s ideas in Scotland. Isis, 60(1):91–103, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229024.

Sambursky:1969:EME

[1232] S. Sambursky. The equivalence of mass and energy. An anticipation by Mendeleev. Isis, 60(1):104–106, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229025.

Greene:1969:N

[1233] John C. Greene. News. Isis, 60(1):107–108, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229026.

Coleman:1969:BRBa

[1234] William Coleman. Book review: Mesmerism and the End of the Enlightenment in France by Robert Darnton. Isis, 60(1):109–111, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229027.

Hall:1969:BRB

[1235] Marie Boas Hall. Book review: Metaphysics and Measurement. Essays in Scientific Revolution by Alexandre Koyré; Études d’histoire de la pensée scientifique by Alexandre Koyré. Isis, 60(1):111–112, Spring 1969. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229028.

May:1969:BRB

[1236] Kenneth O. May. Book review: Mathematics and Logic. Retrospect and Prospect by Mark Kac; Stanislaw Ulam. Isis, 60(1):112–113, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229029.

Brock:1969:BRB

[1237] W. H. Brock. Book review: Discovery of the Elements by Mary Elvira Weeks; Henry M. Leicester. Isis, 60(1):113–114, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229030.

Kantor:1969:BRB

[1238] J. R. Kantor. Book review: The Great Psychologists, from Aristotle to Freud by Robert I. Watson. Isis, 60(1):114–115, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229031.

Beer:1969:BRB

[1239] John J. Beer. Book review: Ideas in Exile. A History of Canadian Invention by J. J. Brown. Isis, 60(1):115–117, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229032.

Rosen:1969:BRBa

[1240] Edward Rosen. Book review: Georg Joachim Rhetikus, 1514–1574. Eine Bio-Bibliographie. Band II: Quellen und Bibliographie by Karl Heinz Burmeister. Isis, 60(1):117–119, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229033.

Lindberg:1969:BRB

[1241] David C. Lindberg. Book review: Physico-mathesis de lumine, coloribus, et iride by Francesco Maria Grimaldi. Isis, 60(1):119, Spring 1969. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229034.

Middleton:1969:BRB

[1242] W. E. Knowles Middleton. Book review: L’expérience du mouvement. Jean-Baptiste Baliani, disciple et critique de Galilée by Serge Moscovici. Isis, 60(1):120–121, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229035.

Smith:1969:BRBa

[1243] T. Arthur Smith. Book review: Obras Completas. Tomo III by Carlos J. Finlay. Isis, 60(1):121–122, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229036.

Egerton:1969:BRB

[1244] Frank N. Egerton, III. Book review: Alexander Wilson, Wanderer in the Wilderness by Robert Plate. Isis, 60(1):122–123, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229037.

Fox:1969:BRB

[1245] Robert Fox. Book review: Raspail: Scientist and Reformer by Dora B. Weiner. Isis, 60(1):123–124, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229038.

Brozek:1969:BRB

[1246] Josef Brožek. Book review: The Soviet Academy of Sciences and the Communist Party, 1927–1932 by Loren Graham. Isis, 60(1):124–126, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229039.

Farber:1969:BRB

[1247] Eduard Farber. Book review: Source Book in Chemistry 1900–1950 by Henry M. Leicester. Isis, 60(1):126, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229040.

Anonymous:1969:BMa

[1248] Anonymous. Back matter. Isis, 60(1):127–128, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229041.

Anonymous:1969:FMb

[1249] Anonymous. Front matter. Isis, 60(2):129–131, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229906.

Holton:1969:EMC

[1250] Gerald Holton. Einstein, Michelson, and the “crucial” experiment. Isis, 60(2):133–197, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229907.

Hoare:1969:ATQ

[1251] Michael E. Hoare. “All things are queer and opposite”: Scientific societies in Tasmania in the 1840’s. Isis, 60(2):198–209, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229908.

Suter:1969:SWA

[1252] Rufus Suter and Allesandro Piccolomini. The scientific work of Allesandro Piccolomini. Isis, 60(2):210–222, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229909.

Clagett:1969:PBN

[1253] Marshall Clagett. Prosdocimus de Beldomandis and Nicole Oresme’s proof of the Merton rule of uniformly difform. Isis, 60(2):223–225, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229910.

Crosland:1969:CDM

[1254] Maurice Crosland. The Congress on Definitive Metric Standards, 1798– 1799: The First International Scientific Conference? Isis, 60(2):226–231, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229911.

Jeffries:1969:BNB

[1255] Theodore W. Jeffries. A biographical note on Benjamin Smith Barton (1766–1815). Isis, 60(2):231–232, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229912.

Ferreira:1969:PM

[1256] Ricardo Ferreira. Photographs of Moseley. Isis, 60(2):233, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229913.

Schrynemakers:1969:DWD

[1257] Arthur H. Schrynemakers. Descartes and the weight-driven chain-clock. Isis, 60(2):233–236, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229914.

Crouch:1969:QCB

[1258] James Crouch. Query: Coomaraswamy bibliography. Isis, 60(2):236, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229915.

Silverstein:1969:QEC

[1259] Albert Silverstein. Query: The endocrinologist in “the creeping man”. Isis, 60(2):237, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229916.

Anonymous:1969:Na

[1260] Anonymous. News. Isis, 60(2):238–239, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229917.

Burckhardt:1969:BRB

[1261] J. J. Burckhardt. Book review: Ibn al-Muthanna’s Commentary on the Astronomical Tables of al-Khwarizmi. Isis, 60(2):240–242, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229918.

Eklund:1969:BRB

[1262] Jon B. Eklund. Book review: A Dictionary of English Weights and Measures. From Anglo-Saxon Times to the Nineteenth Century by Ronald Edward Zupko. Isis, 60(2):242, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229919.

Chojecka:1969:BRB

[1263] Ewa Chojecka. Book review: The Pentagram as a Medical Symbol. An Iconological Study by J. Schouten. Isis, 60(2):242–244, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229920.

Hellman:1969:BRB

[1264] C. Doris Hellman. Book review: Dialexis de novae et prius incognitae stellae inusitatae magnitudinis & splendidissimi luminis apparitione, & de eiusdem stellae vero loco constituendo by Thaddaeus Hagecius ab Hayck; Zdenek Horský; Thaumantias Liber de arcu coelesti deque colorum apparentium natura, ortu et causis by Johannes Marcus Marci; Jiří Marek. Isis, 60(2):244–246, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229921.

Merriman:1969:BRB

[1265] Daniel Merriman. Book review: Premier Congrès International d’Histoire de l’Océanographie. Monaco–1966 by Jacqueline Carpine-Lancre; John Leighly. Isis, 60(2):246–248, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229922.

Barber:1969:BRB

[1266] Bernard Barber. Book review: The Development of Medicine as a Profession. The Contribution of the Medieval University to Modern Medicine by Vern L. Bullough; Medical Licensing in America, 1650–1965 by Richard Harrison Shryock; The Formation of the American Medical Profession. The Role of Institutions, 1780–1860 by Joseph F. Kett. Isis, 60(2):248– 250, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229923.

Davis:1969:BRBa

[1267] Audrey B. Davis. Book review: La Chirurgie moderne–ses débuts en occident: XVIe, XVIIe, XVIIIe siècles by Pierre Huard; Mirko Drazen Grmek; An Introduction to the History of General Surgery by Richard Hardaway Meade. Isis, 60(2):250–251, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229924.

Bullough:1969:BRB

[1268] Vern L. Bullough. Book review: The Royal Apothecaries by Leslie G. Matthews. Isis, 60(2):251–252, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229925.

Galdston:1969:BRB

[1269] Iago Galdston. Book review: Geschichte der Medizin. Einführung in ihre Grundzüge by Alexander Mette; Irena Winter. Isis, 60(2):252–253, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229926.

Condit:1969:BRB

[1270] Carl W. Condit. Book review: Matrix of Man: An Illustrated History of the Urban Environment by Sibyl Moholy-Nagy. Isis, 60(2):253–255, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229927.

Kudlien:1969:BRB

[1271] Fridolf Kudlien. Book review: Ancient Medicine by Ludwig Edelstein; Owsei Temkin; C. Lilian Temkin. Isis, 60(2):255, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229928.

Saidan:1969:BRB

[1272] A. S. Saidan. Book review: Probability and Opinion. A Study of the Medieval Presuppositions of Post-Medieval Theories of Probability by Edmund F. Byrne. Isis, 60(2):255–257, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229929.

Multhauf:1969:BRB

[1273] Robert P. Multhauf. Book review: Book of Minerals by Albertus Magnus; Dorothy Wyckoff. Isis, 60(2):257–258, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229930.

Jones:1969:BRB

[1274] William Powell Jones. Book review: This Long Disease, My Life: Alexander Pope and the Sciences by Marjorie Nicolson; G. S. Rousseau. Isis, 60(2):258–259, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229931.

Veith:1969:BRBa

[1275] Ilza Veith. Book review: The Construction and Government of Lunatic Asylums and Hospitals for the Insane by John Conolly. Isis, 60(2):259– 261, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229932.

Richardson:1969:BRB

[1276] R. Alan Richardson. Book review: Darwinism Comes to America by George Daniels. Isis, 60(2):261, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229933.

Turner:1969:BRB

[1277] G. L’E. Turner. Book review: Alvan Clark & Sons, Artists in Optics by Deborah Jean Warner. Isis, 60(2):261–262, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229934.

Goodfield:1969:BRB

[1278] June Goodfield. Book review: Philosophie et méthodologie scientifiques de Claude Bernard by Etienne Wolff. Isis, 60(2):262–264, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229935.

Berendzen:1969:BRB

[1279] Richard Berendzen. Book review: The Story of Jodrell Bank by Bernard Lovell. Isis, 60(2):264–265, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229936.

Anonymous:1969:BMb

[1280] Anonymous. Back matter. Isis, 60(2):266–268, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229937.

Anonymous:1969:FMc

[1281] Anonymous. Front matter. Isis, 60(3):269–271, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229482.

Geison:1969:PTL

[1282] Gerald L. Geison. The protoplasmic theory of life and the vitalistmechanist debate. Isis, 60(3):272–292, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229483.

McCormmach:1969:HCS

[1283] Russell McCormmach. Henry Cavendish: A study of rational empiricism in Eighteenth-Century natural philosophy. Isis, 60(3):293–306, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229484.

Pursell:1969:FCC

[1284] Carroll W. Pursell, Jr. The Farm Chemurgic Council and the United States Department of Agriculture, 1935–1939. Isis, 60(3):307–317, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229485.

Daub:1969:PTR

[1285] Edward E. Daub. Probability and thermodynamics: The reduction of the second law. Isis, 60(3):318–330, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229486.

Skabelund:1969:WOM

[1286] Donald Skabelund and Phillip Thomas. Walter of Odington’s mathematical treatment of the primary qualities. Isis, 60(3):331–350, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229487.

Parascandola:1969:HPT

[1287] John Parascandola and Aaron J. Ihde. History of the pneumatic trough. Isis, 60(3):351–361, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229488.

Finn:1969:AOF

[1288] Bernard S. Finn. An appraisal of the origins of Franklin’s electrical theory. Isis, 60(3):362–369, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229489.

McKinney:1969:WEO

[1289] H. Lewis McKinney. Wallace’s earliest observations on evolution: 28 December 1845. Isis, 60(3):370–373, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229490.

Morris:1969:LFA

[1290] Robert J. Morris, Jr. and Henry Guerlac. Lavoisier on fire and air: The memoir of July 1772. Isis, 60(3):374–382, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229491.

Clagett:1969:NCF

[1291] Marshall Clagett. A note on the commensurator falsely attributed to Regiomontanus. Isis, 60(3):383–384, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229492.

Forman:1969:AFV

[1292] Paul Forman. The archive of Friedr. Vieweg & Sohn, Braunschweig: Correspondence of Nineteenth-Century physical scientists, especially German chemists. Isis, 60(3):384–386, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229493.

Kilburn:1969:PG

[1293] Paul D. Kilburn. Plants of the Galapagos. Isis, 60(3):386–388, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229494.

Sivin:1969:PPT

[1294] N. Sivin. On the pao p’u tzu nei p’ien and the life of ko hung (283–343). Isis, 60(3):388–391, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229495.

Stuewer:1969:WNW

[1295] Roger H. Stuewer. Was Newton’s “wave-particle duality” consistent with Newton’s observations? Isis, 60(3):392–394, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229496.

Anonymous:1969:Nb

[1296] Anonymous. News. Isis, 60(3):395, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229497.

Rousseau:1969:BRB

[1297] G. S. Rousseau. Book review: Literature and Technology, by Wylie Sypher. The Alien Vision. Isis, 60(3):396–397, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229498.

Maier:1969:BRB

[1298] Clifford Maier. Book review: Steps in the Scientific Tradition. Readings in the History of Science by Richard S. Westfall; Victor E. Thoren. Isis, 60(3):398–399, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229499.

Platt:1969:BRB

[1299] John Platt. Book review: Approaches to a Philosophical Biology by Marjorie Grene. Isis, 60(3):399–400, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229500.

Smith:1969:BRBb

[1300] Cecil O. Smith, Jr. Book review: France in the Age of the Scientific State by Robert Gilpin. Isis, 60(3):400–402, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229501.

Veith:1969:BRBb

[1301] Ilza Veith. Book review: Documentary History of Psychiatry. A Source Book on Historical Principles by Charles E. Goshen. Isis, 60(3):402– 403, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229502.

Weyant:1969:BRB

[1302] Robert G. Weyant. Book review: Madness in Society: Chapters in the Historical Sociology of Mental Illness by George Rosen. Isis, 60(3):403– 404, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229503.

Theodorides:1969:BRBa

[1303] Jean Théodoridès. Book review: Medicine, Science, and Culture. Historical Essays in Honor of Owsei Temkin by Lloyd G. Stevenson; Robert P. Multhauf. Isis, 60(3):404–405, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229504.

Gericke:1969:BRB

[1304] H. Gericke. Book review: Greek Mathematical Philosophy by Edward A. Maziarz; Thomas Greenwood. Isis, 60(3):406, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229505.

McVaugh:1969:BRB

[1305] Michael McVaugh. Book review: Fachliteratur des Mittelalters. Festschrift für Gerhard Eis by Gundolf Keil; Rainer Rudolf; Wolfram Schmitt; Hans J. Vermeer. Isis, 60(3):406–407, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229506.

Eastwood:1969:BRB

[1306] Bruce S. Eastwood. Book review: Ontologie oder Metaphysik? Die Diskussion über den Gegenstand der Metaphysik im 13. und 14. Jahrhundert by Albert Zimmermann. Isis, 60(3):407–408, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229507.

Lorch:1969:BRB

[1307] Jacob Lorch. Book review: Leonhard Rauwolf, Sixteenth-Century Physician, Botanist, and Traveler by Karl H. Dannenfeldt. Isis, 60(3):408– 409, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229508.

Bos:1969:BRB

[1308] H. J. M. Bos. Book review: Petrus Ramus en de Wiskunde by J. J. Verdonk. Isis, 60(3):409–410, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229509.

Morris:1969:BRB

[1309] Robert J. Morris, Jr. Book review: The Collected Works of Count Rumford. Vol. I: The Nature of Heat by Benjamin Thompson; Count Rumford; Sanborn C. Brown. Isis, 60(3):410–411, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229510.

Johnson:1969:BRB

[1310] Robert E. Johnson. Book review: Daniel Carl Solander. Naturalist on the “Endeavour.” by Roy Anthony Rauschenberg. Isis, 60(3):411–412, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229511.

Heilbron:1969:BRB

[1311] John L. Heilbron. Book review: The Apparatus of Science at Harvard 1765–1800 by David P. Wheatland; Barbara Carson. Isis, 60(3):412– 413, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229512.

Loewenberg:1969:BRB

[1312] Bert James Loewenberg. Book review: Darwin and Henslow: The Growth of an Idea. Letters 1831–1860 by Nora Barlow. Isis, 60(3):413– 414, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229513.

Stuewer:1969:BRB

[1313] Roger H. Stuewer. Book review: The House on College Avenue: The Comptons at Wooster, 1891–1913 by James R. Blackwood. Isis, 60(3): 414–415, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229514.

Klein:1969:BRB

[1314] Marc Klein. Book review: Catalogue des manuscrits de Claude Bernard by M. D. Grmek. Isis, 60(3):415–417, Autumn 1969. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229515.

Burstyn:1969:BRBa

[1315] Harold L. Burstyn. Book review: Government in Science. The U.S. Geological Survey 1867–1894 by Thomas G. Manning. Isis, 60(3):417– 418, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229516.

Ravetz:1969:BRB

[1316] J. R. Ravetz. Book review: Science in Modern Society by J. G. Crowther. Isis, 60(3):418–419, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229517.

Badash:1969:BRB

[1317] Lawrence Badash. Book review: History and Development of Radiology in Denmark, 1896–1950 by P. Flemming Møller; A. Engelbreth-Holm. Isis, 60(3):419, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229518.

Anonymous:1969:BMc

[1318] Anonymous. Back matter. Isis, 60(3):420–422, Autumn 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229519.

Anonymous:1969:FMd

[1319] Anonymous. Front matter. Isis, 60(4):423–425, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229102.

Temkin:1969:HRS

[1320] Owsei Temkin. Historical reflections on the Scientist’s virtue. Isis, 60 (4):427–438, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229103.

Schneer:1969:EEF

[1321] Cecil J. Schneer. Ebenezer emmons and the foundations of American geology. Isis, 60(4):439–450, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229104.

Morris:1969:PRD

[1322] John Morris. Pattern recognition in Descartes’ Automata. Isis, 60(4): 451–460, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229105.

Zapffe:1969:GHP

[1323] Carl A. Zapffe. Gustavus Hinrichs, precursor of Mendeleev. Isis, 60 (4):461–476, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229106.

Kapoor:1969:OLO

[1324] Satish C. Kapoor. The origins of Laurent’s organic classification. Isis, 60(4):477–527, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229107.

Boyer:1969:EWH

[1325] Carl B. Boyer. Éloge: William Harris Stahl, 1908–1969. Isis, 60(4):528– 534, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229108.

Thiele:1969:KPE

[1326] Joachim Thiele. Karl Pearson, Ernst Mach, John B. Stallo: Briefe aus den Jahren 1897 bis 1904. (German) [Karl Pearson, Ernst Mach, John B. Stallo: Correspondence from the years 1897 to 1904]. Isis, 60(4):535– 542, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229109.

Kennedy:1969:CCI

[1327] E. S. Kennedy and Walid Ukashah. The chandelier clock of ibn Yūnis. Isis, 60(4):543–545, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229110.

Goran:1969:PPP

[1328] Morris Goran. The proxy parents of pioneer physical scientists. Isis, 60(4):545–546, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229111.

Rabinovitch:1969:BC

[1329] N. L. Rabinovitch. “Berossos the Chaldean”. Isis, 60(4):547, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229112. See [942, 1110].

Anonymous:1969:Nc

[1330] Anonymous. News. Isis, 60(4):548–549, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229113.

Hufbauer:1969:BRB

[1331] Karl Hufbauer. Book review: Lawrence and Oppenheimer. Isis, 60(4): 550–551, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229114.

Theodorides:1969:BRBb

[1332] Jean Théodoridès. Book review: The Art of Organic Forms by Philip C. Ritterbush. Isis, 60(4):551–553, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229115.

Hofmann:1969:BRB

[1333] J. E. Hofmann. Book review: A History of Mathematics by Carl B. Boyer. Isis, 60(4):553, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229116.

Smith:1969:BRBc

[1334] Cyril Stanley Smith. Book review: History of Analytical Chemistry by Ferenc Szabadváry; Gyula Svehla. Isis, 60(4):553–554, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229117.

Nelson:1969:BRB

[1335] Clifford M. Nelson. Book review: Men and Dinosaurs. The Search in Field and Laboratory by Edwin H. Colbert. Isis, 60(4):554–556, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229118.

Rosen:1969:BRBb

[1336] George Rosen. Book review: La Folie à travers les siècles by Michèle Ristich De Groote. Isis, 60(4):556, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229119.

Ferguson:1969:BRB

[1337] Eugene S. Ferguson. Book review: American Building: Materials and Techniques from the First Colonial Settlements to the Present by Carl W. Condit. Isis, 60(4):556–558, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229120.

Smith:1969:BRBd

[1338] N. A. F. Smith. Book review: Bibliography of the History of Technology by Eugene S. Ferguson. Isis, 60(4):558–559, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229121.

Tsouyopoulos:1969:BRB

[1339] Nelly Tsouyopoulos. Book review: Johannes Philoponos, Grammatikos von Alexandrien (6. Jh. n. Chr.) by Walter Böhm. Isis, 60(4):559–561, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229122.

ONeill:1969:BRB

[1340] Ynez Violé O’Neill. Book review: Problemata Varia Anatomica. The University of Bologna MS 1165 by L. R. Lind. Isis, 60(4):561–562, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229123.

Wallace:1969:BRB

[1341] William A. Wallace. Book review: Le Livre du ciel et du monde by Nicole Oresme; Albert D. Menut; Alexander J. Denomy; Nicole Oresme and the Medieval Geometry of Qualities and Motions by Marshall Clagett. Isis, 60(4):562–564, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229124.

Hamarneh:1969:BRB

[1342] Sami Hamarneh. Book review: Das Fortleben der Antike im Islam by Franz Rosenthal; Arabische Geisteswelt by Charles Pellat. Isis, 60(4): 564–565, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229125.

Kren:1969:BRB

[1343] Claudia Kren. Book review: Mechanics in Sixteenth-Century Italy. Selections from Tartaglia, Benedetti, Guido Ubaldo, and Galileo by Stillman Drake; I. E. Drabkin. Isis, 60(4):565–566, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229126.

MacClintock:1969:BRB

[1344] Stuart MacClintock. Book review: Gianfrancesco Pico della Mirandola (1469–1533) and His Critique of Aristotle by Charles B. Schmitt. Isis, 60(4):566–567, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229127.

Gingerich:1969:BRB

[1345] Owen Gingerich. Book review: Bibliographia Kepleriana by Max Caspar. Isis, 60(4):567–568, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229128.

Bodemer:1969:BRB

[1346] Charles W. Bodemer. Book review: Investigations into Generation, 1651–1828 by Elizabeth Gasking. Isis, 60(4):568–569, Winter 1969. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229129.

Chappell:1969:BRB

[1347] John E. Chappell, Jr. Book review: The Figure of the Earth by Tom B. Jones. Isis, 60(4):569–570, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229130.

Davis:1969:BRBb

[1348] Audrey B. Davis. Book review: Joachim Jungius’ Experimente und Gedanken zur Begründung der Chemie als Wissenschaft. Ein Beitrag zur Geistesgeschichte des 17. Jahrhunderts by Hans Kangro. Isis, 60(4):570– 572, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229131.

Burstyn:1969:BRBb

[1349] Harold L. Burstyn. Book review: Atomism in England from Hariot to Newton by Robert Hugh Kargon. Isis, 60(4):572–575, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229132.

Rauschenberg:1969:BRB

[1350] Roy A. Rauschenberg. Book review: The Essential Writings of Erasmus Darwin by Desmond King-Hele. Isis, 60(4):575–576, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229133.

Westfall:1969:BRB

[1351] Richard S. Westfall. Book review: The Mathematical Principles of Natural Philosophy by Isaac Newton; Andrew Motte. Isis, 60(4):576, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229134.

Schiller:1969:BRB

[1352] J. Schiller. Book review: Jan Swammerdam (16 February 1637–17 February 1680). His Life and Works by A. Schierbeek. Isis, 60(4):576–577, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229135.

Dexter:1969:BRB

[1353] Ralph W. Dexter. Book review: The Indians of Texas in 1830 by Jean Louis Berlandier; John C. Ewers; Patricia Reading Leclercq. Isis, 60 (4):577–578, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229136.

Hannaway:1969:BRB

[1354] Owen Hannaway. Book review: The Society of Arcueil. A View of French Science at the Time of Napoleon I by Maurice Crosland. Isis, 60(4):578– 581, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229137.

Cannon:1969:BRB

[1355] Walter F. Cannon. Book review: Herschel at the Cape. Diaries and Correspondence, 1834–1838 by John Herschel; David S. Evans; Terence J. Deeming; Betty Hall Evans; Stephen Goldfarb. Isis, 60(4):581–583, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229138.

Burstyn:1969:BRBc

[1356] Harold L. Burstyn. Book review: The Admiralty Chart. British Naval Hydrography in the Nineteenth Century by G. S. Ritchie. Isis, 60(4):583– 585, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229139.

Carozzi:1969:BRB

[1357] Albert V. Carozzi. Book review: Naturalisti-Esploratori dell’Ottocento Italiano. Antologia scientifica e letteraria by Francesco Rodolico. Isis, 60(4):585–587, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229140.

Coleman:1969:BRBb

[1358] William Coleman. Book review: On University Studies by F. W. J. Schelling; Ella S. Morgan; Norbert Guterman. Isis, 60(4):587–588, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229141.

Forman:1969:BRB

[1359] Paul Forman. Book review: Deutsche Wissenschaft und internationale Zusammenarbeit 1914–1928. Ein Beitrag zum Studium kultureller Beziehungen in politischen Krisenzeiten by Brigitte Schröder–Gudehus. Isis, 60(4):588–590, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229142.

Abrams:1969:BRB

[1360] John W. Abrams. Book review: The U.S. Machine Tool Industry from 1900–1950 by Harless D. Wagoner. Isis, 60(4):590, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229143.

Anonymous:1969:BMd

[1361] Anonymous. Back matter. Isis, 60(4):591–602, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229144.

Anonymous:1969:VI

[1362] Anonymous. Volume information. Isis, 60(4):593–598, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229101.

Anonymous:1969:FMe

[1363] Anonymous. Front matter. Isis, 60(5):1–4, ???? 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229588.

Neu:1969:NFC

[1364] John Neu. Ninety-fourth critical bibliography of the history of science and its cultural influences (to January 1969). Isis, 60(5):5–197+199– 219, ???? 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229589.

Anonymous:1969:BMe

[1365] Anonymous. Back matter. Isis, 60(5):??, ???? 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

      jstor.org/stable/229590.